### v1

# Discrete $N$-particle systems at high temperature through Jack generating functions

Abstract.
We find necessary and sufficient conditions for the Law of Large Numbers for random discrete $N$-particle systems with the deformation (inverse temperature) parameter $\theta$, as their size $N$ tends to infinity simultaneously with the inverse temperature going to zero. Our conditions are expressed in terms of the Jack generating functions, and our analysis is based on the asymptotics of the action of Cherednik operators obtained via Hecke relations. We apply the general framework to obtain the LLN for a large class of Markov chains of $N$ nonintersecting particles with interaction of log-gas type, and the LLN for the multiplication of Jack polynomials, as the inverse temperature tends to zero. We express the answer in terms of novel one-parameter deformations of cumulants and their description provided by us recovers previous work by Bufetov–Gorin [BG15] on quantized free cumulants when $\theta=1$, and by Benaych-Georges–Cuenca–Gorin [BGCG22] after a deformation to continuous space of random matrix eigenvalues. Our methods are robust enough to be applied to the fixed temperature regime, where we recover the LLN of Huang [Hua21].
CC was partially supported by the NSF grant DMS-2348139 and by the Simons Foundation’s Travel Support for Mathematicians grant MP-TSM-00006777. This research was funded in whole or in part by Narodowe Centrum Nauki, grant 2021/42/E/ST1/00162. For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
1. Introduction
1.1. Overview
Motivated by classical problems in statistical physics and their deep connections with random matrix theory, understanding the global asymptotics of various random $N$-particle systems $x_{1}<\dots<x_{N}$ in the real line, at inverse temperature $\beta>0$, has been a central question in probability theory for the last few decades. When the parameter $\beta$ is one of the values $1,2,4$, the random $N$-particle systems usually correspond to the spectra of self-adjoint random matrices. The first insight into their global asymptotics traces back to Wigner’s discovery [Wig58] of the semicircle distribution as the limiting empirical measure for the Gaussian Unitary Ensemble, which corresponds to a specific model at $\beta=2$. In many cases, it has been shown that the global asymptotics of these random $N$-particle systems exhibit the following universality: they display the same behavior regardless of the actual value of $\beta$, as long as it remains fixed. Another important discovery, relevant to us, is due to Voiculescu [Voi86, Voi91], who showed that the behavior of large independent random matrices can be captured by freely independent random variables. In particular, the spectrum of the sum of large independent uniformly-rotated matrices with fixed eigenvalues is described by the free convolution of probability measures.
In the limiting case $\beta=0$, the interaction between random particles disappears and they become independent; then the classical Law of Large Numbers describes the system when the number of particles $N$ tends to infinity. Moreover, in the regime where $\beta$ tends to zero at the same time as $N$ grows, one can observe phenomena interpolating between classical and free probability. This middle ground is known as the high temperature regime and occurs when $\beta N$ tends to a positive constant $\gamma$. The high temperature LLN for the Hermite and Laguerre $\beta$-ensembles, as well as the $\gamma$-deformation of the semicircle and Marchenko-Pastur distributions, were studied in [ABG12, ABMV13, DS15, TT21]. More general ensembles and $\gamma$-dependent limits, including $\beta$-sums of random matrices, were studied in [BGCG22] (see also [MP22]), where the authors provided necessary and sufficient conditions for the high temperature LLN.
To the best of our knowledge, and in sharp contrast with the continuous case, the high temperature regime of the global asymptotics of discrete random $N$-particle systems at inverse temperature $\beta>0$ has not been studied in the literature. Other discrete models of $\beta$-ensembles, however, have only recently been studied in the high and low-temperature regimes, see [DS19, Mol23, CDM23], but the case of discrete random $N$-particle systems was the remaining case to be understood. In this paper, we fill this gap by providing necessary and sufficient conditions for the LLN for a very general class of discrete random $N$-particle systems in the high temperature regime.
1.2. Main results
Motivated by the previous work on discrete random $N$-particle systems at inverse temperature $\beta>0$ (see [GS15, BGG17, Hua21] and references therein), we consider the following model. Let $\{\mathbb{P}_{N}\}_{N\geq 1}$ be a sequence of finite signed measures on the set of $N$-tuples $(\mathscr{L}_{1}>\dots>\mathscr{L}_{N})$ of the form $\mathscr{L}_{i}:=\lambda_{i}-(i-1)\theta$, for $i=1,2,\dots,N$, where
$\lambda=(\lambda_{1},\dots,\lambda_{N})\in\mathbb{S}(N):=\{\lambda=(\lambda_{1}\geq\cdots\geq\lambda_{N})\in\mathbb{Z}^{N}\}$
belongs to the set of highest weights of irreducible $U(N)$-representations, also known as $N$-signatures. Clearly, $\mathbb{P}_{N}$ can also be interpreted as a finite signed measure on the set of $N$-signatures, and we will use this interpretation. At this point, the measure $\mathbb{P}_{N}$ is as general as it can be, and in particular it might [^1] It is more interesting when it does depend on the parameter $\theta$, and in the examples we will consider later this is always the case. In the context of the discrete log-gas model, the parameter $\theta$ is related to the parameter $\beta$ by $2\theta=\beta$. Moreover, most of our applications concern probability measures, but our main result is more general and applies to finite signed measures. or might not depend on the parameter $\theta$. We are interested in the behavior of the empirical measures
[…]
In this paper, we find explicit necessary and sufficient conditions for the empirical measures $\mu_{N}$ to converge (either weakly, or in the sense of moments) to some probability measure, as $N\to\infty$, $N\theta\to\gamma\in\mathbb{R}_{>0}$. The conditions are expressed in terms of certain discrete Fourier-type transforms of $\mathbb{P}_{N}$, called the Jack generating functions.
To motivate the definition, let us mention at this point that in the special case $\theta=1$, the underlying $N$-particles can be treated as random highest weights of irreducible $U(N)$-representations. The Fourier transform on the unitary group $U(N)$ that allows to control natural representation-theoretic operations as $N\to\infty$ is given by the so-called Schur generating function, which turned out to be a very convenient tool to approach many interesting classes of interacting particles beyond representation theoretic models, for instance random tilings (see e.g. [BG15, BG18, BG19]). The Fourier transform on the unitary group $U(N)$ can be naturally interpreted as the Fourier transform on the Gelfand pair $(\operatorname{GL}_{N}(\mathbb{C}),\operatorname{U}_{N}(\mathbb{C}))$, and by switching to classical (skew) fields like $\mathbb{R}$ or the quaternions $\mathbb{H}$, one can study other $N$-particle systems of representation-theoretic origin corresponding to the Gelfand pairs $(\operatorname{GL}_{N}(\mathbb{R}),\operatorname{O}_{N}(\mathbb{R}))$ and $(\operatorname{GL}_{N}(\mathbb{H}),\operatorname{U}_{N}(\mathbb{H}))$, through their Fourier transforms. The Jack generating function is equipped with a continuous parameter $\theta>0$, and interpolates between these three Fourier transforms at $\theta=1,1/2,2$, respectively. Our main result explains that the germ at unity of the Jack generating function contains complete information about the Law of Large Numbers not only for a fixed parameter $\theta>0$, as considered in [Hua21], but also in the high temperature regime.
Another reason for the significance of the Jack generating function is that it can be explicitly computed for a natural subclass of Jack measures (see Definition 5.3). The Jack measures are one-parameter generalizations of the Schur measures [Oko01] and have been studied in connection to asymptotic representation theory [BO05, CDM23], Markov processes with interaction of log-gas type [GS15, Hua21], etc. They generally depend on two specializations $\rho,\rho^{\prime}$ of $\operatorname{Sym}$, but if one of them is $\rho^{\prime}=(1^{N})$, then the corresponding Jack measures are supported on the set $\mathbb{Y}(N)=\{\lambda=(\lambda_{1}\geq\dots\geq\lambda_{N})\in(\mathbb{Z}_{\geq 0})^{N}\}$ of partitions with at most $N$ parts. Natural choices for the other specialization $\rho$ are parametrized by points belonging to the Thoma cone, and we are able to prove the LLN at high temperature due to the fact that the Jack generating function have explicit factorized forms and the conditions for LLN are easily verified. We will furter motivate the definition of the Jack generating function in the following section, where we discuss applications of our criterion for the LLN.
Let $P_{\lambda}(x_{1},\dots,x_{N};\theta)$ be the Jack symmetric polynomial [^2] See Section 2 for a review of the necessary background of the theory of Jack symmetric polynomials. that depends on the Jack parameter $\theta>0$. We associate to any measure $\mathbb{P}_{N}$ on $\mathbb{S}(N)$ the following formal multivariate symmetric power series, to be called the Jack generating function of $\mathbb{P}_{N}$:
(1) $G_{\mathbb{P}_{N},\theta}(x_{1},\dots,x_{N}):=\sum_{\lambda\in\mathbb{S}(N)}{\mathbb{P}_{N}(\lambda)\frac{P_{\lambda}(x_{1},\dots,x_{N};\theta)}{P_{\lambda}(1^{N};\theta)}}.$
Our first main result (Theorem 3.6 in the text) establishes necessary and sufficient conditions for the Law of Large Numbers of particles $(\mathscr{L}_{1}>\dots>\mathscr{L}_{N})$ associated with a sequence of finite signed measures $\{\mathbb{P}_{N}\}_{N\geq 1}$ on $N$-signatures in the high temperature regime:
$N\to\infty,\quad\theta\to 0,\quad N\theta\to\gamma\in(0,\infty).$
For simplicity, we break down this result into the next two theorems in this introduction. Informally speaking, the first part states that all the information about the LLN of the empirical measures in the high temperature regime is encoded in the germ of the Jack generating function at the point $(1^{N})$ in this regime.
Theorem 1.1 (First part of Theorem 3.6 in the text).
Let $\{\mathbb{P}_{N}\}_{N\geq 1}$ be a sequence of finite signed measures on $N$-signatures (that satisfies a natural condition of Definition 2.2 on the growth rate of $G_{\mathbb{P}_{N},\theta}$) and let $\{\mu_{N}\}_{N\geq 1}$ be their empirical measures. Then $\mathbb{E}_{\mathbb{P}_{N}}\mu_{N}$ converge in the sense of moments, in the high temperature regime, i.e. there exist $m_{1},m_{2},\dots\in\mathbb{R}$ such that, for all $s\in\mathbb{Z}_{\geq 1}$ and $k_{1},\dots,k_{s}\in\mathbb{Z}_{\geq 1}$, we have
$\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\frac{1}{N^{s}}\,\mathbb{E}_{\mathbb{P}_{N}}\left[\prod_{j=1}^{s}{\sum_{i=1}^{N}{\mathscr{L}_{i}^{k_{j}}}}\right]=\prod_{j=1}^{s}{m_{k_{j}}},$
if and only if the following two conditions are satisfied, for some constants $\kappa_{1}^{\gamma},\kappa_{2}^{\gamma},\cdots$:
  1. (1)
    $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\frac{1}{(n-1)!}\frac{\partial^{n}}{\partial x_{1}^{n}}\ln\big{(}G_{\mathbb{P}_{N},\theta}\big{)}\Big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\kappa_{n}^{\gamma}$ exists and is finite, for all $n\in\mathbb{Z}_{\geq 1}$,
  2. (2)
    $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\frac{\partial^{r}}{\partial x_{i_{1}}\cdots\partial x_{i_{r}}}\ln\big{(}G_{\mathbb{P}_{N},\theta}\big{)}\Big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=0$, for all $r\geq 2$ and $i_{1},\dots,i_{r}\in\mathbb{Z}_{\geq 1}$ with at least two distinct indices among $i_{1},\dots,i_{r}$.
Moreover, if $\mathbb{P}_{N}$ are probability measures and $\sup_{n\geq 1}\big{|}\kappa_{n}^{\gamma}\big{|}^{1/n}<\infty$, then $\mu_{N}$ converge in the sense of moments, in probability, to a probability measure $\mu_{\gamma}$ uniquely determined by its moments.
The values $\kappa_{1}^{\gamma},\kappa_{2}^{\gamma},\cdots$ from part (1) of Theorem 1.1 will be called the quantized $\gamma$-cumulants of the limiting measure $\mu_{\gamma}$; they uniquely determine and, at the same time, are determined by the moments $m_{1},m_{2},\dots$ of $\mu_{\gamma}$. The second part of our main theorem provides a combinatorial transform for calculating the moments of the limiting measure in terms of the quantized $\gamma$-cumulants. It refines several classical and recently discovered moment-cumulant-type formulas and fits into the universal class of combinatorial formulas expressed as the weighted generating function of Łukasiewicz lattice paths (see Section 1.4 for references on these other works). The exact relation is the following.
Theorem 1.2 (Second part of Theorem 3.6 in the text).
Assume that we are in the setting of Theorem 1.1. If $(m_{\ell})_{\ell\geq 1}$ is the sequence of moments of $\mu_{\gamma}$, then
(2) $m_{\ell}=\sum_{\Gamma\in\mathbf{L}(\ell)}\frac{\Delta_{\gamma}\left(x^{1+\text{\#\,horizontal steps at height $0$ in $\Gamma$}}\right)\big{|}_{x=\kappa_{1}^{\gamma}}}{1+\text{\#\,horizontal steps at height $0$ in $\Gamma$}}\cdot\prod_{i\geq 1}\big{(}\kappa_{1}^{\gamma}+i\big{)}^{\text{\#\,horizontal steps at height $i$ in $\Gamma$}}\\ \cdot\prod_{j\geq 1}\big{(}\kappa_{j}^{\gamma}+\kappa_{j+1}^{\gamma}\big{)}^{\text{\#\,steps $(1,j)$ in $\Gamma$}}(j+\gamma)^{\text{\#\,down steps from height $j$ in $\Gamma$}},$
for all $\ell\in\mathbb{Z}_{\geq 1}$, where $\mathbf{L}(\ell)$ is the set of Łukasiewicz paths of length $\ell$ (certain lattice paths in $\mathbb{Z}^{2}$ whose definition is recalled in Definition 3.3), and $\Delta_{\gamma}(f)(x):=\frac{1}{\gamma}(f(x)-f(x-\gamma))$.
The term quantized $\gamma$-cumulant is used because, upon certain normalization followed by the limit $\gamma\to\infty$, the quantities $\kappa_{n}^{\gamma}$ turn into the coefficients of the quantized $R$-transform from [BG15], as shown in Section 7. We will call the generating series
$R^{\gamma}_{\mu}(z):=\sum_{n\geq 1}{\frac{\kappa_{n}^{\gamma}[\mu]z^{n}}{n}},$
the $\gamma$-quantized $R$-transform of $\mu$. Our Theorem 1.2 provides a combinatorial recipe for calculating such transform for measures $\mu$ with finite moments of all orders. There is a natural question on how to extend the definition to larger classes of measures. In that generality, Free Probability defines the $R$-transform as an analytic function that satisfies a certain functional equation involving the Cauchy transform. Our next main theorem is a step in this direction and gives an alternative expression between quantized $\gamma$-cumulants and moments in a form of a functional equation between the $\gamma$-quantized $R$-transform and the moment generating function. We believe that this formal power series identity can become a key tool for extending the $\gamma$-quantized $R$-transform to measures with heavy tails in future research.
Theorem 1.3 (Theorem 6.1 in the text).
Assume that the sequences $(m_{n})_{n\geq 1}$ and $\big{(}\kappa^{\gamma}_{n}\big{)}_{n\geq 1}$ satisfy the relations (2). Then there exists a unique sequence $\big{(}c^{\gamma}_{n}\big{)}_{n\geq 1}$ such that we have the following equalities of formal power series in $z$ and $z^{-1}$, respectively,
(3) $\exp\left(\sum_{n\geq 1}\frac{\kappa_{n}^{\gamma}z^{n}}{n}\right)$ $=1+\sum_{n\geq 1}\frac{c_{n}^{\gamma}}{\gamma^{\uparrow n}}z^{n},$
$1+\sum_{n\geq 1}\frac{(-1)^{n}c_{n}^{\gamma}}{z^{\uparrow n}}$ $=\exp\left(\gamma\cdot\widetilde{\mathcal{L}}\left\{\frac{t}{1-e^{-t}}\sum_{n\geq 1}\left(\frac{(-1)^{n}m_{n}}{n!}-\frac{\gamma^{n}}{(n+1)!}\right)t^{n-1}\right\}(z)\right),$
where $z^{\uparrow n}:=\prod_{i=0}^{n-1}{(z+i)}$ and ~$\widetilde{\mathcal{L}}\left\{\sum_{n\geq 0}{\frac{a_{n}}{n!}t^{n}}\right\}(z):=\sum_{n\geq 0}{a_{n}z^{-n-1}}$.
In the following, we discuss various applications of our results.
1.3. Applications
1.3.1. Quantized $\gamma$-convolution and the Jack Littlewood–Richardson coefficients
Recall that classical cumulants linearize the classical convolution of two probability measures. Similarly, free cumulants linearize the free convolution of probability measures, which, as shown by Voiculescu in his seminal papers [Voi86, Voi91], describes the spectrum of the sum of large independent uniformly-rotated matrices. An analogous result was shown by Bufetov and Gorin [BG15], who demonstrated that their quantized $R$-transform linearizes the operation of decomposing the tensor product of two random irreducible representations of the unitary group $U(N)$ into irreducible components, and they established the LLN for this operation as $N\to\infty$. We will show that $\gamma$-deformed cumulants linearize the Jack-deformed version of the aforementioned operation of tensoring representations, and we will prove the associated LLN in the high temperature regime. Interestingly, we will show that the existence of quantized $\gamma$-convolution for probability measures is strictly related to the long-standing open problem posed by Stanley [Sta89] regarding the structure constants associated to the Jack symmetric functions.
Let $\{\mathbb{P}_{N}^{(1)}\}_{N\geq 1}$ and $\{\mathbb{P}_{N}^{(2)}\}_{N\geq 1}$ be sequences of probability measures on $N$-signatures that satisfy the growth-rate condition from Definition 2.2, and such that (see Section 5.1 for more details):
  • Their empirical measures $\big{\{}\mu_{N}^{(1)}\big{\}}_{N\geq 1}$ and $\big{\{}\mu_{N}^{(2)}\big{\}}_{N\geq 1}$ are supported on the same compact set in the high temperature regime.
  • $\mu_{N}^{(1)}\to\mu^{(1)}$ and $\mu_{N}^{(2)}\to\mu^{(2)}$ as $N\to\infty$, weakly, in probability.
Let $\kappa_{n}^{\gamma}[\mu^{(1)}]$ and $\kappa_{n}^{\gamma}[\mu^{(2)}]$ be the quantized $\gamma$-cumulants of $\mu^{(1)}$ and $\mu^{(2)}$, respectively. We are interested in whether the sequence of sums
$\tilde{\kappa}_{n}^{\gamma}:=\kappa_{n}^{\gamma}[\mu^{(1)}]+\kappa_{n}^{\gamma}[\mu^{(2)}],\quad n\in\mathbb{Z}_{\geq 1},$
and the associated sequence $\tilde{m}_{1},\tilde{m}_{2},\dots$ of ”moments” obtained from $\big{(}\tilde{\kappa}_{n}^{\gamma}\big{)}_{n\geq 1}$ by the combinatorial formula (2) correspond to some measure that arises as a limit of certain natural ”convolution” in the high temperature regime.
Define the finite signed measures $\mathbb{P}_{N}^{(1)}\boxplus_{\theta}\mathbb{P}_{N}^{(2)}$ on $\mathbb{S}(N)$ by
$\mathbb{P}_{N}^{(1)}\boxplus_{\theta}\mathbb{P}_{N}^{(2)}(\lambda):=\sum_{\mu,\,\nu\in\mathbb{Y}(N)}c^{\lambda}_{\mu,\,\nu}(\theta)\mathbb{P}_{N}^{(1)}(\mu)\mathbb{P}_{N}^{(2)}(\nu),\qquad\lambda\in\mathbb{S}(N),$
where $c^{\lambda}_{\mu,\,\nu}(\theta)$ are the Jack Littlewood–Richardson coefficients, defined by Eq. 82. Equivalently, $\mathbb{P}_{N}^{(1)}\boxplus_{\theta}\mathbb{P}_{N}^{(2)}$ is the unique measure on $\mathbb{S}(N)$ whose Jack generating function is the product of Jack generating functions of $\mathbb{P}_{N}^{(1)}$ and $\mathbb{P}_{N}^{(2)}$. We prove in Section 5.1 that $\mathbb{E}_{\mathbb{P}_{N}^{(1)}\boxplus_{\theta}\mathbb{P}_{N}^{(2)}}\mu_{N}$ converges in the sense of moments, in the high temperature regime, and the limiting moment sequence is $\tilde{m}_{1},\tilde{m}_{2},\cdots$. Moreover, if we additionally knew that $\mathbb{P}_{N}^{(1)}\boxplus_{\theta}\mathbb{P}_{N}^{(2)}$ is a sequence of probability measures, then we would have a natural operation of quantized $\gamma$-convolution that produces a probability measure $\mu^{(1)}\boxplus^{(\gamma)}\mu^{(2)}$ uniquely determined by the property that its quantized $\gamma$-cumulants are:
$\kappa_{n}^{\gamma}\big{[}\mu^{(1)}\boxplus^{(\gamma)}\mu^{(2)}\big{]}=\kappa_{n}^{\gamma}\big{[}\mu^{(1)}\big{]}+\kappa_{n}^{\gamma}\big{[}\mu^{(2)}\big{]},\quad\text{for all }n\in\mathbb{Z}_{\geq 1}.$
In that case, the empirical measures $\mu_{N}$ of $\mathbb{P}_{N}^{(1)}\boxplus_{\theta}\mathbb{P}_{N}^{(2)}$ would converge in moments, in probability to the convolution $\mu^{(1)}\boxplus^{(\gamma)}\mu^{(2)}$.
One of the celebrated conjectures of Stanley [Sta89] states that the Jack Littlewood–Richardson coefficients are (up to a proper normalization) polynomials in $\theta$ with nonnegative integer coefficients. This conjecture is much stronger than the assumption that $\mathbb{P}_{N}^{(1)}\boxplus_{\theta}\mathbb{P}_{N}^{(2)}$ are probability measures. Even though Stanley’s conjecture is still wide open, we will actually show the existence of the quantized $\gamma$-convolution for a large class of measures. In fact, next we demonstrate that it naturally appears in the study of global asymptotics of the law of certain Markov chains on $N$ nonintersecting particles with log-gas type interaction in the high temperature regime.
1.3.2. Nonintersecting particle systems in the high temperature regime through Jack measures
In this section, we assume that we have a sequence of Jack parameters $\theta_{N}>0$ such that $N\theta_{N}\to\gamma$, as $N\to\infty$, for some $\gamma>0$, as well as a sequence of specializations $\rho_{N}\colon\operatorname{Sym}\to\mathbb{C}$ such that $P_{\lambda}\big{(}\rho_{N};\theta_{N}\big{)}\geq 0$, for all $\lambda\in\mathbb{Y}(N)$ and all $N\in\mathbb{Z}_{\geq 1}$. Then if the condition given in Definition 5.4 is satisfied, then there exist unique probability measures with weights proportional to
$\mathbb{P}_{N}(\lambda)\propto Q_{\lambda}\big{(}\rho;\theta_{N}\big{)}P_{\lambda}\big{(}1^{N};\theta_{N}\big{)},\quad\lambda\in\mathbb{Y}(N),$
where $Q_{\lambda}$ is defined in (14).
(a) $\theta=1,N=60$
(b) $\theta=\frac{2}{N},N=60$
(c) $\theta=\frac{1}{2N},N=60$
Figure 1. Three simulations of the Gorin–Shkolnikov process with $N=60$ particles and initial configuration $\mathscr{L}^{(0)}_{i}=\theta(1-i)$, for $1\leq i\leq N$. The simulations show the asymptotic behavior in the fixed temperature regime on the left (with $\theta=1$), and in the high temperature regime with $\gamma=2$ in the middle, and with $\gamma=\frac{1}{2}$ on the right. The $x$-axis represents the space and the $y$-axis represents the time.
These measures comprise a distinguished class of Jack measures, see e.g. [BO05, CDM23, GS15]. It is known that the specializations that are nonnegative on Jack symmetric functions are parametrized by points of the form $(\alpha,\beta,\delta)\in(\mathbb{R}_{\geq 0})^{\infty}\times(\mathbb{R}_{\geq 0})^{\infty}\times\mathbb{R}_{\geq 0}$ in the infinite-dimensional Thoma cone (see 2.4 for the precise statement). When $\alpha_{i}=\beta_{i}=0$, for all $i$, and only $\delta$ is nonzero, the associated Jack measure (called the pure-Plancherel Jack measures in Definition 5.16) was used by Gorin–Shkolnikov [GS15] to construct a continuous Markov process on a discrete space of $N$-tuples of particles, which has a diffusive limit to the $\theta$-Dyson Brownian motion, and which can be interpreted as a one-parameter deformation, nonintersecting $N$-particle version of the Poisson random walk on $\mathbb{Z}$ (see Fig. 1 for a sample of this process). Based on [GS15], Huang defined in [Hua21] a large class of discrete Markov chains $\mathscr{L}^{(0)}(N),\mathscr{L}^{(1)}(N),\dots$, of $N$ nonintersecting discrete particles with log-gas type interaction, by replacing the pure Plancherel specialization with more general Jack-positive specializations (see Definitions 5.4 and 5.7), and proved the LLN at a fixed parameter $\theta$ for the joint law at appropriately chosen sequences of ”times” $\{t_{N}\in\mathbb{Z}_{\geq 1}\}_{N\geq 1}$. We will generalize these results by providing an analogous result in the high temperature regime. Moreover, we will show that the limiting measure can be naturally expressed as the quantized $\gamma$-convolution of two probability measures.
Theorem 1.4 (Theorem 5.9 in the text).
Let $\rho=\{\rho_{N}\}_{N\geq 1}$ be a sequence of stable Jack-positive specialization. Under a natural assumption on the limiting behavior (see (95)) of the sequence of ”times” $\{t_{N}\in\mathbb{Z}_{\geq 1}\}_{N\geq 1}$, we have that:
(1) The sequence of Jack measures $\mathbb{P}_{t_{N}\rho_{N};N}^{(\theta_{N})}$ satisfies the LLN and the associated empirical measures $\mu_{N}$ converge in moments, in probability, as $N\to\infty$, to a probability measure $\mu_{\rho}$.
(2) Additionally, let $\mu^{(0)}_{N}$ be the empirical distributions of the law of the initial configuration $\mathscr{L}^{(0)}(N)$ in the Markov chain $\mathscr{L}^{(0)}(N),\mathscr{L}^{(1)}(N),\mathscr{L}^{(2)}(N),\dots$, with a compactly supported weak limit $\mu^{(0)}$ (in probability, as $N\to\infty$). Then the empirical measure of the law of $\mathscr{L}^{(t_{N})}(N)$ from the Markov chain converges in moments, in probability, as $N\to\infty$, to the quantized $\gamma$-convolution $\mu^{(0)}\boxplus^{(\gamma)}\mu_{\rho}$.
We specifically discuss the high temperature LLN above for the three kinds of “pure specializations”, namely, the case when each $\rho_{N}$ depends only on finitely many parameters of the same kind ($\alpha_{i}$’s, or $\beta_{i}$’s, or $\delta$). The associated Markov chains can be interpreted as one-parameter deformations, nonintersecting $N$-particle versions of the geometric, Bernoulli, and Poisson random walks on $\mathbb{Z}$. While the LLN in the generic, but fixed $\theta$ regime, proved by Huang [Hua21], did not depend on the parameter $\theta$ after the appropriate scaling, we will see that in the high temperature regime the limiting law depends on the parameter $\gamma=\lim_{N}{N\theta_{N}}$ describing the high temperature scaling (compare the left with the middle and the right pictures in Fig. 1).
Let us also mention that the pure-Plancherel Jack measure is a discrete $\beta$-ensemble, in the sense of [BGG17]. That paper proves the LLN for fixed $\theta$, but does not consider the high temperature regime; moreover, it expresses the limit measures as unique solutions to variational problems, while our characterizations are as unique solutions to certain moment problems. Lastly, we mention that the proof of Theorem 5.17 shows that the limiting measure in the appropriate scaling for the law of the Gorin–Shkolnikov process starting from the empty partition has all vanishing quantized $\gamma$-cumulants, except for the first one. As a result, this theorem can be regarded as a generalization of the classical Wigner theorem with the semicircle distribution as the limit (which has only one free cumulant being nonzero) and of the high temperature limit theorem for Hermite $\beta$-ensembles with the limiting measure having only one nonzero $\gamma$-cumulant; see [BGCG22, MP22].
1.4. Previous related works
Motivated by asymptotic representation theory, Vershik–Kerov [VK77] and Logan–Shepp [LS77] initiated the study of random partitions by proving the limit shape phenomenon for Plancherel-distributed Young diagrams. The connection to random matrices was later hinted at by Kerov [Ker93a, Ker93b], who linked this shape to Wigner’s semicircle distribution. This parallelism between the Plancherel measure and Gaussian Unitary Ensembles has been shown to be much deeper in a series of groundbreaking works [BDJ99, Oko00, BOO00, Joh01], which catalyzed further investigations and insights into the parallelism between random matrices and partitions that continue to this day. In particular, motivated by the one-parameter $\beta$-extrapolations of random (continuous) eigenvalue ensembles, it is natural to ask: what are the parallel one-parameter random partition ensembles?
Many articles have addressed this question by studying several alternatives; see, e.g., [Ker00, BO05, DF16, BGG17, DS19, GH19, Hua21, Mol23, CDM23]. The common theme is that all these partition ensembles are based on the theory of Jack symmetric functions and depend on the Jack parameter $\theta>0$, which is related to the inverse temperature $\beta$ from the continuous world by $\theta=\beta/2$. It is worth noting that the random $\theta$-deformed partitions arise not only as natural discrete versions of $\beta$-eigenvalues, but they also have intrinsic interest in various areas, such as algebraic geometry [Oko03, LQW03, PY24], Seiberg–Witten theory [NO06], etc.
Unlike the continuous case, there are many different models of random partitions, and they differ quite drastically, even regarding the state space: some ensembles are defined on the set of all partitions $\mathbb{Y}$, e.g., the Jack measures [GS15, Mol23]; some on the set $\mathbb{Y}_{d}$ of partitions of size $d$, e.g., the Jack–Plancherel measure [Ker00, DF16], and more general measures from Jack characters with the Approximate Factorization Property [DS19]; finally, some ensembles are defined on the set $\mathbb{Y}(N)$ of partitions with at most $N$ parts, e.g., discrete $\beta$-ensembles [BGG17]. In the case of fixed $\theta$, the LLN and CLT for the global asymptotics of the various models of random $\theta$-partitions have been shown to be independent of the value of $\theta$, as in the continuous case [DF16, BGG17, DS19, Hua21, Mol23, CDM23].
In the high and low temperature regimes, Dołęga–Śniady [DS19] proved LLN and CLT for the global asymptotics of a model of measures on partitions of a fixed size, while Moll [Mol23] proved the LLN for a subclass of Jack measures on the set of all partitions. Further, Cuenca–Dołęga–Moll [CDM23] studied both of these models simultaneously and exposed a deep connection between them, which in particular led to a universal moment-cumulant-type formula deforming the celebrated free version of Speicher [Spe94]. This formula is naturally expressed as a weighted generating function of Łukasiewicz paths (see Definition 3.3), and it fits into the generalized framework of the combinatorial approach of Viennot [Vie85] to orthogonal polynomials.
The universal character of the combinatorial moment-cumulant-type formula from [CDM23] motivates the question of whether a similar universal moment-cumulant-type formula exists for the aforementioned models of random partitions with at most $N$ parts. In this paper, we provide such a formula. The connection between the global asymptotics of random partitions with at most $N$ parts and free probability was observed by Biane [Bia95] and Collins–Śniady [CŚ09] in the context of the tensor product of large unitary representations. However, the asymptotic behavior of their model was studied under a different scaling than the one studied in this paper. In the $\theta=1$ case of the model studied here, Bufetov and Gorin noticed that the free version of cumulants from the previous works of Biane and Collins–Śniady should be replaced by the so-called quantized free cumulants [BG15]. Further, the global asymptotics for a fixed, but generic $\theta$, studied in Huang [Hua21], turned out to be the same as for $\theta=1$ after the appropriate normalization. The aforementioned papers did not provide a combinatorial moment-cumulant formula and described the LLN in analytic terms. In Section 7, we show that the analytic description of LLN from [Hua21], in the regime when $\theta$ is fixed, can be recovered from our combinatorial moment-cumulant formula. It is worth mentioning that a different limiting case of our combinatorial formula recovers another recent moment-cumulant formula from [BGCG22].
1.5. Methodology
The main technique employed in this article is the method of moments. More explicitly, we compute the moments of the prelimit empirical measures $\mu_{N}$ of $\mathbb{P}_{N}$ by extracting them from the Jack generating functions using certain differential-difference operators of representation-theoretic origin and then study the asymptotics of the resulting expression. This general approach was pioneered for random partition ensembles by Bufetov and Gorin [BG15, BG18, BG19], who considered the Schur generating functions and differential operators with the Schur polynomials as eigenfunctions. Some applications and extensions followed swiftly by Bufetov–Knizel [BK18], Gorin–Sun [GS22], etc.
In the continuous framework, Benaych-Georges, Cuenca, and Gorin [BGCG22] utilized the multivariate Bessel functions and Dunkl operators to study the large-scale asymptotics of one-parameter extrapolations of random matrix eigenvalues in the high temperature regime. These ideas have been extended to the BC-type by [Xu23], and they have already found several applications [Yao21, KX24, GXZ24]. In the discrete case, Huang [Hua21] explored similar ideas for ensembles of partitions of a given length, for a generic but fixed finite temperature. He used the Nazarov–Sklyanin operators [NS13] and a different version of Jack generating function than the one used by us, whose expansion in the power-sum basis was crucial, as the Nazarov–Sklyanin operators behave well with respect to this basis.
The present research was mainly motivated by [BGCG22, Hua21, BG15] and, simultaneously, it generalizes them all. On the one hand, the random partition ensembles that we study degenerate in a certain regime to the random eigenvalues considered in [BGCG22]; at the technical level, our Jack generating functions and Cherednik operators also degenerate to the Bessel generating functions and Dunkl operators used in that reference. Moreover, we recover their main result from our proof of the LLN by picking up the highest degree terms in all polynomial equations. On the other hand, though we focus on the high temperature regime, we can use the very same analysis to prove analogues of our results at fixed temperature, allowing us to recover the LLN from [Hua21]. As we have mentioned before, although a version of the Jack generating function (JGF) also appears in that paper, there are at least two advantages of our version that employs polynomials and not symmetric functions: (a) our JGF admits explicit product forms for a wide class of Jack measures, as shown in Section 5.2; (b) the power-sum expansion of the JGF used in [Hua21] was an artifact of the used tools, rather than a natural choice of the monomial basis directly related to the model of partitions with a fixed number of parts; the latter choice allows for a direct analysis, which was not possible in [Hua21] and required many technicalities.
A naive approach of mimicking the analysis of the action of Dunkl operators on the Bessel function performed in [BGCG22] and applied to the action of Cherednik operators on the JGF fails. One of the key ideas in [BGCG22] was a decomposition of Dunkl operators as a sum of operators, whose action on the Bessel functions always produces the same result after the specialization of all variables to zero. In fact, this kind of symmetry was always present explicitly or implicitly in all previous analyses of similar problems. Such symmetry does not hold in the case of Cherednik operators and requires new ideas. Our new key insight is that the representation-theoretic Hecke relations allow us to reduce an analysis of generic terms to the analysis of some special terms, and this reduction is also the source of the appearance of some new ingredients in the moment-cumulant formula (see Section 4.4, and Section 7 for the comparison with the formula from [BGCG22]). We hope that this idea gets exploited for further applications.
Acknowledgements
CC is grateful to Vadim Gorin for many useful conversations over the years. CC would like to thank the Institute of Mathematics of the Polish Academy of Sciences (IMPAN) for hosting him during his research visits, and also the Royal Swedish Academy of Sciences and Institut Mittag–Leffler, where part of this research was conducted during the special program on Random Matrices and Scaling Limits in the Fall of 2024.
2. Jack polynomial theory and Jack generating functions
2.1. Partitions and signatures
In this paper, $N$ denotes a positive integer that will later tend to infinity.
A partition is an infinite sequence $\lambda=(\lambda_{1}\geq\lambda_{2}\geq\cdots)\in(\mathbb{Z}_{\geq 0})^{\infty}$ of weakly decreasing nonnegative integers such that $\lambda_{i}=0$, for all large $i$. The size of $\lambda$ is defined to be $|\lambda|:=\sum_{i\geq 1}{\lambda_{i}}$, while its length is $\ell(\lambda):=\max\{i\geq 1\colon\lambda_{i}\neq 0\}$. For the empty partition $\emptyset=(0,0,0,\cdots)$, the size and length are both zero, by convention. For a partition $\lambda$ with $\ell(\lambda)\leq N$, we can write $\lambda=(\lambda_{1},\dots,\lambda_{N})$, with the understanding that infinitely many zeroes are being omitted. We denote the set of all partitions by $\mathbb{Y}$. For any $N\in\mathbb{Z}_{\geq 1}$, we also denote
$\mathbb{Y}(N):=\{\lambda\in\mathbb{Y}\colon\ell(\lambda)\leq N\}.$
For any $N\in\mathbb{Z}_{\geq 1}$, we say that an $N$-tuple $\lambda=(\lambda_{1}\geq\cdots\geq\lambda_{N})\in\mathbb{Z}^{N}$ of weakly decreasing integers is an $N$-signature. By analogy to partitions, we say that an $N$-signature $\lambda=(\lambda_{1},\cdots,\lambda_{N})$ has length $N$. The set of all $N$-signatures is denoted by
$\mathbb{S}(N):=\{\lambda=(\lambda_{1}\geq\cdots\geq\lambda_{N})\in\mathbb{Z}^{N}\}.$
Observe that $\mathbb{Y}(N)\subseteq\mathbb{S}(N)$.
2.2. The Jack parameter
In this paper, there is a positive real $\theta>0$ that will later tend to zero, simultaneously as $N$ tends to infinity. The parameter $\theta$ is called the Jack parameter, because of its relationship with Jack symmetric polynomials. In the literature, e.g. [Sta89, Mac95], sometimes a different Jack parameter $\alpha>0$ is used; the relation between them is $\theta=\frac{1}{\alpha}$.
To any $N$-signature $\lambda=(\lambda_{1}\geq\cdots\geq\lambda_{N})\in\mathbb{S}(N)$, we associate the $N$-tuple of reals $(\mathscr{L}_{1}>\cdots>\mathscr{L}_{N})$ defined by
(4) $\mathscr{L}_{i}:=\lambda_{i}-(i-1)\theta,\text{ for }i=1,2,\dots,N.$
This gives the following bijection that will be used throughout the paper:
$\mathbb{S}(N)\longleftrightarrow\{(\mathscr{L}_{1},\dots,\mathscr{L}_{N})\in\mathbb{R}^{N}\colon\mathscr{L}_{1}>\dots>\mathscr{L}_{N}\text{ and }\mathscr{L}_{i}+(i-1)\theta\in\mathbb{Z},\text{ for }i=1,\dots,N\}.$
2.3. Jack Laurent polynomials
Given any $N\in\mathbb{Z}_{\geq 1}$, the Jack (symmetric) polynomials $P_{\lambda}(x_{1},\dots,x_{N};\theta)$, for $\lambda\in\mathbb{Y}(N)$, are defined as the unique homogeneous symmetric polynomials of degree $|\lambda|$ of the form
(5) $P_{\lambda}(x_{1},\dots,x_{N};\theta)=x_{1}^{\lambda_{1}}\cdots x_{N}^{\lambda_{N}}+\text{ monomials of the form }c_{\mu}x_{1}^{\mu_{1}}\cdots x_{N}^{\mu_{N}}\\ \text{ with }\mu\leq\lambda\text{ in the lexicographic order},$
that are orthogonal with respect to certain explicit weight function; we do not discuss the details further since they will not be needed, but see [Mac95, Ch. VI.10]. If $\lambda\in\mathbb{Y}$ has $\ell(\lambda)>N$, we use the convention $P_{\lambda}(x_{1},\dots,x_{N};\theta):=0$. We have the index-stability property:
$P_{(\lambda_{1}+1,\dots,\lambda_{N}+1)}(x_{1},\dots,x_{N};\theta)=(x_{1}\cdots x_{N})\cdot P_{\lambda}(x_{1},\dots,x_{N};\theta),\quad\lambda\in\mathbb{Y}(N).$
This property allows us to define the Jack Laurent polynomials, for any $\lambda\in\mathbb{S}(N)$, by taking any $M>0$ with $\lambda_{N}+M>0$, so that $(\lambda_{1}+M,\dots,\lambda_{N}+M)\in\mathbb{Y}(N)$, and setting
(6) $P_{\lambda}(x_{1},\dots,x_{N};\theta):=(x_{1}\cdots x_{N})^{-M}\cdot P_{(\lambda_{1}+M,\dots,\lambda_{N}+M)}(x_{1},\dots,x_{N};\theta).$
The definition (6) does not depend on the choice of $M>0$, because of the index-stability.
From (5), all monomials $x_{1}^{\mu_{1}}\cdots x_{N}^{\mu_{N}}$ in the expansion of the Jack Laurent polynomial $P_{\lambda}(x_{1},\dots,x_{N};\theta)$, $\lambda\in\mathbb{S}(N)$, are such that $\lambda_{1}\geq\max_{i}{\mu_{i}}\geq\min_{i}{\mu_{i}}\geq\lambda_{N}$. We will also need the fact that for any $\theta>0$ and $\lambda\in\mathbb{S}(N)$, the coefficients of $P_{\lambda}(x_{1},\dots,x_{N};\theta)$ are nonnegative, see [Mac95, Ch. VI, (10.12)-(10.13)].
2.4. Cherednik operators
The Cherednik operators are the differential-difference operators $\xi_{1},\dots,\xi_{N}$ acting on functions with $N$ variables $x_{1},\dots,x_{N}$, given by
(7) $\xi_{i}:=x_{i}\partial_{i}-\theta(i-1)+\theta\sum_{j=1}^{i-1}\frac{x_{i}}{x_{i}-x_{j}}(1-s_{i,j})+\theta\sum_{j=i+1}^{N}\frac{x_{j}}{x_{i}-x_{j}}(1-s_{i,j}),$
where $s_{i,j}f(x_{1},\dots,x_{i},\dots,x_{j},\dots,x_{N}):=f(x_{1},\dots,x_{j},\dots,x_{i},\dots,x_{N})$ acts by transposing the variables $x_{i}$ and $x_{j}$, and we denote throughout $\partial_{i}:=\frac{\partial}{\partial x_{i}}$, for simplicity. The operators $\xi_{i}$ were introduced by Cherednik in [Che91].
Evidently, each $\xi_{i}$ preserves the space of $N$-variate polynomials of degree $d$, for any $d\in\mathbb{Z}_{\geq 0}$. They also preserve the space of functions that are analytic in some neighborhood of any point of the form $(x_{0}^{N}):=(x_{0},\dots,x_{0})$, for example $(1^{N})$.
The following theorem is due to Opdam [Opd95, (2.7)]; see also [KS97, Corollary 3.2].
Theorem 2.1.
The Cherednik operators $\xi_{1},\dots,\xi_{N}$ satisfy the following Hecke relations:
(8) $\xi_{i}s_{i}-s_{i}\xi_{i+1}$ $=\theta,$
$\xi_{i+1}s_{i}-s_{i}\xi_{i}$ $=-\theta,$
$\xi_{i}s_{j}-s_{j}\xi_{i}$ $=0,\text{ for all }j\neq i-1,i,$
where $s_{i}$ denotes the transposition $s_{i}:=s_{i,i+1}$ of variables $x_{i}$ and $x_{i+1}$. Moreover, the operators $\xi_{1},\dots,\xi_{N}$ pairwise commute and, for all $i=1,\dots,N$, and all $\lambda\in\mathbb{S}(N)$, we have
(9) $\xi_{i}\,P_{\lambda}(x_{1},\dots,x_{N};\theta)=\mathscr{L}_{i}\cdot P_{\lambda}(x_{1},\dots,x_{N};\theta),$
where $\mathscr{L}_{i}:=\lambda_{i}-(i-1)\theta$.
We point out that (9) is stated for $\lambda\in\mathbb{Y}(N)$ in the references [Opd95, KS97], but this identity can be extended to all $\lambda\in\mathbb{S}(N)$, by using
$\xi_{i}\left[(x_{1}\cdots x_{N})^{-1}\cdot f(x_{1},\dots,x_{N})\right]=(x_{1}\cdots x_{N})^{-1}\cdot\left[-f(x_{1},\dots,x_{N})+(\xi_{i}f)(x_{1},\dots,x_{N})\right]$
and the definition (6) of Jack Laurent polynomials.
2.5. Measures on $N$-signatures and Jack generating functions
In this paper, we will generally consider sequences of finite signed measures $\mathbb{P}_{N}$ on the sets $\mathbb{S}(N)$ of $N$-signatures; moreover, we will always assume that $\mathbb{P}_{N}$ have total mass equal to $1$. Sometimes, for conciseness, we may omit the qualifiers finite or signed or of total mass one, and simply say measures. In applications, however, we will usually restrict ourselves to the case when $\mathbb{P}_{N}$ are probability measures, except in Section 5.1.
Given a finite signed measure $\mathbb{P}_{N}$ on $N$-signatures $\lambda=(\lambda_{1}\geq\dots\geq\lambda_{N})\in\mathbb{S}(N)$, with total mass equal to $1$, define its Jack generating function as the following formal Laurent series on $N$ variables:
(10) $G_{\mathbb{P}_{N},\theta}(x_{1},\dots,x_{N}):=\sum_{\lambda\in\mathbb{S}(N)}\mathbb{P}_{N}(\lambda)\frac{P_{\lambda}(x_{1},\dots,x_{N};\theta)}{P_{\lambda}(1^{N};\theta)}.$
We implicitly assume that the measure $\mathbb{P}_{N}$ might additionally depend on the Jack parameter $\theta$, so that the Jack generating function $G_{\mathbb{P}_{N},\theta}$ is defined with respect to the measure $\mathbb{P}_{N}$ evaluated at $\theta>0$, see Section 5 for examples. Next, we identify a large family of measures $\mathbb{P}_{N}$ for which $G_{\mathbb{P}_{N},\theta}$ is an analytic function in a neighborhood of $(1^{N})$.
Definition 2.2.
We say that the finite signed measure $\mathbb{P}_{N}$ on $\mathbb{S}(N)$ has small tails if the Laurent series
(11) $\sum_{\lambda\in\mathbb{S}(N)}{\mathbb{P}_{N}(\lambda)\frac{P_{\lambda}(z_{1},\dots,z_{N};\theta)}{P_{\lambda}(1^{N};\theta)}}$
converges absolutely on an $N$-dimensional annulus of the form
(12) $A_{N;R}:=\{(z_{1},\dots,z_{N})\in\mathbb{C}^{N}\colon R^{-1}<|z_{i}|<R,\text{ for all }i=1,\dots,N\},$
for some $R>1$.
Example 1.
If $\mathbb{P}_{N}$ is supported on a finite subset of $\mathbb{S}(N)$, then $\mathbb{P}_{N}$ has small tails.
If $\mathbb{P}_{N}$ has small tails, then evidently its Jack generating function $G_{\mathbb{P}_{N},\theta}$ is an analytic function in the annulus (12), and moreover
$G_{\mathbb{P}_{N},\theta}(1^{N})=1,$
since $\mathbb{P}_{N}$ has total mass equal to $1$.
Remark 1.
If $\mathbb{P}_{N}$ is a probability measure with small tails and is supported on the set $\mathbb{Y}(N)\subseteq\mathbb{S}(N)$ of partitions of length $\leq N$, then the series (11) is a power series (not a Laurent series) and it converges absolutely on an $N$-dimensional disk
$D_{N;R}:=\{(z_{1},\dots,z_{N})\in\mathbb{C}^{N}\colon|z_{i}|<R,\text{ for all }i=1,\dots,N\},$
for some $R>1$. This follows from the fact that the coefficients of any Jack symmetric polynomial $P_{\lambda}(z_{1},\dots,z_{N};\theta)$ are nonnegative.
For our examples in Section 5, Definition 2.2 is easy to check.
2.6. Jack symmetric functions
We recall some material from [Mac95, Ch. VI.10], [Sta89]. Let $\operatorname{Sym}$ be the real algebra of symmetric functions, that is, the projective limit of the chain
$\cdots\xrightarrow{\ x_{N+2}=0\ }\mathbb{R}[x_{1},\dots,x_{N},x_{N+1}]^{\mathfrak{S}_{N+1}}\xrightarrow{\ x_{N+1}=0\ }\mathbb{R}[x_{1},\dots,x_{N}]^{\mathfrak{S}_{N}}\xrightarrow{\ x_{N}=0\ }\cdots$
of symmetric polynomials in finitely many variables. Let $\mathbb{Y}:=\bigcup_{N\geq 1}{\mathbb{Y}(N)}$ be the set of all partitions. For any $\lambda\in\mathbb{Y}$, we have the stability property of Jack polynomials
$P_{\lambda}(x_{1},\dots,x_{N},x_{N+1};\theta)\big{|}_{x_{N+1}=0}=P_{\lambda}(x_{1},\dots,x_{N};\theta),$
[…]
(13) $\sum_{\lambda\in\mathbb{Y}}P_{\lambda}(\mathbf{x};\theta)Q_{\lambda}(\mathbf{y};\theta)=\exp\Bigg{\{}\theta\sum_{k\geq 1}{\frac{p_{k}(\mathbf{x})p_{k}(\mathbf{y})}{k}}\Bigg{\}}=:H_{\theta}(\mathbf{x};\mathbf{y}),$
where $\mathbf{x}=(x_{1},x_{2},\dots)$, $\mathbf{y}=(y_{1},y_{2},\dots)$, and
(14) $Q_{\lambda}(;\theta):=\prod_{(i,j)\in\lambda}\frac{(\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+\theta}{(\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+1}\cdot P_{\lambda}(;\theta)$
is the $Q$-normalization of Jack polynomials/symmetric functions. In Eq. 14, we denoted $\lambda^{\prime}$ the conjugate partition to $\lambda$, so that $\lambda_{j}^{\prime}$ is the length of the $j$-th largest column of the Young diagram of $\lambda$.
For any partitions $\lambda,\mu\in\mathbb{Y}$, we also define the skew-Jack symmetric function $Q_{\lambda/\mu}(\mathbf{x};\theta)$ (and analogously, the $P$-normalized version $P_{\lambda/\mu}(\mathbf{x};\theta)$) by the following identity:
(15) $Q_{\lambda}(\mathbf{x},\mathbf{y};\theta):=\sum_{\mu\in\mathbb{Y}}Q_{\mu}(\mathbf{y};\theta)Q_{\lambda/\mu}(\mathbf{x};\theta).$
We remark that the sum in the right hand side of (15) is actually finite, since it is known that the skew-Jack symmetric functions $Q_{\lambda/\mu}(\mathbf{x};\theta)$ vanish unless $\mu\subseteq\lambda$. Moreover, $Q_{\lambda/\mu}(\mathbf{x};\theta)$ coincides with $Q_{\lambda}(\mathbf{x};\theta)$ when $\mu=\emptyset$. We will also need the following skew-Cauchy identity:
(16) $\sum_{\lambda\in\mathbb{Y}}P_{\lambda/\mu}(\mathbf{x};\theta)Q_{\lambda/\rho}(\mathbf{y};\theta)=H_{\theta}(\mathbf{x};\mathbf{y})\sum_{\lambda\in\mathbb{Y}}Q_{\mu/\lambda}(\mathbf{y};\theta)P_{\rho/\lambda}(\mathbf{x};\theta).$
2.7. Jack-positive specializations
Definition 2.3.
A specialization of $\operatorname{Sym}$ is a unital algebra homomorphism $\rho\colon\operatorname{Sym}\to\mathbb{R}$. For any $f\in\operatorname{Sym}$, we denote the image of $f$ under $\rho$ by $f(\rho)$. The specialization $\rho\colon\operatorname{Sym}\to\mathbb{R}$ is said to be Jack-positive if
$P_{\lambda}(\rho;\theta)\geq 0,\quad\text{for all }\lambda\in\mathbb{Y}.$
Note that this definition depends on the value of $\theta$, which is usually understood from the context.
The Cauchy identity (13) turns into a numerical equality if $\mathbf{x},\mathbf{y}$ are replaced by two specializations $\rho_{1},\rho_{2}$, as long as the resulting LHS in (13) converges absolutely. Furthermore, if $\rho_{1},\rho_{2}$ are Jack-positive, then each term $P_{\lambda}(\rho_{1};\theta)Q_{\lambda}(\rho_{2};\theta)$ is nonnegative, hence (13) is a series of nonnegative real numbers that converges to $H_{\theta}(\rho_{1};\rho_{2})$ and this value is strictly positive, as $H_{\theta}(\rho_{1};\rho_{2})\geq P_{\emptyset}(\rho_{1};\theta)Q_{\emptyset}(\rho_{2};\theta)=1>0$.
Theorem-Definition 2.4 ([KOO98]).
Given $\theta>0$ fixed, the set of Jack-positive specializations is in bijection with points of the following Thoma cone:
$\Omega:=\Big{\{}(\alpha,\beta,\delta)\in(\mathbb{R}_{\geq 0})^{\infty}\times(\mathbb{R}_{\geq 0})^{\infty}\times\mathbb{R}_{\geq 0}\ \mid\ \alpha=(\alpha_{1},\alpha_{2},\dots),\ \beta=(\beta_{1},\beta_{2},\dots),\\ \alpha_{1}\geq\alpha_{2}\geq\dots\geq 0,\quad\beta_{1}\geq\beta_{2}\geq\dots\geq 0,\quad\sum_{i=1}^{\infty}(\alpha_{i}+\beta_{i})\leq\delta\Big{\}}.$
For any $\omega=(\alpha,\beta,\delta)\in\Omega$, the corresponding Jack-positive specialization $\rho_{\omega}\colon\operatorname{Sym}\to\mathbb{R}$ is given by
$p_{1}(\rho_{\omega})=\delta,\qquad p_{k}(\rho_{\omega})=\sum_{i=1}^{\infty}\Big{(}\alpha_{i}^{k}+(-\theta)^{k-1}\beta_{i}^{k}\Big{)},\quad k=2,3,\cdots.$
If $\omega=(\alpha,\beta,\delta)\in\Omega$, we say that $\alpha_{i},\beta_{i},\delta$ are the Thoma parameters of $\rho_{\omega}$.
Example 2.
If all Thoma parameters of $\omega=(\alpha,\beta,\delta)\in\Omega$ vanish, except possibly $\alpha_{1},\dots,\alpha_{n}$ and $\delta=\sum_{i=1}^{n}\alpha_{i}$, then we say that $\rho_{\omega}$ is a pure alpha specialization and we denote it by $\operatorname{Alpha}(\alpha_{1},\dots,\alpha_{n})$. For any $f=f(x_{1},x_{2},\dots)\in\operatorname{Sym}$, we have
$f\big{(}\operatorname{Alpha}(\alpha_{1},\dots,\alpha_{n})\big{)}=f(\alpha_{1},\dots,\alpha_{n}),$
i.e. $\operatorname{Alpha}(\alpha_{1},\dots,\alpha_{n})$ specializes $n$ variables $x_{i}\mapsto\alpha_{i}$ and the rest of $x_{i}$’s are set to zero.
Example 3.
If the only possibly nonzero Thoma parameters of $\omega=(\alpha,\beta,\delta)\in\Omega$ are $\beta_{1},\dots,\beta_{n}$ and $\delta=\sum_{i=1}^{n}\beta_{i}$, then we say that $\rho_{\omega}$ is a pure beta specialization and we denote it by $\operatorname{Beta}(\beta_{1},\dots,\beta_{n})$. For any $f\in\operatorname{Sym}$, we have
$f\big{(}\operatorname{Beta}(\beta_{1},\dots,\beta_{n})\big{)}=(\omega_{\theta^{-1}}f)(\theta\beta_{1},\dots,\theta\beta_{n}),$
where $\omega_{\theta^{-1}}$ is the automorphism of $\operatorname{Sym}$ defined by $\omega_{\theta^{-1}}(p_{r}):=(-1)^{r-1}\theta^{-1}p_{r}$, for all $r\geq 1$. This means that $\operatorname{Beta}(\beta_{1},\dots,\beta_{n})$ is the composition of $\omega_{\theta^{-1}}$ and the pure alpha specialization $\operatorname{Alpha}(\theta\beta_{1},\dots,\theta\beta_{n})$.
Example 4.
If all Thoma parameters of $\omega=(\alpha,\beta,\delta)\in\Omega$ vanish, except possibly $\delta$, then we say that $\rho_{\omega}$ is a Plancherel specialization and we denote it by $\operatorname{Planch}(\delta)$. Equivalently, it is uniquely defined by $p_{k}(\operatorname{Planch}(\delta))=\delta\cdot\mathbf{1}_{k=1}$.
For any specializations $\rho_{1},\rho_{2}$ of $\operatorname{Sym}$, we define the union specialization, to be denoted simply $(\rho_{1},\rho_{2})$, by the formula
$p_{k}(\rho_{1},\rho_{2}):=p_{k}(\rho_{1})+p_{k}(\rho_{2}),\quad\text{for all }k\geq 1.$
If $\rho_{1},\rho_{2}$ are Jack-positive specializations, then so is their union $(\rho_{1},\rho_{2})$, by virtue of 2.4. Furthermore, for any specialization $\rho$, define the $n$-fold union $n\rho:=(\underbrace{\rho,\dots,\rho}_{n})$; again, if $\rho$ is Jack-positive, then so is $n\rho$.
3. Statement of the main theorem: Law of Large Numbers at high temperature
We are going to study sequences of finite signed measures $\{\mathbb{P}_{N}\}_{N\geq 1}$ on $N$-signatures in the high temperature regime
(17) $N\to\infty,\quad\theta\to 0,\quad N\theta\to\gamma,$
where $\gamma\in(0,\infty)$ is a fixed constant.
Let us assume for the rest of this paper that all measures $\mathbb{P}_{N}$ have small tails in the sense of Definition 2.2. Then the Jack generating functions, denoted by
(18) $G_{N,\theta}(x_{1},\dots,x_{N}):=\sum_{\lambda\in\mathbb{S}(N)}\mathbb{P}_{N}(\lambda)\frac{P_{\lambda}(x_{1},\dots,x_{N};\theta)}{P_{\lambda}(1^{N};\theta)},$
are analytic in a neighborhood of $(1^{N})$ and $G_{N,\theta}(1^{N})=1$. Consequently, the logarithm
(19) $F_{N,\theta}(x_{1},\dots,x_{N}):=\ln(G_{N,\theta}(x_{1},\dots,x_{N}))$
also defines an analytic function in a neighborhood of $(1^{N})$ and $F_{N,\theta}(1^{N})=0$.
Definition 3.1 (HT-appropriateness).
Let $\{\mathbb{P}_{N}\}_{N\geq 1}$ be a sequence of finite signed measures on $N$-signatures with small tails and let $F_{N,\theta}=F_{N,\theta}(x_{1},\dots,x_{N})$ be the logarithms of the Jack generating functions of $\mathbb{P}_{N}$, i.e. the functions $F_{N,\theta}(x_{1},\dots,x_{N})$ are defined by (18)–(19).
We say that the sequence $\{\mathbb{P}_{N}\}_{N\geq 1}$ is HT-appropriate [^3] HT stands for high temperature. if
  1. (1)
    $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\frac{\partial_{1}^{n}}{(n-1)!}F_{N,\theta}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\kappa_{n}$ exists and is finite, for all $n\in\mathbb{Z}_{\geq 1}$,
  2. (2)
    $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\partial_{i_{1}}\cdots\partial_{i_{r}}F_{N,\theta}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=0$, for all $r\geq 2$ and $i_{1},\dots,i_{r}\in\mathbb{Z}_{\geq 1}$ with at least two distinct indices among $i_{1},\dots,i_{r}$.
We also say that the sequence $\{F_{N,\theta}\}_{N\geq 1}$ is HT-appropriate if these conditions hold.
The measure $\mathbb{P}_{N}$ on $N$-signatures $\lambda=(\lambda_{1}\geq\dots\geq\lambda_{N})\in\mathbb{S}(N)$ can also be regarded as a measure on real $N$-tuples $(\mathscr{L}_{1}>\dots>\mathscr{L}_{N})$, where $\mathscr{L}_{i}+(i-1)\theta\in\mathbb{Z}$, for all $i=1,\dots,N$, by using the shifted coordinates $\mathscr{L}_{i}=\lambda_{i}-(i-1)\theta$, as explained in Section 2.2.
Definition 3.2 (LLN-satisfaction).
Let $\{\mathbb{P}_{N}\}_{N\geq 1}$ be a sequence of finite signed measures on $N$-signatures with small tails. We say that $\{\mathbb{P}_{N}\}_{N\geq 1}$ satisfies the LLN (Law of Large Numbers) if there exist $m_{1},m_{2},\dots\in\mathbb{R}$ such that, for all $s\in\mathbb{Z}_{\geq 1}$ and $k_{1},\dots,k_{s}\geq 1$, we have
(20) $\lim_{\begin{subarray}{c}N\to\infty,\\ N\theta\to\gamma\end{subarray}}\frac{1}{N^{s}}\,\mathbb{E}_{\mathbb{P}_{N}}\left[\prod_{j=1}^{s}{\sum_{i=1}^{N}{\mathscr{L}_{i}^{k_{j}}}}\right]=\prod_{j=1}^{s}{m_{k_{j}}},$
where, in the LHS, the $N$-tuples $(\mathscr{L}_{1}>\dots>\mathscr{L}_{N})$ are $\mathbb{P}_{N}$-distributed. [^4] We use here the probabilistic terminology and notation, as if $\mathbb{P}_{N}$ were a probability measure, but in general, for a function $f\colon\mathbb{S}(N)\to\mathbb{R}$, the expression $\mathbb{E}_{\mathbb{P}_{N}}[f(\lambda)]$ means $\sum_{\lambda\in\mathbb{Y}(N)}{f(\lambda)\mathbb{P}_{N}(\lambda)}$.
If all $\mathbb{P}_{N}$ are probability measures, then the LHS of (20) can be interpreted in terms of the empirical measures $\mu_{N}$ of $\mathbb{P}_{N}$, which by definition are the (random) probability measures on $\mathbb{R}$, given by
[…]
Indeed, the moments of the empirical measures are
$m_{k}(\mu_{N}):=\int_{\mathbb{R}}x^{k}\mu_{N}(\text{d}x)=\frac{1}{N}\sum_{i=1}^{N}{\mathscr{L}_{i}^{k}},$
then (20) is equivalent to:
$\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\mathbb{E}_{\mathbb{P}_{N}}\left[\prod_{j=1}^{s}{m_{k_{j}}(\mu_{N})}\right]=\prod_{j=1}^{s}{m_{k_{j}}}.$
As a result, LLN-satisfaction is equivalent to the convergence $\mu_{N}\to\mu$ in the sense of moments, in probability, to a probability measure $\mu$ with moments $m_{1},m_{2},\cdots$. In the case that $m_{1},m_{2},\dots$ are the moments of a unique probability measure, then LLN-satisfaction implies $\mu_{N}\to\mu$ weakly, in probability.
Our main result below shows that a sequence $\{\mathbb{P}_{N}\}_{N\geq 1}$ is HT-appropriate if and only if it satisfies the LLN. However, this result would not be complete if we did not show how the sequences $(\kappa_{n})_{n\geq 1}$ and $(m_{n})_{n\geq 1}$ in these definitions are related to each other. This relation depends on the combinatorial gadgets called Łukasiewicz paths.
Definition 3.3.
A Łukasiewicz path $\Gamma$ of length $\ell$ is a lattice path on $\mathbb{Z}^{2}$ that starts at the origin $(0,0)$, ends at $(\ell,0)$, stays above the $x$-axis, and which has steps $(1,0)$ (horizontal steps), $(1,-1)$ (down steps) and steps of the form $(1,j)$, for some $j\in\mathbb{Z}_{\geq 1}$ (up steps). We denote by $\mathbf{L}(\ell)$ the set of all Łukasiewicz paths of length $\ell$.
It is known that the cardinality of $\mathbf{L}(\ell)$ is the Catalan number $C_{\ell}=\frac{1}{\ell+1}{2\ell\choose\ell}$. For example, the five Łukasiewicz paths of length $3$ are depicted in Fig. 2. It turns out that the transformation $(\kappa_{n})_{n\geq 1}\mapsto(m_{n})_{n\geq 1}$ can be described in terms of weighted Łukasiewicz paths.
Definition 3.4.
Given a constant $\gamma\in(0,\infty)$ and $\vec{\kappa}=(\kappa_{n})_{n\geq 1}$, define $\vec{m}=(m_{n})_{n\geq 1}$ by
(22) $m_{\ell}:=\sum_{\Gamma\in\mathbf{L}(\ell)}\frac{\Delta_{\gamma}\left(x^{1+\text{\#\,horizontal steps at height $0$ in $\Gamma$}}\right)(\kappa_{1})}{1+\text{\#\,horizontal steps at height $0$ in $\Gamma$}}\cdot\prod_{i\geq 1}(\kappa_{1}+i)^{\text{\#\,horizontal steps at height $i$ in $\Gamma$}}\\ \cdot\prod_{j\geq 1}(\kappa_{j}+\kappa_{j+1})^{\text{\#\,steps $(1,j)$ in $\Gamma$}}(j+\gamma)^{\text{\#\,down steps from height $j$ in $\Gamma$}},\quad\text{for all $\ell\in\mathbb{Z}_{\geq 1}$},$
where the term $\Delta_{\gamma}$ is the divided difference operator, defined by
(23) $\Delta_{\gamma}(f)(x):=\frac{f(x)-f(x-\gamma)}{\gamma},$
and then evaluated at $x=\kappa_{1}$. We denote the map $\vec{\kappa}\mapsto\vec{m}$ by $\mathcal{J}_{\gamma}^{\kappa\mapsto m}\colon\mathbb{R}^{\infty}\to\mathbb{R}^{\infty}$ and the compositional inverse $\vec{m}\mapsto\vec{\kappa}$ by $\mathcal{J}_{\gamma}^{m\mapsto\kappa}:=(\mathcal{J}_{\gamma}^{\kappa\mapsto m})^{(-1)}$.
Note that
$\Delta_{\gamma}\left(\frac{x^{1+n}}{1+n}\right)(\gamma)=\frac{\gamma^{n}}{1+n},\quad\text{for all }n\in\mathbb{Z}_{\geq 0},$
and consequently (22) leads to the special case stated in the following lemma.
Lemma 3.5.
If $\vec{m}=\mathcal{J}^{\kappa\mapsto m}_{\gamma}(\vec{\kappa})$, in the sense of Definition 3.4, and $\kappa_{1}=\gamma$, then
$m_{\ell}=\sum_{\Gamma\in\mathbf{L}(\ell)}\frac{1}{1+\text{\#\,horizontal stepsat height $0$ in $\Gamma$}}\cdot\prod_{i\geq 0}(i+\gamma)^{\text{\#\,horizontal steps at height $i$ in $\Gamma$}}\\ \cdot\prod_{j\geq 1}(\kappa_{j}+\kappa_{j+1})^{\text{\#\,up steps $(1,j)$ in $\Gamma$}}(j+\gamma)^{\text{\#\,down steps from height $j$ in $\Gamma$}},\quad\text{for all $\ell\in\mathbb{Z}_{\geq 1}$}.$
It is not hard to see that Eq. 22 has the form
(24) $m_{\ell}=(\gamma+1)^{\uparrow(\ell-1)}\cdot\kappa_{\ell}\,+\text{ polynomial over }\mathbb{Q}[\gamma]\text{ in }\kappa_{1},\cdots,\kappa_{\ell-1},$
where
(25) $g^{\uparrow 0}:=1,\qquad g^{\uparrow n}:=\prod_{i=1}^{n}{(g+i-1)},\quad n\in\mathbb{Z}_{\geq 1},$
is the raising factorial. Therefore the system of equations (22) is invertible,
(26) $\kappa_{\ell}=\frac{1}{(\gamma+1)^{\uparrow(\ell-1)}}\cdot m_{\ell}\,+\text{ polynomial over }\mathbb{Q}(\gamma)\text{ in }m_{1},\cdots,m_{\ell-1},$
which indeed guarantees the existence of the compositional inverse map $\mathcal{J}_{\gamma}^{m\mapsto\kappa}$.
Example 5.
The first few entries of the sequence $\vec{m}=\mathcal{J}_{\gamma}^{\kappa\mapsto m}(\vec{\kappa})$ are:
$m_{1}$ $=\kappa_{1}-\frac{\gamma}{2},$
$m_{2}$ $=(\gamma+1)\kappa_{2}+\kappa_{1}^{2}+\kappa_{1}+\frac{\gamma^{2}}{3},$
$m_{3}$ $=(\gamma+1)(\gamma+2)\kappa_{3}+3(\gamma+1)\kappa_{2}\kappa_{1}+\kappa_{1}^{3}+3(\gamma+1)\kappa_{2}+\frac{3}{2}(\gamma+2)\kappa_{1}^{2}+\kappa_{1}-\frac{\gamma^{3}}{4}.$
See Fig. 2 for the calculation of $m_{3}$. From these equalities, we can deduce the first few entries of the inverse map $\vec{\kappa}=\mathcal{J}_{\gamma}^{m\mapsto\kappa}(\vec{m})$:
$\kappa_{1}$ $=m_{1}+\frac{\gamma}{2},$
$\kappa_{2}$ $=\frac{m_{2}-m_{1}^{2}-(\gamma+1)m_{1}-\frac{\gamma(7\gamma+6)}{12}}{\gamma+1},$
$\kappa_{3}$ $=\frac{m_{3}-3m_{2}m_{1}+2m_{1}^{3}-\frac{3}{2}(\gamma+2)m_{2}+\frac{3}{2}(\gamma+2)m_{1}^{2}+(\gamma+1)(\gamma+2)m_{1}+\frac{\gamma(5\gamma+4)(\gamma+2)}{8}}{(\gamma+1)(\gamma+2)}.$
Figure 2. There are five Łukasiewicz paths of length $3$. We present them with the weights associated to each of their steps, and the overall contribution after applying the divided difference operator $\Delta_{\gamma}$. The horizontal steps at height 00 and their weight after applying the divided difference operator are marked in blue.
We remark that the following computation
(27) $\Delta_{\gamma}\left(\frac{x^{1+n}}{1+n}\right)(\kappa_{1})=\frac{1}{1+n}\sum_{i=0}^{n}{\kappa_{1}^{i}(\kappa_{1}-\gamma)^{n-i}},\quad\text{for all }n\in\mathbb{Z}_{\geq 0},$
allows to alternatively express $m_{\ell}$ as a $\kappa$-weighted sum of Łukasiewicz paths of length at most $\ell$.
Theorem 3.6 (The main theorem: HT-appropriateness is equivalent to LLN-satisfaction).
Let $\{\mathbb{P}_{N}\}_{N\geq 1}$ be a sequence of finite signed measures with total mass $1$ on $N$-signatures and with small tails. Then $\{\mathbb{P}_{N}\}_{N\geq 1}$ is HT-appropriate (Definition 3.1) if and only if it satisfies the LLN (Definition 3.2). If these equivalent conditions hold, then the sequences $\vec{\kappa}=(\kappa_{n})_{n\geq 1}$ and $\vec{m}=(m_{n})_{n\geq 1}$ are related to each other by $\vec{m}=\mathcal{J}_{\gamma}^{\kappa\mapsto m}(\vec{\kappa})$, given by Definition 3.4. Furthermore, if all $\mathbb{P}_{N}$ are probability measures and moreover there exists a constant $C>0$ such that $|\kappa_{n}|\leq C^{n}$, for all $n\geq 1$, then there exists a unique probability measure $\mu^{\gamma}$ on $\mathbb{R}$ with moments $m_{\ell}$:
$m_{\ell}=\int_{\mathbb{R}}x^{\ell}\mu^{\gamma}(\text{d}x),\quad\text{for all }\ell\in\mathbb{Z}_{\geq 1}.$
In the case where all $\mathbb{P}_{N}$ are probability measures and the conditions of the previous theorem are satisfied, then the empirical measures $\mu_{N}$ of $\mathbb{P}_{N}$, defined by (21), converge to $\mu^{\gamma}$ weakly and in the sense of moments, in probability, as discussed earlier.
4. Proof of the LLN at high temperature
4.1. The action of Cherednik operators at high temperature
It is convenient to define
$\mathcal{P}_{\ell}:=\sum_{i=1}^{N}{\xi_{i}^{\ell}},\quad\ell\in\mathbb{Z}_{\geq 1}.$
Let $\{F_{N}(x_{1},\dots,x_{N})\}_{N\geq 1}$ be a sequence of smooth functions in some neighborhoods of $(1^{N})=(1,\dots,1)$ such that $F_{N}(1^{N})=0$, for all $N\geq 1$. Let $k\in\mathbb{Z}_{\geq 1}$ be arbitrary and let the $k$-th order Taylor expansion of $F_{N}$ at $(1^{N})$ be
(28) $F_{N}(x_{1},\dots,x_{N})=\sum_{\lambda\in\mathbb{Y}(N)\colon|\lambda|\leq k}{c_{F_{N}}^{\lambda}\,\mathfrak{m}_{\lambda}(x_{1}-1,\dots,x_{N}-1)}+O(\|\mathbf{x}-\mathbf{1}\|^{k+1}),$
where $\mathbf{x}-\mathbf{1}:=(x_{1}-1,\dots,x_{N}-1)$ and $\mathfrak{m}_{\lambda}$ are the monomial symmetric polynomials. Note that the coefficients corresponding to the empty partition vanish: $c_{F_{N}}^{\emptyset}=0$, for all $N\geq 1$.
The proof of Theorem 3.6 employs the next three helper theorems. The first Theorem 4.1 is a refinement of [BGCG22, Thm. 5.1], but the argument is similar, so we will keep the proof here fairly brief. In all these three theorems, we regard $\big{\{}c_{F_{N}}^{\lambda}\big{\}}_{\lambda\in\mathbb{Y}(N)}$ as a sequence of variables of degrees $\deg\big{(}c_{F_{N}}^{\lambda}\big{)}:=|\lambda|$.
Theorem 4.1.
Let $k\in\mathbb{Z}_{\geq 1}$, let $\{F_{N}=F_{N}(x_{1},\dots,x_{N})\}_{N\geq 1}$ be a sequence of smooth functions near $(1^{N})$ such that $F_{N}(1^{N})=0$, with $k$-th order Taylor expansions of $F_{N}$ given by (28), and let $\lambda=(\lambda_{1}\geq\dots\geq\lambda_{s}>0)$ be a partition with $|\lambda|=k$ and $\ell(\lambda)=s$. Then
(29) $N^{-s}\left[\prod_{j=1}^{s}{\mathcal{P}_{\lambda_{j}}}\right]e^{F_{N}}\bigg{|}_{(x_{1},\cdots,x_{N})=(1^{N})}=b^{\lambda}_{\lambda}\,c^{\lambda}_{F_{N}}+\sum_{\mu\colon|\mu|=k,\,\ell(\mu)>s}{b^{\lambda}_{\mu}\,c^{\mu}_{F_{N}}}\\ +\prod_{j=1}^{s}{\frac{1}{N}\cdot\mathcal{P}_{\lambda_{j}}e^{F_{N}}\bigg{|}_{(x_{1},\cdots,x_{N})=(1^{N})}}-\mathbf{1}_{\{s=1\}}\cdot k(\gamma+1)_{k-1}\cdot c^{(k)}_{F_{N}}\\ +R_{1}\left(c^{\nu}_{F_{N}}\colon|\nu|<k\right)+R_{2}\left(c^{\nu}_{F_{N}}\colon|\nu|\leq k\right),$
where $b^{\lambda}_{\mu}$ are uniformly bounded coefficients in the high temperature regime (17). In particular,
(30) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{b^{\lambda}_{\lambda}}=\prod_{j=1}^{s}{\lambda_{j}(\gamma+1)^{\uparrow(\lambda_{j}-1)}}.$
where we used the notation of raising factorial (25). Moreover, $R_{1}(c^{\nu}_{F_{N}}\colon|\nu|<k)$ is a polynomial in the variables $c^{\nu}_{F_{N}}$, $|\nu|<k$, of degree at most $k$ with uniformly bounded coefficients in the regime (17) and such that each monomial has at least one factor $c^{\nu}_{F_{N}}$ with $\ell(\nu)>1$. Finally, $R_{2}(c^{\nu}_{F_{N}}\colon|\nu|\leq k)$ is a polynomial in $c^{\nu}_{F_{N}}$, $|\nu|\leq k$, of degree at most $k$ with coefficients of order $o(1)$ in the regime (17).
Theorem 4.2.
Let $\ell\in\mathbb{Z}_{\geq 1}$ and let $\{F_{N}=F_{N}(x_{1},\dots,x_{N})\}_{N\geq 1}$ be a sequence of smooth functions such that $F_{N}(1^{N})=0$ with $\ell$-th order Taylor expansions at $(1^{N})$:
(31) $F_{N}(x_{1},\dots,x_{N})=\sum_{\lambda\in\mathbb{Y}(N)\colon|\lambda|\leq\ell}{c_{F_{N}}^{\lambda}\,\mathfrak{m}_{\lambda}(x_{1}-1,\dots,x_{N}-1)}+O(\|\mathbf{x}-\mathbf{1}\|^{\ell+1}).$
[…]
(32) $\frac{1}{N}\cdot\mathcal{P}_{\ell}e^{F_{N}}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=m_{\ell}\left(c^{(1)}_{F_{N}},\cdots,\ell c^{(\ell)}_{F_{N}}\right)+S_{1}\left(c^{\nu}_{F_{N}}\colon|\nu|<\ell\right)+S_{2}\left(c^{\nu}_{F_{N}}\colon|\nu|\leq\ell\right),$
where $S_{1}\big{(}c^{\nu}_{F_{N}}\colon|\nu|<\ell\big{)}$ is a polynomial in $c^{\nu}_{F_{N}}$, $|\nu|<\ell$, of degree at most $\ell$ with uniformly bounded coefficients in the regime (17), and each monomial has at least one factor $c^{\nu}_{F_{N}}$ with $\ell(\nu)>1$. Similarly, $S_{2}\big{(}c^{\nu}_{F_{N}}\colon|\nu|\leq\ell\big{)}$ is a polynomial in $c^{\nu}_{F_{N}}$, $|\nu|\leq\ell$, of degree at most $\ell$ with coefficients of order $o(1)$ in the regime (17).
An immediate corollary of the previous theorem is the following.
Corollary 4.3.
Assume that the sequence of smooth functions $F_{N}=F_{N}(x_{1},\dots,x_{N})$ in Theorem 4.2 satisfies
  1. (a)
    $\lim_{N\to\infty}\frac{\partial_{1}^{n}}{(n-1)!}F_{N}\,\Big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=:\kappa_{n}$ exists and is finite, for all $n\in\mathbb{Z}_{\geq 1}$,
  2. (b)
    $\lim_{N\to\infty}\partial_{i_{1}}\cdots\partial_{i_{r}}F_{N}\,\Big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=0$, for any $r\geq 2$ and any $i_{1},\dots,i_{r}\in\mathbb{Z}_{\geq 1}$ with at least two distinct indices among $i_{1},\dots,i_{r}$.
If we let $(m_{1},m_{2},\dots):=\mathcal{J}_{\gamma}^{\kappa\mapsto m}(\kappa_{1},\kappa_{2},\dots)$, then
(33) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\frac{1}{N}\cdot\mathcal{P}_{\ell}e^{F_{N}}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}}=m_{\ell},\quad\text{for all }\ell\in\mathbb{Z}_{\geq 1}.$
Remark 2.
The assumptions (a)-(b) in Corollary 4.3 are the same as the ones in Definition 3.1 for $F_{N}:=F_{N,\theta_{N}}$, where $N\theta_{N}\to\gamma$, as $N\to\infty$. In terms of the Taylor coefficients $c^{\nu}_{F_{N}}$ of $F_{N}$, these conditions can be expressed as:
  1. (a’)
    $\lim_{N\to\infty}{\big{(}n\cdot c^{(n)}_{F_{N}}\big{)}}=:\kappa_{n}$ exists and is finite, for all $n\in\mathbb{Z}_{\geq 1}$,
  2. (b’)
    $\lim_{N\to\infty}{c^{\nu}_{F_{N}}}=0$, for all $\nu\in\mathbb{Y}(N)$ with $\ell(\nu)>1$.
This reformulation makes the fact that Theorem 4.2 implies Corollary 4.3 more transparent.
Theorem 4.4.
Assume that $\{\mathbb{P}_{N}\}_{N\geq 1}$ is a sequence of probability measures that is HT-appropriate in the sense of Definition 3.1 and satisfies the LLN in the sense of Definition 3.2. Let $(\kappa_{n})_{n\geq 1}$ and $(m_{n})_{n\geq 1}$ be the sequences that arise from these definitions. Moreover, assume that there exists a constant $C>0$ such that $|\kappa_{n}|\leq C^{n}$, for all $n\geq 1$. Then $(m_{n})_{n\geq 1}$ is the sequence of moments of a unique probability measure on $\mathbb{R}$.
4.2. Proof of the main Theorem 3.6
Assuming the validity of Theorems 4.1, 4.2 and 4.4, we prove Theorem 3.6 here. The argument is similar to the proof of [BGCG22, Thm. 5.1]. The three helper theorems will be proved in later subsections.
Lemma 4.5.
Let $\mathbb{P}_{N}$ be any finite signed measure on $N$-signatures with small tails and with Jack generating function $G_{N,\theta}(x_{1},\dots,x_{N})$. For any $s\in\mathbb{Z}_{\geq 1}$ and $k_{1},\dots,k_{s}\geq 1$, we have
(34) $\left[\prod_{j=1}^{s}{\mathcal{P}_{k_{j}}}\right]G_{N,\theta}\,\bigg{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\mathbb{E}_{\mathbb{P}_{N}}\left[\prod_{j=1}^{s}{\sum_{i=1}^{N}{\mathscr{L}_{i}^{k_{j}}}}\right].$
Since the Cherednik operators $\xi_{1},\dots,\xi_{N}$ pairwise commute, by Theorem 2.1, then the operators $\mathcal{P}_{1},\mathcal{P}_{2},\dots$ also pairwise commute. Consequently, the order of application of the operators $\mathcal{P}_{k_{1}},\dots,\mathcal{P}_{k_{s}}$ in the LHS of (34) does not matter.
Proof.
The desired equation (34) follows from writing $G_{N,\theta}$ as the Laurent series (10), then exchanging the order between the operator $\prod_{j=1}^{s}{\mathcal{P}_{k_{j}}}$ and the infinite sum defining $G_{N,\theta}$, and finally applying Theorem 2.1, which shows that $\prod_{j=1}^{s}{\mathcal{P}_{k_{j}}}$ acts on each $P_{\lambda}(x_{1},\dots,x_{N};\theta)$ as the eigenvalue $\prod_{j=1}^{s}{\sum_{i=1}^{N}{\mathscr{L}_{i}^{k_{j}}}}$. The condition of $\mathbb{P}_{N}$ having small tails is necessary to argue that the series (10) converges and that we can exchange the order between the operators $\mathcal{P}_{k_{j}}$ and the sum. ∎
HT-appropriateness implies LLN-satisfaction. We prove that if $\{\mathbb{P}_{N}\}_{N\geq 1}$ is HT-appropriate with sequence $\vec{\kappa}=(\kappa_{n})_{n\geq 1}$, then it satisfies the LLN with $\vec{m}=(m_{n})_{n\geq 1}:=\mathcal{J}^{\kappa\mapsto m}(\vec{\kappa})$, i.e. we verify the condition from Definition 3.2, for any $k_{1},\dots,k_{s}\geq 1$.
Assume w.l.o.g. that $k_{1}\geq\dots\geq k_{s}$. By Lemma 4.5 and Theorem 4.1, $N^{-s}\cdot\mathbb{E}_{\mathbb{P}_{N}}\left[\prod_{j=1}^{s}{\sum_{i=1}^{N}{\mathscr{L}_{i}^{k_{j}}}}\right]$ is equal to the RHS of (29) with $\lambda=(k_{1},\dots,k_{s})$ and $F_{N}$ replaced by $F_{N,\theta}=\ln(G_{N,\theta})$ (we will drop the dependence of $\theta$ on $N$ in the notation for brevity, but we should keep in mind that we choose any sequence $\theta_{N}$ such that $N\theta_{N}\to\gamma$, as $N\to\infty$, as in Remark 2). It remains to prove that this converges to $\prod_{j=1}^{s}{m_{k_{j}}}$ in the regime (17). Indeed, $R_{2}\to 0$ because of the hypothesis on its coefficients, whereas $R_{1}\to 0$ because each monomial contains a factor $c^{\nu}_{F_{N,\theta}}$ with $\ell(\nu)>1$ and this converges to zero by HT-appropriateness. The terms $b^{\lambda}_{\mu}\,c^{\mu}_{F_{N,\theta}}$ in the sum have $\ell(\mu)>1$, so they converge to zero. The term $b^{\lambda}_{\lambda}\,c^{\lambda}_{F_{N,\theta}}$ converges to zero, except when $\ell(\lambda)=1$, in which case $\lambda=(k_{1})$ and so $b^{\lambda}_{\lambda}\,c^{\lambda}_{F_{N,\theta}}\to(\gamma+1)_{k_{1}-1}\cdot\kappa_{k_{1}}$ by (30) and the assumption $k_{1}\cdot c^{(k_{1})}_{F_{N,\theta}}\to\kappa_{k_{1}}$. Finally, by Corollary 4.3, the second line in the RHS of (29) converges to $\prod_{j=1}^{s}{m_{k_{j}}}-\mathbf{1}_{\{\ell(\lambda)=1\}}\cdot(\gamma+1)_{k_{1}-1}\cdot\kappa_{k_{1}}$. By putting everything together, the RHS of  (29) converges to $\prod_{j=1}^{s}{m_{k_{j}}}$, as desired.
LLN-satisfaction implies HT-appropriateness. Assume that $\{\mathbb{P}_{N}\}_{N\geq 1}$ satisfies the LLN with $\vec{m}=(m_{n})_{n\geq 1}$; let us prove HT-appropriateness with $\vec{\kappa}=(\kappa_{n})_{n\geq 1}:=\mathcal{J}_{\gamma}^{m\mapsto\kappa}(\vec{m})$. By Remark 2, we need to prove the following limits:
  1. (a’)
    $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{c^{(n)}_{F_{N,\theta}}}=\frac{\kappa_{n}}{n}$,
  2. (b’)
    $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{c^{\nu}_{F_{N,\theta}}}=0$, for all $\nu\in\mathbb{Y}(N)$ with $\ell(\nu)>1$ and $|\nu|=n$.
The proof is by induction on $n$. Let us start with the base case $n=1$. We only need to verify (a’) for $n=1$, as condition (b’) is non-existent for $n=1$. By the assumption of LLN-satisfaction and Lemma 4.5, we have that
(35) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\frac{1}{N}\,\mathcal{P}_{1}e^{F_{N,\theta}}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}}=m_{1}.$
By definition, $\mathcal{P}_{1}=\xi_{1}+\dots+\xi_{N}$, where $\xi_{i}$’s are the Cherednik operators (7). Since $F_{N,\theta}$ is symmetric in its arguments, it is killed by operators $1-s_{i,j}$, therefore $\mathcal{P}_{1}e^{F_{N,\theta}}=\sum_{i=1}^{N}(x_{i}\partial_{i}-\theta(i-1))e^{F_{N,\theta}}$. The symmetry also shows that $x_{1}\partial_{1}F_{N,\theta}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=x_{i}\partial_{i}F_{N,\theta}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}$, for all $i=1,\dots,N$. Together with the fact that $F_{N,\theta}(1^{N})=0$, we deduce
(36) $\frac{1}{N}\,\mathcal{P}_{1}e^{F_{N,\theta}}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}$ $=\frac{1}{N}\Big{(}Nx_{1}\partial_{1}-\theta\,\frac{N(N-1)}{2}\Big{)}e^{F_{N,\theta}}\Big{|}_{(x_{1},\dots,x_{N})=(1^{N})}$
$=\partial_{1}F_{N,\theta}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}-\frac{\theta(N-1)}{2}.$
From (35) and (36), we obtain
(37) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\partial_{1}F_{N,\theta}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=m_{1}+\frac{\gamma}{2}.$
Since $\kappa_{1}=m_{1}+\gamma/2$ (see Example 5), and the Taylor coefficient $c^{(1)}_{F_{N,\theta}}$ can be expressed as $c^{(1)}_{F_{N,\theta}}=\partial_{1}F_{N,\theta}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}$ (see Equation (28)), it follows that (37) is precisely (a’) for $n=1$.
For the inductive step, take any $k\geq 2$ and assume that (a’)-(b’) above are true for all $n\leq k-1$; to complete the argument, it remains to prove these conditions for $n=k$.
Take any $\lambda=(\lambda_{1}\geq\dots\geq\lambda_{s})$ with $\ell(\lambda)=s>1$, $|\lambda|=k$, and write Equation (29) for this partition. Note that the LHS of this equation converges to $\prod_{j=1}^{s}{m_{\lambda_{j}}}$, by Lemma 4.5 and LLN-satisfaction. Next, consider equations (32) for $\ell=\lambda_{1},\dots,\lambda_{s}$ and multiply them (note that the LHS of the product also converges to $\prod_{j=1}^{s}{m_{\lambda_{j}}}$) and plug the result back into the RHS of (29). The result is that the linear combination of the variables $c^{\mu}_{F_{N,\theta}}$, $|\mu|=k$, $\ell(\mu)>1$, with coefficients $b^{\lambda}_{\mu}$, up to an additive error $o(1)$, is equal to zero. Let $p(k)$ be the number of partitions of size $k$. By letting $\lambda$ range over all partitions $\lambda$ with $|\lambda|=k$ and $\ell(\lambda)>1$, we obtain a system of $p(k)-1$ equations on $p(k)$ variables $c^{\mu}_{F_{N,\theta}}$, $|\mu|=k$. We need one last equation. Take Eq. 32 with $\ell=k$ and note that the LHS converges to $m_{k}$. Moreover, since $kc^{(k)}_{F_{N,\theta}}$ only appears in $m_{k}(c^{(1)}_{F_{N,\theta}},\cdots,kc^{(k)}_{F_{N,\theta}})$ as the linear monomial $(\gamma+1)_{k-1}\cdot kc^{(k)}_{F_{N,\theta}}$, according to (24), then the term $m_{k}(c^{(1)}_{F_{N,\theta}},\cdots,kc^{(k)}_{F_{N,\theta}})$ in the RHS of (32) is asymptotically equal to $m_{k}-(\gamma+1)_{k-1}\cdot kc^{(k)}_{F_{N,\theta}}+(\gamma+1)_{k-1}\cdot\kappa_{k}$. Thus, upon rearranging, this equation is (up to an additive $o(1)$ error) of the form $\frac{\kappa_{k}}{k}=c^{(k)}_{F_{N,\theta}}+$ linear combination of $c^{\nu}_{F_{N,\theta}}$, $|\nu|=k$, with coefficients of order $o(1)$.
The previous paragraph shows how to construct a system of $p(k)$ equations in the $p(k)$ variables $c^{\mu}_{F_{N,\theta}}$, $|\mu|=k$, for which the matrix of coefficients consists of the $b^{\lambda}_{\mu}$, up to an additive error $o(1)$. But $b^{\lambda}_{\mu}\neq o(1)$ only when $\ell(\mu)>\ell(\lambda)$ or $\mu=\lambda$; also, the diagonal elements $b^{\lambda}_{\lambda}$ are asymptotically nonzero, due to (30), so if we label the $p(k)$ equations by their associated partitions $\lambda$, and associated column vectors $(\lambda_{1}^{\prime},\lambda_{2}^{\prime},\dots)$, and order them w.r.t. lexicographic order of vectors, it follows that the system of equations is asymptotically lower-triangular with nonzero diagonal entries. In particular, the solution has a limit which must be given exactly by equations (a’)-(b’) for $n=k$. This completes the proof.
The moment problem. The fact that if $|\kappa_{n}|\leq C^{n}$, for some constant $C>0$, then $m_{1},m_{2},\dots$ are the moments of a unique probability measure on $\mathbb{R}$ is exactly the content of Theorem 4.4. This ends the proof of Theorem 3.6.
4.3. Method of moments: Proof of Theorem 4.1
Consider the $k$-th order truncations
(38) $\tilde{F}_{N}(x_{1},\dots,x_{N}):=\sum_{\lambda\colon|\lambda|\leq k,\,\ell(\lambda)\leq N}c_{F_{N}}^{\lambda}\,\mathfrak{m}_{\lambda}(x_{1}-1,\dots,x_{N}-1).$
The operators $x_{i}\partial_{i}$ and $\frac{x_{\ell}}{x_{i}-x_{j}}(1-s_{i,j})$ for some $i,j,\ell$, are constituents of the Cherednik operators (7). Their actions on shifted monomials, that is, monomials in the variables $x_{1}-1,\dots,x_{N}-1$, are the following:
(39) $x_{i}\partial_{i}\left[(x_{1}-1)^{\alpha_{1}}\cdots(x_{N}-1)^{\alpha_{N}}\right]=\alpha_{i}\bigg{(}(x_{1}-1)^{\alpha_{1}}\cdots(x_{N}-1)^{\alpha_{N}}+\\ \hfill+(x_{1}-1)^{\alpha_{1}}\cdots(x_{i}-1)^{\alpha_{i}-1}\cdots(x_{N}-1)^{\alpha_{N}}\bigg{)},\\ \frac{x_{\ell}}{x_{i}-x_{j}}(1-s_{i,j})\left[(x_{1}-1)^{\alpha_{1}}\cdots(x_{N}-1)^{\alpha_{N}}\right]=\pm x_{\ell}(x_{1}-1)^{\alpha_{1}}\cdots(x_{i}-1)^{a}\cdots(x_{j}-1)^{a}\cdots(x_{N}-1)^{\alpha_{N}}\\ \times\sum_{c=0}^{b-a-1}{(x_{i}-1)^{c}(x_{j}-1)^{b-a-1-c}},$
where $a:=\min(\alpha_{i},\alpha_{j})$, $b:=\max(\alpha_{i},\alpha_{j})$, and the sign is positive (negative, respectively) if $a=\alpha_{j}$ ($a=\alpha_{i}$, respectively). This leads to the following lemma.
Lemma 4.6.
Let $1\leq i_{1},\dots,i_{k}\leq N$ be arbitrary integers, not necessarily distinct. Then
(40) $\left(\prod_{j=1}^{k}\xi_{i_{j}}\right)\exp(F_{N})\Big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\left(\prod_{j=1}^{k}\xi_{i_{j}}\right)\exp(\tilde{F}_{N})\Big{|}_{(x_{1},\dots,x_{N})=(1^{N})}.$
Moreover, this is a polynomial in the variables $c_{F_{N}}^{\lambda}$, $|\lambda|\leq k$, of degree at most $k$, and with coefficients that are uniformly bounded over all $i_{1},\dots,i_{k}$ in the regime (17).
Proof.
Note that
$\exp(F_{N}(x_{1},\dots,x_{N}))=\exp(\tilde{F}_{N}(x_{1},\dots,x_{N}))+R(x_{1},\dots,x_{N}),$
where $R(x_{1},\dots,x_{N})=O(\|\mathbf{x}-\mathbf{1}\|^{k+1})$. This means that $R(x_{1},\dots,x_{N})$ vanishes up to degree $(k+1)$ at $(1^{N})$. After applying any sequence of $k$ operators $\theta(i-1)$, $x_{i}\partial_{i}$ or $\frac{x_{\ell}}{x_{i}-x_{j}}(1-s_{i,j})$, we obtain a continuously differentiable function that still vanishes at $(1^{N})$. Since $\prod_{j=1}^{k}\xi_{i_{j}}$ is a linear combination of sequences of $k$ operators as just described, we obtain the desired (40).
We now prove the polynomiality part. Consider any product of $k$ operators, each of the form $\theta(i-1)$, $x_{i}\partial_{i}$ or $\frac{x_{\ell}}{x_{i}-x_{j}}(1-s_{i,j})$, for some $i,j,\ell$. We want to apply them inductively to $\exp(\tilde{F}_{N})$ using the following rules:
(41) $\begin{gathered}x_{i}\partial_{i}\left(H(\mathbf{x})\cdot\exp(\tilde{F}_{N}(\mathbf{x}))\right)=\left(x_{i}\big{(}\partial_{i}H(\mathbf{x})+H(\mathbf{x})\cdot\partial_{i}\tilde{F}_{N}(\mathbf{x})\big{)}\right)\cdot\exp(\tilde{F}_{N}(\mathbf{x})),\\ \frac{x_{\ell}}{x_{i}-x_{j}}(1-s_{i,j})\left(H(\mathbf{x})\cdot\exp(\tilde{F}_{N}(\mathbf{x}))\right)=\left(\frac{x_{\ell}}{x_{i}-x_{j}}(1-s_{i,j})H(\mathbf{x})\right)\cdot\exp(\tilde{F}_{N}(\mathbf{x})).\end{gathered}$
Taking into account that $\tilde{F}_{N}(1^{N})=0$, the result of having $k$ such operators act on $\exp(\tilde{F}_{N}(\mathbf{x}))$ and then setting all variables equal to $1$ is a finite linear combination of products of actions of $\theta(i-1)$, $x_{i}\partial_{i}$ and $\frac{x_{\ell}}{x_{i}-x_{j}}(1-s_{i,j})$ on $\tilde{F}_{N}(\mathbf{x})$, followed by evaluation at $\mathbf{x}=(1^{N})$. Since $\tilde{F}_{N}(\mathbf{x})$ is a polynomial of the shifted variables $x_{1}-1,\dots,x_{N}-1$ with coefficients $c^{\lambda}_{F_{N}}$, and the actions of $x_{i}\partial_{i}$, $\frac{x_{\ell}}{x_{i}-x_{j}}(1-s_{i,j})$ on shifted monomials are given by (39), we conclude the polynomiality. It remains to verify the statements about the degree and uniform boundedness.
The previous argument shows in fact that the rules (39) and (41) imply that the action of $\xi_{i_{q}}\cdots\xi_{i_{1}}$ on $\exp(\tilde{F}_{N})$ gives an expression of the form
$\xi_{i_{q}}\cdots\xi_{i_{1}}\exp(\tilde{F}_{N})=\left(\sum_{r_{1},\dots,r_{N}\geq 0}{h^{(q)}_{r_{1},\dots,r_{N}}(x_{1}-1)^{r_{1}}\cdots(x_{N}-1)^{r_{N}}}\right)\cdot\exp(\tilde{F}_{N}),$
for all $q=1,\dots,k$, and some coefficients $h^{(q)}_{r_{1},\dots,r_{N}}$. At the end, since $\tilde{F}_{N}(1^{N})=0$, we have
$\xi_{i_{k}}\cdots\xi_{i_{1}}\exp(\tilde{F}_{N})\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=h^{(k)}_{0,\dots,0}.$
We claim that each $h^{(q)}_{r_{1},\dots,r_{N}}$, with $1\leq q\leq k$ and $r_{1},\dots,r_{N}\geq 0$, is a polynomial in the $c^{\lambda}_{F_{N}}$, $|\lambda|\leq k$, of degree at most $r_{1}+\dots+r_{k}+q$, and with coefficients that are uniformly bounded in the regime (17). In particular, this will be true of $h^{(k)}_{0,\dots,0}$ and will end the proof.
The previous claim follows from the explicit expression for the Cherednik operators (7). Indeed, the action of Cherednik operators on a homogeneous polynomial of degree $r$ in the shifted variables $x_{1}-1,\dots,x_{N}-1$ can be expressed as an inhomogeneous polynomial in these shifted variables, with nonzero components only in degrees $r-1$ and $r$; this is a consequence of the calculations (39). This observation and a simple induction on $q$ proves the claim about the degree of $h^{(q)}_{r_{1},\dots,r_{N}}$.
For the claim about uniform boundedness, note that the rules (41) allow one to express each $h^{(q+1)}_{r_{1},\dots,r_{N}}$ as a linear combination of $h^{(q)}_{s_{1},\dots,s_{N}}$, over $s_{1},\dots,s_{N}\geq 0$. In this linear combination, the number of coefficients coming from the action of constituents $x_{i_{q+1}}\partial_{i_{q+1}}$ and $-\theta(i_{q+1}-1)$ is $O(1)$, whereas the number of coefficients coming from the action of constituents of the form $\theta\frac{x_{\ell}}{x_{i}-x_{j}}(1-s_{i,j})$, with $i$ or $j$ being $i_{q+1}$, is $O(N)$. Since $|-\theta(i_{q+1}-1)|\leq\theta N$ and $\theta N\sim\gamma$, it follows by induction that for $q=1,\dots,k$, the coefficients of $h^{(q)}_{s_{1},\dots,s_{N}}$, as a polynomial on $c^{\lambda}_{F_{N}}$, can be bounded by a constant that only depends on $\gamma$, as desired. ∎
By virtue of Lemma 4.6 we can, and will, assume without loss of generality that $F_{N}(x_{1},\dots,x_{N})$ is the polynomial given by the RHS of (38):
(42) $F_{N}(x_{1},\dots,x_{N})=\sum_{\lambda\colon|\lambda|\leq k,\,\ell(\lambda)\leq N}c_{F_{N}}^{\lambda}\,\mathfrak{m}_{\lambda}(x_{1}-1,\dots,x_{N}-1).$
Lemma 4.7.
Let $s\in\mathbb{Z}_{\geq 1}$ be arbitrary and let $\alpha_{1},\dots,\alpha_{s}\in\mathbb{Z}_{\geq 1}$ be such that $\alpha_{1}+\cdots+\alpha_{s}\leq k$. Further, let $1\leq i_{1},\dots,i_{s}\leq N$ be any distinct integers. Then
$\left(\prod_{j=1}^{s}\xi_{i_{j}}^{\alpha_{j}}\right)\exp(F_{N})\bigg{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\prod_{j=1}^{s}\left(\xi_{i_{j}}^{\alpha_{j}}\exp(F_{N})\right)\Big{|}_{(x_{1},\dots,x_{N})=(1^{N})}+R+O\left(\frac{1}{N}\right),$
where $R$ is a polynomial in the $c_{F_{N}}^{\lambda}$ such that $\deg(R)\leq k$, it has coefficients of order $O(1)$ in the regime (17), uniformly over all $1\leq i_{1},\dots,i_{s}\leq N$, and such that each monomial in it contains at least a factor $c_{F_{N}}^{\nu}$ with $\ell(\nu)>1$. Moreover, $O(\frac{1}{N})$ is a polynomial in the $c_{F_{N}}^{\lambda}$ of degree $\leq k$, whose coefficients are $O(\frac{1}{N})$ in the regime (17), uniformly over all $1\leq i_{1},\dots,i_{s}\leq N$.
Proof.
Since the coefficients $c_{F_{N}}^{\nu}$ with $\ell(\nu)>1$ will be swallowed by $R$, and the uniformity of bounds on the coefficients of $R$ is already proved in Lemma 4.6, for this lemma we can replace $F_{N}(\mathbf{x})$ in (42) by
(43) $H_{N}(\mathbf{x}):=\sum_{n=0}^{k}{c_{F_{N}}^{(n)}\mathfrak{m}_{(n)}(x_{1}-1,\dots,x_{N}-1)}=\sum_{l=1}^{N}\sum_{n=0}^{k}{c_{F_{N}}^{(n)}(x_{l}-1)^{n}}.$
Then the desired statement reduces to proving:
(44) $\xi_{i_{s}}^{\alpha_{s}}\cdots\xi_{i_{1}}^{\alpha_{1}}\exp(H_{N})\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}\\ =\prod_{j=1}^{s}\left[\xi_{i_{j}}^{\alpha_{j}}\exp(H_{N})\right]\bigg{|}_{(x_{1},\dots,x_{N})=(1^{N})}+O\left(\frac{1}{N}\right).$
Note that $\xi_{i_{s}}^{\alpha_{s}}\cdots\xi_{i_{1}}^{\alpha_{1}}$ will be a sum of $(N+1)^{\sum_{j=1}^{s}{\alpha_{j}}}$ terms $w_{1}^{s}\cdots w_{\alpha_{s}}^{s}\cdots w_{1}^{1}\cdots w_{\alpha_{1}}^{1}$, where each $w_{b}^{a}$ is of the form $x_{i_{a}}\partial_{i_{a}}$ or $-\theta(i_{a}-1)$ or $\frac{\theta x_{\ell}}{x_{i_{a}}-x_{j}}(1-s_{i_{a},j})$, for some $1\leq j<i_{a}$ and $\ell=i_{a}$, or $i_{a}<j\leq N$ and $\ell=j$. If a term $w_{1}^{s}\cdots w_{\alpha_{s}}^{s}\cdots w_{1}^{1}\cdots w_{\alpha_{1}}^{1}$ satisfies:
$\bullet$ some $w^{a}_{b}=\frac{\theta x_{\ell}}{x_{i}-x_{j}}(1-s_{i,j})$ and $j\in\{i_{1},\dots,i_{s}\}$, or
$\bullet$ some $w^{a}_{b}=\frac{\theta x_{\ell}}{x_{i}-x_{j}}(1-s_{i,j}),\ w^{a^{\prime}}_{b^{\prime}}=\frac{\theta x_{\ell^{\prime}}}{x_{i^{\prime}}-x_{j^{\prime}}}(1-s_{i^{\prime},j^{\prime}})$, such that $a\neq a^{\prime}$, and $j=j^{\prime}$,
then we call it singular and if it does not satisfy any of the two conditions above, we call it generic. There are at least $(N-s)(N-s-1)\cdots(N-s-\sum_{j=1}^{s}{\alpha_{j}}+1)$ generic terms, where each $w^{a}_{b}$ is of the form $\frac{\theta x_{\ell}}{x_{i}-x_{j}}(1-s_{i,j})$, as can be seen by picking the letters $w_{1}^{s},w_{2}^{s},\cdots,w^{1}_{\alpha_{1}}$ in that order, preventing at each step both conditions for singularity. Then the number of singular terms is at most $(N+1)^{\sum_{j=1}^{s}{\alpha_{j}}}-(N-s)(N-s-1)\cdots(N-s-\sum_{j=1}^{s}{\alpha_{j}}+1)=O(N^{\sum_{j=1}^{s}{\alpha_{j}}-1})$, which is smaller by a factor of $N$ than the number of generic terms. Thus, to analyze the LHS of (44), it suffices to look at the generic terms $w_{1}^{s}\cdots w_{\alpha_{s}}^{s}\cdots w_{1}^{1}\cdots w_{\alpha_{1}}^{1}\exp(H_{N})\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}$, since the singular ones will be part of $O(\frac{1}{N})$ in (44).
By the rules (39), (41), we deduce
(45) $w_{1}^{1}\cdots w_{\alpha_{1}}^{1}\exp(H_{N})=h_{1}\cdot\exp(H_{N}),$
where $h_{1}=h_{1}(\mathbf{x})$ is a polynomial that involves only variables $x_{i_{1}}$ and $x_{j}$, for those $j$ that are part of some operator $w^{1}_{a}=\frac{\theta x_{\ell}}{x_{i_{1}}-x_{j}}(1-s_{i_{1},j})$; this part of the argument is where the special form of $H_{N}(\mathbf{x})$ in (43) is used. Then the next step is to apply $w_{1}^{2}\cdots w_{\alpha_{2}}^{2}$ to (45). By the condition of being generic and the rules (41), it turns out that these $w^{2}_{a}$ do not act on $h_{1}$, but only on $\exp(H_{N})$, thus we conclude that $w_{1}^{2}\cdots w_{\alpha_{2}}^{2}w_{1}^{1}\cdots w_{\alpha_{1}}^{1}\exp(H_{N})=w_{1}^{2}\cdots w_{\alpha_{2}}^{2}\exp(H_{N})\times w_{1}^{1}\cdots w_{\alpha_{1}}^{1}\exp(H_{N})$. By repeating this argument $s$ times, we deduce that for generic terms, we have the exact equality
$w_{1}^{s}\cdots w_{\alpha_{s}}^{s}\cdots w_{1}^{1}\cdots w_{\alpha_{1}}^{1}\exp(H_{N})=\prod_{j=1}^{s}{w_{1}^{j}\cdots w_{\alpha_{j}}^{j}\exp(H_{N})}.$
After setting all variables $x_{1},\dots,x_{N}$ equal to $1$, then adding all these generic terms and noting that the special terms amount to an additive uniform error of $O(\frac{1}{N})$, as discussed earlier, we conclude the desired (44). ∎
Lemma 4.8.
Let $s\in\mathbb{Z}_{\geq 1}$ be arbitrary and let $1\leq i_{1},\dots,i_{s}\leq N$ be distinct integers. Further, let $\lambda=(\alpha_{1}\geq\dots\geq\alpha_{s}>0)$ be a partition with $\ell(\lambda)=s$ and $|\lambda|=\alpha_{1}+\cdots+\alpha_{s}=k$. Then
(46) $\left(\prod_{j=1}^{s}\xi_{i_{j}}^{\alpha_{j}}\right)\exp(F_{N})\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=b^{\lambda}_{\lambda}c^{\lambda}_{F_{N}}+\sum_{\mu\colon|\mu|=k,\,\ell(\mu)>s}{b^{\lambda}_{\mu}c^{\mu}_{F_{N}}}+R+O\left(\frac{1}{N}\right),$
where the coefficients $b^{\lambda}_{\mu}$ are uniformly bounded in the regime (17); in particular,
(47) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{b^{\lambda}_{\lambda}}=\prod_{j=1}^{s}{\alpha_{j}(\gamma+1)^{\uparrow(\alpha_{j}-1)}}.$
Moreover, $R$ is a polynomial in the $c^{\lambda}_{F_{N}}$, $|\lambda|<k$ (i.e. it does not involve $c^{\nu}_{F_{N}}$ with $|\nu|=k$) with $\deg(R)\leq k$, and $O(\frac{1}{N})$ is a polynomial in the $c^{\lambda}_{F_{N}}$ of degree $\leq k$, whose coefficients are $O(\frac{1}{N})$ in the regime (17).
Proof.
According to Lemma 4.6, the LHS of (46) is a polynomial of degree at most $k$ in the coefficients $c^{\lambda}_{F_{N}}$, so its linear degree $k$ component is of the form
(48) $\sum_{\mu\colon|\mu|=k}{b^{\lambda}_{\mu}c^{\mu}_{F_{N}}}.$
It remains to prove that $b^{\lambda}_{\mu}=O(\frac{1}{N})$, unless $\ell(\mu)>s$ or $\mu=\lambda$, as well as the limit (47).
Consider the polynomial
(49) $\hat{F}_{N}(x_{1},\dots,x_{N}):=\sum_{\lambda\colon|\lambda|\leq k,\,\ell(\lambda)\leq N}{c_{F_{N}}^{\lambda}\,\mathfrak{m}_{\lambda}(x_{1},\dots,x_{N})},$
which is related to $F_{N}$ by $F_{N}(x_{1},\dots,x_{N})=\hat{F}_{N}(x_{1}-1,\dots,x_{N}-1)$. We also consider the following Dunkl operators [Dun89],
(50) $\tilde{\xi}_{i}:=\partial_{i}+\theta\sum_{j\neq i}\frac{1}{x_{i}-x_{j}}(1-s_{i,j}),\quad i=1,2,\dots,N.$
Note that the Dunkl operators decrease the degree of polynomials by $1$, whereas the Cherednik operators (7) preserve the degree of polynomials. Moreover, by looking at the rules (41) of how the Cherednik operators act, and at the analogous relations for Dunkl operators [BGCG22, Eqns. (5.8)-(5.9)], we deduce
$\left(\prod_{j=1}^{s}\xi_{i_{j}}^{\alpha_{j}}\right)\exp(F_{N})=\left(\prod_{j=1}^{s}(\xi_{i_{j}}+\tilde{\xi}_{i_{j}})^{\alpha_{j}}\right)\exp(\hat{F}_{N})\bigg{|}_{x_{i}\mapsto x_{i}-1,\ \forall i=1,\dots,N}$
and therefore
(51) $\left(\prod_{j=1}^{s}\xi_{i_{j}}^{\alpha_{j}}\right)\exp(F_{N})\bigg{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\left(\prod_{j=1}^{s}(\xi_{i_{j}}+\tilde{\xi}_{i_{j}})^{\alpha_{j}}\right)\exp(\hat{F}_{N})\bigg{|}_{(x_{1},\dots,x_{N})=(0^{N})}.$
To calculate the RHS of (51), write $\prod_{j=1}^{s}(\xi_{i_{j}}+\tilde{\xi}_{i_{j}})^{\alpha_{j}}$ as a sum of $2^{k}$ sequences of $k$ operators, each being either $\xi_{i_{j}}$ or $\tilde{\xi}_{i_{j}}$, then apply each sequence of operators to $\exp(\hat{F}_{N})$ in order to obtain an expression of the form $h\cdot\exp(\hat{F}_{N})$, and finally pick up the constant term of $h=h(\mathbf{x})$. Since we are only interested in the degree $k$ linear component (48) and the Cherednik operators preserve the degree of polynomials, any sequence involving some Cherednik operator $\xi_{i_{j}}$ will not contribute to (48). As a result, the part of the RHS of (51) that contributes to (48) comes from
(52) $\left(\prod_{j=1}^{s}\tilde{\xi}_{i_{j}}^{\,\alpha_{j}}\right)\exp(\hat{F}_{N})\bigg{|}_{(x_{1},\dots,x_{N})=(0^{N})}.$
It was precisely proved in [BGCG22, Prop. 5.5] that the coefficients of the linear component $\sum_{\mu\colon|\mu|=k}{b^{\lambda}_{\mu}c^{\mu}_{F_{N}}}$ of (52) are of order $b^{\lambda}_{\mu}=O(\frac{1}{N})$, unless $\ell(\mu)>\ell(\lambda)$ or $\mu=\lambda$; the limit (47) was also shown there. Hence, the proof is now finished. ∎
Proof of Theorem 4.1.
The theorem is an easy consequence of Lemmas 4.6, 4.7 and 4.8; since the argument is very similar to the proof of [BGCG22, Thm. 5.1], we omit the details. Note however that [BGCG22, Thm. 5.1] has a polynomial $R_{2}$ with coefficients of order $O(\frac{1}{N})$, while in our theorem the polynomial $R_{2}$ has coefficients of order $o(1)$. [^5] It seems that [BGCG22, Thm. 5.1] has a mistake, namely, the difference between the two sides of [BGCG22, Eqn. (5.3)] is not of order $O(\frac{1}{N})$, but of order $o(1)$. However, the applications of [BGCG22, Thm. 5.1] only use that the contribution of $R_{2}$ vanishes in the regime (17), so their mistake does not affect their other results.
4.4. Limiting moments via Łukasiewicz paths: Proof of Theorem 4.2
Throughout this subsection, assume that we are in the setting of Theorem 4.2. For the proof, it is convenient to introduce some notation. Let :=$\mathcal{F}:=\{F_{N}=F_{N}(x_{1},\dots,x_{N})\}_{N\geq 1}$ be the sequence of smooth functions with Taylor expansions (31) and such that $F_{N}(1^{N})=0$. The Taylor coefficients $c^{\nu}_{F_{N}}$, $\nu\in\mathbb{Y}$, will be treated as variables of degrees $\deg(c^{\nu}_{F_{N}})=|\nu|$, as before.
For any $\ell\in\mathbb{Z}_{\geq 1}$, define $I_{1}(\mathcal{F},\ell)$ as the set of polynomials $R_{1}=R_{1}\big{(}c^{\nu}_{F_{N}}\colon\nu\in\mathbb{Y}\big{)}$ in the variables $c^{\nu}_{F_{N}}$ with real coefficients, possibly depending on $N$ and $\theta$, and such that
  • $\deg(R_{1})\leq\ell$;
  • $R_{1}$ has uniformly bounded coefficients in the regime (17);
  • each monomial of $R_{1}$ has at least one factor $c^{\nu}_{F_{N}}$ with $\ell(\nu)>1$.
Similarly, define $I_{2}(\mathcal{F},\ell)$ as the set of polynomials $R_{2}=R_{2}\big{(}c^{\nu}_{F_{N}}\colon\nu\in\mathbb{Y}\big{)}$ such that
  • $\deg(R_{2})\leq\ell$;
  • $R_{2}$ has coefficients of order $o(1)$ in the regime (17).
Finally we will need the following formal power series, for all $i=1,2,\dots,N$:
(53) $\Xi_{i}(z):=1+\sum_{\ell\geq 1}\xi_{i}^{\ell}\exp(F_{N})\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}\cdot z^{\ell}.$
Lemma 4.9.
Assume that we are in the setting of Theorem 4.2, in particular, $\ell\in\mathbb{Z}_{\geq 1}$ and :=$\mathcal{F}:=\{F_{N}=F_{N}(x_{1},\dots,x_{N})\}_{N\geq 1}$ is a sequence of smooth functions with Taylor expansions (31), such that $F_{N}(1^{N})=0$. Also, let $i=1,2,\dots,N$. Then
(54) $\xi_{i}^{\ell}\exp(F_{N})\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=[z^{\ell}]\Xi_{1}(z)\sum_{k\geq 0}\big{(}(1-i)\theta z\cdot\Xi_{1}(z)\big{)}^{k}+R_{1}+R_{2},$
for some $R_{1}\in I_{1}(\mathcal{F},\ell)$ and $R_{2}\in I_{2}(\mathcal{F},\ell)$. Moreover, the bounds on the coefficients of $R_{1},R_{2}$ are uniform over $i=1,2,\dots,N$.
Proof.
The statement is obvious when $i=1$, so assume that $2\leq i\leq N$ and let ev denote the operator of evaluation of a function in variables $x_{1},\dots,x_{N}$ at $(1^{N})$. Since $\text{ev}=\text{ev}\circ s_{i-1}$, therefore by using the Hecke relation (8) repeatedly, we get that the LHS of (54) is equal to
$\text{ev}\xi_{i}^{\ell}\exp(F_{N})=\text{ev}s_{i-1}\xi_{i}^{\ell}\exp(F_{N})=\text{ev}(\xi_{i-1}s_{i-1}-\theta)\xi_{i}^{\ell-1}\exp(F_{N})=\\ \text{ev}\bigg{(}\xi_{i-1}(\xi_{i-1}s_{i-1}-\theta)\xi_{i}^{\ell-2}-\theta\xi_{i}^{\ell-1}\bigg{)}\exp(F_{N})=\cdots=\text{ev}\xi_{i-1}^{\ell}\exp(F_{N})-\theta\sum_{\begin{subarray}{c}a,b\geq 0\\ a+b=\ell-1\end{subarray}}\text{ev}\xi_{i-1}^{a}\xi_{i}^{b}\exp(F_{N}).$
Applying Lemma 4.7 to each summand $\text{ev}\xi_{i-1}^{a}\xi_{i}^{b}\exp(F_{N})$ with $a,b\geq 1$, we have
(55) $\text{ev}\xi_{i-1}^{a}\xi_{i}^{b}\exp(F_{N})=\text{ev}\xi_{i-1}^{a}\exp(F_{N})\cdot\text{ev}\xi_{i}^{b}\exp(F_{N})-R_{i;a,b}-\frac{1}{N}R^{\prime}_{i;a,b},$
for $R_{i;a,b}$, $R^{\prime}_{i;a,b}$ as described in that lemma. If $a=0$ or $b=0$ (or both), then (55) is also true if we set both $R_{i;a,b}$ and $R^{\prime}_{i;a,b}$ be the zero polynomials. This leads to the power series identity
(56) $\Xi_{i}(z)=\left(\Xi_{i-1}(z)+\theta zR_{i}(z)\right)\left(1+\theta z\cdot\Xi_{i-1}(z)\right)^{-1},$
where
(57) $R_{i}(z):=\sum_{a,b\geq 1}\left(R_{i;a,b}+\frac{1}{N}R^{\prime}_{i;a,b}\right)z^{a+b}.$
Taking into account that (56) holds true for any $2\leq i\leq N$, we obtain by iterating this equation that for any $11\leq k\leq i-$ 1, one has
(58) $\Xi_{i}(z)=\left(\Xi_{k}(z)+\sum_{m=1}^{i-k}(\theta z)^{m}f^{(k)}_{m}\right)\left(1+(i-k)\theta z\cdot\Xi_{k}(z)+\sum_{m=2}^{i-k}(\theta z)^{m}g^{(k)}_{m}\right)^{-1},$
where
$f^{(k)}_{2j}$ $=\Xi_{k}(z)\cdot\sum_{k+1\leq i_{1}<\cdots<i_{j}\leq i}(i_{1}-k-1)\prod_{s=1}^{j-1}(i_{s+1}-i_{s}-1)\cdot R_{i_{1}}(z)\cdots R_{i_{j}}(z),$
$f^{(k)}_{2j-1}$ $=\sum_{k+1\leq i_{1}<\cdots<i_{j}\leq i}\,\prod_{s=1}^{j-1}(i_{s+1}-i_{s}-1)\cdot R_{i_{1}}(z)\cdots R_{i_{j}}(z),$
$g^{(k)}_{2j}$ $=\sum_{k+1\leq i_{1}<\cdots<i_{j}\leq i}(i-i_{j})\prod_{s=1}^{j-1}(i_{s+1}-i_{s}-1)\cdot R_{i_{1}}(z)\cdots R_{i_{j}}(z),$
$g^{(k)}_{2j+1}$ $=\Xi_{k}(z)\cdot\sum_{k+1\leq i_{1}<\cdots<i_{j}\leq i}(i-i_{j})(i_{1}-k-1)\prod_{s=1}^{j-1}(i_{s+1}-i_{s}-1)\cdot R_{i_{1}}(z)\cdots R_{i_{j}}(z).$
By taking $k=1$ in (58) we obtain that
(59) $\Xi_{i}(z)=\left(\Xi_{1}(z)+z\tilde{R}(z)\right)\left(1+(i-1)\theta z\cdot\Xi_{1}(z)+z\tilde{R}^{\prime}(z)\right)^{-1},$
where
(60) $\tilde{R}(z):=\theta\sum_{m=1}^{i-1}(\theta z)^{m-1}f^{(1)}_{m},\quad\tilde{R}^{\prime}(z):=\theta\sum_{m=2}^{i-1}(\theta z)^{m-1}g^{(1)}_{m}.$
Claim 4.10.
The constant coefficients $[z^{0}]\tilde{R}(z)$, $[z^{0}]\tilde{R}^{\prime}(z)$ are equal to zero. Moreover, for any $p\in\mathbb{Z}_{\geq 1}$, both $[z^{p}]\tilde{R}(z)$ and $[z^{p}]\tilde{R}^{\prime}(z)$ are of the form $S_{1}+S_{2}$, where $S_{1}\in I_{1}(\mathcal{F},p)$ and $S_{2}\in I_{2}(\mathcal{F},p)$.
Let us defer the proof of the claim until later. The desired conclusion is now a consequence of the following chain of equalities
$\xi_{i}^{\ell}\exp(F_{N})\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}$ $=[z^{\ell}]\Xi_{i}(z)$
(61) $=[z^{\ell}]\left(\Xi_{1}(z)+z\tilde{R}(z)\right)\sum_{k\geq 0}\left((1-i)\theta z\cdot\Xi_{1}(z)-z\tilde{R}^{\prime}(z)\right)^{k}$
(62) $=[z^{\ell}]\,\Xi_{1}(z)\sum_{k\geq 0}\big{(}(1-i)\theta z\cdot\Xi_{1}(z)\big{)}^{k}+R_{1}+R_{2},$
for some $R_{1}\in I_{1}(\mathcal{F},p)$ and $R_{2}\in I_{2}(\mathcal{F},p)$. Indeed, the first equality above is due to the definition (53) of $\Xi_{i}(z)$, whereas the second one is due to Equation (59), so only the last equality remains to be justified. For this, let the difference between the coefficient of $z^{\ell}$ in line (61) and the coefficient of $z^{\ell}$ in line (62) be denoted by $[z^{\ell}]A(z)$; the task is to verify that $[z^{\ell}]A(z)$ has the form $R_{1}+R_{2}$, for some $R_{1}\in I_{1}(\mathcal{F},p)$ and $R_{2}\in I_{2}(\mathcal{F},p)$. Evidently, $A(z)$ is a finite sum of products, each of which possibly contains the factors $\Xi_{1}(z)$ and $\big{(}(1-i)\theta\big{)}^{j}\cdot\big{(}z\cdot\Xi_{1}(z)\big{)}^{j}$, for some $1\leq j\leq\ell$, and it always contains a factor of the form $z\tilde{R}(z)$ or $\big{(}z\tilde{R}^{\prime}(z)\big{)}^{j}$, for some $1\leq j\leq\ell$ (or both of them). By Lemma 4.6, applied to $\Xi_{1}(z)=\xi_{1}^{\ell}\exp(F_{N})\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}$, and 4.10, it follows that $[z^{\ell}]A(z)$ is a polynomial of the $c^{\lambda}_{F_{N}}$ of degree at most $\ell$. Then we can gather those monomials that have some factor $c^{\nu}_{F_{N}}$ with $\ell(\nu)>1$ into $R_{1}$ and those that do not contain such factors into $R_{2}$. The fact that the coefficients of $R_{1}$ and $R_{2}$ are of orders $O(1)$ and $o(1)$, respectively, in the regime (17) also follow from Lemma 4.6 and 4.10; finally, the uniformity claim stated in the lemma follows because the factors $\big{(}(1-i)\theta\big{)}^{j}$, for some $1\leq j\leq\ell$, that could appear are bounded by
$\big{|}((1-i)\theta)^{j}\big{|}\leq(N\theta)^{j}\sim\gamma^{j}\leq\max\{1,\gamma^{\ell}\}$
and this is independent of $i=1,2,\dots,N$. It only remains to show 4.10.
Proof of 4.10. From (60), we have
$[z^{p}]\tilde{R}(z)=\sum_{m=1}^{p+1}\theta^{m}[z^{p+1-m}]f_{m}^{(1)},\qquad[z^{p}]\tilde{R}^{\prime}(z)=\sum_{m=2}^{p+1}\theta^{m}[z^{p+1-m}]g_{m}^{(1)}.$
All terms inside the sums defining $f_{m}^{(1)}$, $m\geq 1$, and $g_{m}^{(1)}$, $m\geq 2$, contain at least some factor $R_{i}(z)$. Since $[z^{0}]R_{i}(z)=0$, by (57), it follows that $[z^{0}]\tilde{R}(z)=[z^{0}]\tilde{R}(z)=0$.
Next, assume $p\geq 1$. Since $\theta\sim\frac{\gamma}{N}$ in the regime (17), it is enough to check that the coefficients $[z^{p+1-m}]f_{m}^{(1)}$, $[z^{p+1-m}]g_{m}^{(1)}$ are polynomials in the $c_{F_{N}}^{\lambda}$ of degree at most $p$ such that the coefficients of the monomials containing at least one $c_{F_{N}}^{\nu}$ with $\ell(\nu)>1$ are $O(N^{m})$ in the regime (17) and all the other coefficients are $o(N^{m})$ and, in fact, we check that these coefficients are $O(N^{m-1})$. We will only verify these bounds for $f^{(1)}_{m}$ indexed by an even integer $m$, as the proofs of the remaining cases are fully analogous.
Let $m=2j$, for some $j\in\mathbb{Z}_{\geq 1}$. For any $k\in\mathbb{Z}_{\geq 0}$, the coefficient $[z^{k}]\Xi_{1}(z)$ is a polynomial in the $c_{F_{N}}^{\lambda}$ of degree at most $k$ with uniformly bounded coefficients in the regime (17) by Lemma 4.6, and for any $1\leq i_{s}<N$, the cofficient $[z^{k}]R_{i_{s}}(z)$ is a polynomial in the $c_{F_{N}}^{\lambda}$ with degree at most $k$ such that the coefficients of the monomials containing at least one $c_{F_{N}}^{\nu}$ with $\ell(\nu)>1$ are $O(1)$ and all the other coefficients are $O(\frac{1}{N})$ in the regime (17). Therefore, for any sequence $2\leq i_{1}<\cdots<i_{j}\leq i$, we have that $[z^{p+1-2j}]\big{(}\Xi_{1}(z)\cdot R_{i_{1}}(z)\cdots R_{i_{j}}(z)\big{)}$ is a polynomial in the $c_{F_{N}}^{\lambda}$ with degree at most $p+1-2j<p$ and such that the coefficients of the monomials containing at least one $c_{F_{N}}^{\nu}$ with $\ell(\nu)>1$ are $O(1)$ and all the other coefficients are $O(\frac{1}{N})$ in the regime (17). For each such sequence, we have that $(i_{1}-2)\prod_{s=1}^{j-1}(i_{s+1}-i_{s}-1)=O(N^{j})$ and the number of such sequences is also $O(N^{j})$, therefore $[z^{p+1-2j}]f_{2j}^{(1)}$ is a polynomial in the $c_{F_{N}}^{\lambda}$ with degree at most $p$ and such that the coefficients of the monomials containing at least one $c_{F_{N}}^{\nu}$ with $\ell(\nu)>1$ are $O(N^{2j})$ and all the other coefficients are $O(N^{2j-1})$ in the regime (17), as desired.
The previous argument ends the proof of 4.10 and, henceforth, we are done. ∎
Lemma 4.11.
Assume that we are in the setting of Theorem 4.2. Then
(63) $\xi_{1}^{\ell}\exp(F_{N})\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\sum_{\Gamma\in\mathbf{L}(\ell)}\bigg{(}\prod_{i\geq 0}\big{(}c_{F_{N}}^{(1)}+i\big{)}^{\text{\# horizontal steps at height $i$ in $\Gamma$}}\\ \cdot\prod_{j\geq 1}\big{(}jc_{F_{N}}^{(j)}+(j+1)c_{F_{N}}^{(j+1)}\big{)}^{\text{\# steps $(1,j)$ in $\Gamma$}}(j+\gamma)^{\text{\# down steps from height $j$ in $\Gamma$}}\bigg{)}+R_{1}^{\prime}+R_{2}^{\prime},$
for some $R_{1}^{\prime}\in I_{1}(\mathcal{F},\ell)$ and $R_{2}^{\prime}\in I_{2}(\mathcal{F},\ell)$.
Proof.
Lemma 4.6 implies that the LHS of (63) is a polynomial in $c_{F_{N}}^{\lambda}$ of degree at most $\ell$ with uniformly bounded coefficients in the regime (17). Therefore, similarly as in the proof of Lemma 4.7, it suffices to prove (63) with $F_{N}(\mathbf{x})$ replaced by
$H_{N}(\mathbf{x})=\sum_{i=1}^{N}{f_{N}(x_{i}-1)},\quad\text{ where }\quad f_{N}(x):=\sum_{k\geq 1}c_{F_{N}}^{(k)}x^{k}.$
With this assumption, it is enough to prove that
(64) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\xi_{1}^{\ell}\exp(H_{N})\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\sum_{\Gamma\in\mathbf{L}(\ell)}\bigg{(}\prod_{i\geq 0}(\kappa_{1}+i)^{\text{\# horizontal steps at height $i$ in $\Gamma$}}\\ \cdot\prod_{j\geq 1}(\kappa_{j}+\kappa_{j+1})^{\text{\# steps $(1,j)$ in $\Gamma$}}(j+\gamma)^{\text{\# down steps from height $j$ in $\Gamma$}}\bigg{)},$
provided that $\lim_{N\to\infty}{\big{(}n\cdot c_{F_{N}}^{(n)}\big{)}}=\kappa_{n}$, for all $n\geq 1$. Let us define the formal power series
(65) $\kappa(x):=\kappa_{1}+\sum_{n\geq 1}(\kappa_{n}+\kappa_{n+1})x^{n}.$
We claim first that (64) follows from the following formula, that will be proved later:
(66) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\xi_{1}^{\ell}\exp(H_{N})\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=(x\partial_{x}+\partial_{x}+\gamma\!\operatorname{d}+\kappa)^{\ell-1}\kappa(x)\big{|}_{x=0},$
where $\kappa$ is the operator of multiplication by $\kappa(x)$, while $\partial_{x}=\frac{d}{dx}$ is the derivative, and $\operatorname{d}$ is the lowering operator:
$\operatorname{d}\left(\sum_{i\geq 0}f_{i}x^{i}\right):=\sum_{i\geq 0}{f_{i+1}x^{i}}.$
Indeed, note that the RHS of (66) can be interpreted as the constant term of a formal power series obtained as a sum of terms obtained by $\ell-1$ applications of operators $x\partial_{x},\partial_{x},\gamma\!\operatorname{d},\kappa$ to $\kappa(x)$. Let $w^{\prime}=w_{\ell}\cdots w_{2}$ be a word of length $\ell-1$, where each $w_{i}$ is one of the operators $\partial_{x},\gamma\!\operatorname{d},x\partial_{x},\kappa_{1}$, or $(\kappa_{j}+\kappa_{j+1})x^{j}$, for some $j\geq 1$. Note that these operators are homogeneous of degrees $-1,-1,0,0,j$, respectively, therefore $w^{\prime}x^{n}\big{|}_{x=0}$ vanishes unless
  1. (A)
    for each $2\leq j\leq\ell$, one has $n+\deg(w_{2})+\cdots+\deg(w_{j})\geq 0$,
  2. (B)
    $\deg(w^{\prime}):=\sum_{j=2}^{\ell}\deg(w_{j})=-n$.
Define $w_{1}:=(\delta_{n>0}\kappa_{n}+\kappa_{n+1})x^{n}$ and $w:=w_{\ell}\cdots w_{1}$, so that $\deg(w)=0$. For each $w$, associate a lattice path $\Gamma$ in $\mathbb{Z}^{2}$ starting at $(0,0)$ with the $i$-th step being $(1,\deg(w_{i}))$; then conditions (A)-(B) are equivalent to $\Gamma$ being a Łukasiewicz path. Note that the height of $\Gamma$ after the $i$-th step is equal to $j$ if and only if $\deg(w_{i}\cdots w_{1})=j$. Together with the following observations:
$(\partial_{x}+\gamma\!\operatorname{d})x^{j}=(j+\gamma)x^{j-1},\quad j\geq 1,\qquad\quad(\kappa_{1}+x\partial_{x})x^{i}=(\kappa_{1}+i)x^{i},\quad i\geq 0,$
this gives that $w^{\prime}\kappa(x)\big{|}_{x=0}=w(1)\big{|}_{x=0}$ is equal to
$\prod_{i\geq 0}{(\kappa_{1}+i)^{\text{\# horizontal steps at height $i$ in $\Gamma$}}}\\ \times\prod_{j\geq 1}{(\kappa_{j}+\kappa_{j+1})^{\text{\# steps $(1,j)$ in $\Gamma$}}(j+\gamma)^{\text{\# down steps from height $j$ in $\Gamma$}}}.$
Together with (66), this proves the desired (64), however it remains to prove (66).
Proof of Equation (66). Recall that the Dunkl operators $\tilde{\xi}_{i}$ were defined in the proof of Lemma 4.8, namely in equation (50). Moreover, we need the following special case of (51):
(67) $\xi_{1}^{\ell}\exp(H_{N}(\mathbf{x}))\Big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\Big{(}\xi_{1}+\tilde{\xi}_{1}\Big{)}^{\ell}\exp\big{(}H_{N}\big{(}\mathbf{x}+1^{N}\big{)}\big{)}\Big{|}_{(x_{1},\dots,x_{N})=(0^{N})}.$
Let us define
(68) $\kappa_{N}(x):=(x_{1}\partial_{1}+\partial_{1})H_{N}(\mathbf{x}+1^{N})\big{|}_{x_{1}\mapsto x}=(x+1)\cdot f_{N}^{\prime}(x)\\ =c_{F_{N}}^{(1)}+\sum_{n\geq 1}\big{(}n\cdot c_{F_{N}}^{(n)}+(n+1)\cdot c_{F_{N}}^{(n+1)}\big{)}x^{n},$
so that, by comparing with (65), we have $\lim_{N\to\infty}{\kappa_{N}(x)}=\kappa(x)$ coefficient-wise. Since $H_{N}$ is symmetric in its variables, then all operators $1-s_{i,j}$ kill it. It then follows from the definition (68) of $\kappa_{N}(x)$ and the definitions (7), (50) of $\xi_{1},\tilde{\xi}_{1}$, that
(69) $\Big{(}\xi_{1}+\tilde{\xi}_{1}\Big{)}^{\ell}\exp(H_{N}(\mathbf{x}+1^{N}))=\Big{(}\xi_{1}+\tilde{\xi}_{1}\Big{)}^{\ell-1}(x_{1}\partial_{1}+\partial_{1})\exp(H_{N}(\mathbf{x}+1^{N}))\\ =\Big{(}\xi_{1}+\tilde{\xi}_{1}\Big{)}^{\ell-1}\left[\kappa_{N}(x_{1})\exp(H_{N}(\mathbf{x}+1^{N}))\right].$
By (67), we are interested in the limit of (69) in the regime (17). We note that $\big{(}\xi_{1}+\tilde{\xi}_{1}\big{)}^{\ell-1}$ is a sum of products or $\ell-1$ operators, each of the form $(x_{1}+1)\partial_{1}$, $\frac{\theta}{x_{1}-x_{j}}(1-s_{1,j})$, of $\frac{\theta x_{j}}{x_{1}-x_{j}}(1-s_{1,j})$, for some $1<j\leq N$. A generic term involves products of $\ell-1$ operators, where all the indices $j$ are distinct. However, (41) implies that all the terms that involve some $\frac{\theta x_{j}}{x_{1}-x_{j}}(1-s_{1,j})$ will not contribute to the limit (69), because the factor $x_{j}$ in it will survive after the action of the operators $(x_{1}+1)\partial_{1}$, $\frac{\theta}{x_{1}-x_{j^{\prime}}}(1-s_{1,j^{\prime}})$, and $\frac{\theta x_{j^{\prime}}}{x_{1}-x_{j^{\prime}}}(1-s_{1,j^{\prime}})$, where $j^{\prime}\neq j$, and therefore the whole contribution will vanish after the substitution $\mathbf{x}=(0^{N})$. Finally, (41) implies that $\frac{\theta}{x_{1}-x_{j}}(1-s_{1,j})=\theta\!\operatorname{d}_{1}+O(x_{j})$, where
$\operatorname{d}_{1}(x_{1}^{\alpha_{1}}\cdots x_{N}^{\alpha_{N}}):=\mathbf{1}_{\alpha_{1}>0}\cdot x_{1}^{\alpha_{1}-1}\cdots x_{N}^{\alpha_{N}},$
therefore
(70) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\left(\xi_{1}+\tilde{\xi}_{1}\right)^{\ell-1}\left[\kappa_{N}(x_{1})\exp(H_{N}(\mathbf{x}+1^{N}))\right]\Big{|}_{(x_{1},\dots,x_{N})=(0^{N})}\\ =\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\big{(}(x_{1}+1)\partial_{1}+(N-1)\theta\!\operatorname{d}_{1}+\kappa_{N}(x_{1})\big{)}^{\ell-1}\left[\kappa_{N}(x_{1})\right]\Big{|}_{(x_{1},\dots,x_{N})=(0^{N})}\\ =\big{(}(x+1)\partial_{x}+\gamma\!\operatorname{d}+\kappa\big{)}^{\ell-1}\left[\kappa(x)\right]\big{|}_{x=0}.$
The desired (66) follows by combining (67), (69) and (70). The proof is now complete. ∎
Proof of Theorem 4.2.
Equation (54) implies that
(71) $\frac{1}{N}\cdot\mathcal{P}_{\ell}e^{F_{N}}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=[z^{\ell}]\frac{1}{N}\Xi_{1}(z)\sum_{i=1}^{N}\sum_{k\geq 0}\left((1-i)\theta z\cdot\Xi_{1}(z)\right)^{k}+R_{1}^{\prime\prime}+R_{2}^{\prime\prime},$
where $R_{1}^{\prime\prime}\in I_{1}(\mathcal{F},\ell)$ and $R_{2}^{\prime\prime}\in I_{2}(\mathcal{F},\ell)$. Next, define the formal power series
(72) $X(z):=1+\sum_{\ell\geq 1}\sum_{\Gamma\in\mathbf{L}(\ell)}\bigg{(}\prod_{i\geq 0}\big{(}c_{F_{N}}^{(1)}+i\big{)}^{\text{\# horizontal steps at height $i$ in $\Gamma$}}\\ \cdot\prod_{j\geq 1}\big{(}jc_{F_{N}}^{(j)}+(j+1)c_{F_{N}}^{(j+1)}\big{)}^{\text{\# steps $(1,j)$ in $\Gamma$}}(j+\gamma)^{\text{\# down steps from height $j$ in $\Gamma$}}\bigg{)}\cdot z^{\ell},$
so that (63) can be rewritten as
(73) $[z^{\ell}]\Xi_{1}(z)=[z^{\ell}]X(z)+R_{1}^{\prime}+R_{2}^{\prime},$
where $R_{1}^{\prime}\in I_{1}(\mathcal{F},\ell)$ and $R_{2}^{\prime}\in I_{2}(\mathcal{F},\ell)$. Moreover, we have
[…]
in the regime (17). Then, by substituting (73) and (74) into (71), we obtain
(75) $\frac{1}{N}\cdot\mathcal{P}_{\ell}e^{F_{N}}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=[z^{\ell}]\left(X(z)\sum_{k\geq 0}\frac{\left(-\gamma z\cdot X(z)\right)^{k}}{1+k}\right)+S_{1}+S_{2},$
where $S_{1}\in I_{1}(\mathcal{F},\ell)$ and $S_{2}\in I_{2}(\mathcal{F},\ell)$. We claim that
(76) $[z^{\ell}]\left(X(z)\sum_{k\geq 0}\frac{\left(-\gamma z\cdot X(z)\right)^{k}}{1+k}\right)=m_{\ell}\left(c^{(1)}_{F_{N}},\cdots,\ell c^{(\ell)}_{F_{N}}\right).$
Evidently, plugging (76) into (75) would finish the proof, so it only remains to prove the claimed equality (76) and we do this next.
For any $\ell\in\mathbb{Z}_{\geq 1}$ and $\Gamma\in\mathbf{L}(\ell)$, consider the $x$-dependent weight of $\Gamma$:
(77) $\operatorname{wt}_{x}(\Gamma):=\big{(}c_{F_{N}}^{(1)}-x\big{)}^{\text{\# horizontal steps at height $0$ in $\Gamma$}}\cdot\prod_{i\geq 1}\big{(}c_{F_{N}}^{(1)}+i\big{)}^{\text{\#\,horizontal steps at height $i$ in $\Gamma$}}\\ \cdot\prod_{j\geq 1}\big{(}jc_{F_{N}}^{(j)}+(j+1)c_{F_{N}}^{(j+1)}\big{)}^{\text{\# steps $(1,j)$ in $\Gamma$}}(j+\gamma)^{\text{\# down steps from height $j$ in $\Gamma$}}$
and let $\operatorname{wt}_{0}(\Gamma)$ be the same weight (77), but with $x=0$. In particular, by the definition (72), we have [^6] We denote by $\ell(\Gamma)$ the length of the Łukasiewicz path $\Gamma$.
(78) $X(z)=\sum_{\ell\geq 0}\sum_{\Gamma\in\mathbf{L}(\ell)}{\operatorname{wt}_{0}(\Gamma)z^{\ell(\Gamma)}}.$
On the other hand, if we let $s\in\mathbb{Z}_{\geq 0}$ be the number of horizontal steps at height 00 of $\Gamma$, and let $1\leq i_{1}<\dots<i_{s}\leq\ell$ be the integers such that $(i_{j}-1,0)\to(i_{j},0)$ are the horizontal steps of $\Gamma$, then by expanding the first factor $(c_{F_{N}}^{(1)}-x)^{s}$ in (77), we deduce
(79) $\operatorname{wt}_{x}(\Gamma)z^{\ell(\Gamma)}=\sum_{A=\{j_{1}<\cdots<j_{m}\}\subseteq\{1,\dots,s\}}\operatorname{wt}_{0}\big{(}\Gamma_{0,\,i_{j_{1}}-1}\big{)}z^{\ell\left(\Gamma_{i_{j_{a}},\,i_{j_{a+1}}-1}\right)}\\ \cdot\prod_{a=1}^{m}{\left(-xz\cdot\operatorname{wt}_{0}\big{(}\Gamma_{i_{j_{a}},\,i_{j_{a+1}}-1}\big{)}z^{\ell\left(\Gamma_{i_{j_{a}},\,i_{j_{a+1}}-1}\right)}\right)},$
where for any subset $A=\{j_{1}<\cdots<j_{m}\}\subset\{1,\dots,s\}$ with cardinality $|A|=m$, we denoted $\Gamma_{i_{j_{a}},\,i_{j_{a+1}}-1}$ the Łukasiewicz path of length $i_{j_{a+1}}-i_{j_{a}}-1$, which is the subpath of $\Gamma$ starting at $(i_{j_{a}},0)$ and finishing at $(i_{j_{a+1}}-1,0)$, and we set $i_{j_{m+1}}:=\ell+1$, so that $\Gamma_{i_{j_{m}},\,i_{j_{m+1}}-1}$ is the subpath of $\Gamma$ starting at $(i_{j_{m}},0)$ and finishing at $(\ell,0)=(\ell(\Gamma),0)$. From (79) and (78), one readily deduces the following identity of formal power series
(80) $\sum_{\ell\geq 0}\sum_{\Gamma\in\mathbf{L}(\ell)}\operatorname{wt}_{x}(\Gamma)z^{\ell(\Gamma)}=X(z)\sum_{m\geq 0}\big{(}-xz\cdot X(z)\big{)}^{m}.$
On the other hand, note that
[…]
Finally, we obtain
[…]
where the first equality follows from Eq. 22 of Definition 3.4, the definition (77) of $\operatorname{wt}_{x}(\Gamma)$ and Eq. 81, while the second equality follows from Eq. 80. This concludes the proof of the desired (76) and, hence, finishing the proof of Theorem 4.2. ∎
4.5. Carleman’s condition: Proof of Theorem 4.4
As $\vec{m}=(m_{n})_{n\geq 1}$ comes from the definition of LLN-satisfaction and we are assuming that the $\mathbb{P}_{N}$ are probability measures in this case, then the $m_{\ell}$’s are the moments of some probability measure. In order to prove the uniqueness, we verify the Carleman’s condition, i.e. we prove that $\sum_{\ell\geq 1}m_{2\ell}^{-\frac{1}{2\ell}}$ diverges. We will use the condition $\vec{m}=\mathcal{J}_{\gamma}^{\kappa\mapsto m}(\vec{\kappa})$ that expresses each moment as a $\kappa$-weighted sum of Łukasiewicz paths, in order to upper bound each $m_{2\ell}$.
Proof of Theorem 4.4.
We use Equation (22) that expresses each $m_{2\ell}$ as a weighted sum over Łukasiewicz paths of length $2\ell$. Take any $\Gamma\in\mathbf{L}(2\ell)$ and let us upper bound the corresponding term in the sum in the RHS of (22).
First, from Eq. 27 for the divided difference operator $\Delta_{\gamma}$, we deduce that for any $\gamma>0$, $n\in\mathbb{Z}_{\geq 0}$, we have
$\bigg{|}\frac{\Delta_{\gamma}\left(x^{1+\text{\#\,horizontal steps at height $0$}}\right)(\kappa_{1})}{1+\text{\#\,horizontal steps at height $0$}}\bigg{|}\leq C_{0}^{\,\text{\# horizontal steps at height $0$}},$
for some sufficiently large $C_{0}>0$ that depends on $\kappa_{1}$ and $\gamma$, but not on $\ell$ or the specific $\Gamma$.
Second, note that the height of any $\Gamma\in\mathbf{L}(2\ell)$ never exceeds $12\ell-$ 1, therefore
$\bigg{|}\prod_{i\geq 1}(\kappa_{1}+i)^{\text{\# horizontal steps at height $i$\,}}\bigg{|}=\prod_{i=1}^{2\ell}|\kappa_{1}+i|^{\text{\,\# horizontal steps atheight $i$}}\\ \leq\prod_{i=1}^{2\ell}(|\kappa_{1}|+2\ell)^{\text{\# horizontal steps at height $i$}}\leq(2\ell C_{1})^{\text{\# horizontal steps at heights $\geq 1$}},$
for some constant $C_{1}>0$. Similarly,
$\bigg{|}\prod_{j\geq 1}(j+\gamma)^{\text{\# down steps from height $j$}}\bigg{|}=\prod_{j=1}^{2\ell}|j+\gamma|^{\text{\,\# down steps from height $j$}}\leq(2\ell C_{2})^{\text{\# down steps}},$
for some constant $C_{2}>0$. Finally, by using the condition of the theorem, we obtain
$\bigg{|}\prod_{j\geq 1}(\kappa_{j}+\kappa_{j+1})^{\text{\# steps $(1,j)$}}\bigg{|}\leq\prod_{j\geq 1}(|\kappa_{j}|+|\kappa_{j+1}|)^{\text{\# steps $(1,j)$}}\leq\prod_{j\geq 1}(C^{j}+C^{j+1})^{\text{\# steps $(1,j)$}}\\ \leq C_{3}^{\,\sum_{j\geq 1}j\,\cdot\,\text{\# steps $(1,j)$}}=C_{3}^{\,\text{\# down steps}},$
for some constant $C_{3}>0$. To explain the last equality, observe that $(1,j)$ is an up step of size $j$, so the sum $\sum_{j\geq 1}j\cdot\text{\# steps$(1,j)$}$ is the total (upward) vertical displacement caused by the up steps in the Łukasiewicz path. Since the path starts and ends at the same horizontal line $y=0$, then the total (downward) vertical displacement caused by down steps should equal this sum and, simultaneously, equal the number of down steps, since each down step goes down exactly by $1$; this justifies the last equality.
From the previous inequalities, the absolute value of the weight of $\Gamma$ in (22) is upper bounded by $(2\ell A)^{\text{\# horizontal steps}\,+\,\text{\# down steps}}\leq(2\ell A)^{2\ell}$, where $A:=\max\{C_{0},C_{1},C_{2}C_{3}\}>0$, for all large $\ell$. It is known that the number of Łukasiewicz paths of length $\ell$ is the $\ell$-th Catalan number:
$|\mathbf{L}(\ell)|=\frac{1}{\ell+1}\binom{2\ell}{\ell}\sim\frac{4^{\ell}}{\ell^{3/2}\sqrt{\pi}},\quad\text{ as }\ell\to\infty,$
therefore
$m_{2\ell}\leq|\mathbf{L}(2\ell)|\cdot(2\ell A)^{2\ell}\leq(4\ell A)^{2\ell},$
for all large $\ell$. Evidently, $m_{2\ell}\geq 0$ because $m_{2\ell}$ is an even moment, consequently, $m_{2\ell}^{-\frac{1}{2\ell}}\geq\frac{1}{4\ell A}$, for all large $\ell$. Since $\sum_{\ell\geq 1}{\frac{1}{4\ell A}}$ diverges, so does $\sum_{\ell\geq 1}m_{2\ell}^{-\frac{1}{2\ell}}$. Thus, the Carleman’s condition is satisfied and the proof is finished. ∎
5. Applications
5.1. Application 1: Jack Littlewood–Richardson coefficients and quantized $\gamma$-convolution
5.1.1. Jack Littlewood–Richardson coefficients
Given partitions $\lambda,\mu,\nu\in\mathbb{Y}(N)$, we will consider the normalization of Jack Littlewood–Richardson coefficients, to be denoted by $c^{\lambda}_{\mu,\nu}(\theta)$, uniquely determined by the equality
(82) $\frac{P_{\mu}(x_{1},\dots,x_{N};\theta)}{P_{\mu}(1^{N};\theta)}\frac{P_{\nu}(x_{1},\dots,x_{N};\theta)}{P_{\nu}(1^{N};\theta)}=\sum_{\lambda\in\mathbb{Y}(N)}{c^{\lambda}_{\mu,\nu}(\theta)\,\frac{P_{\lambda}(x_{1},\dots,x_{N};\theta)}{P_{\lambda}(1^{N};\theta)}}.$
From the definition, it is clear that $c^{\lambda}_{\mu,\nu}(\theta)\neq 0$ implies $|\lambda|=|\mu|+|\nu|$. Also, for any fixed $\mu,\nu\in\mathbb{Y}(N)$, we have
(83) $\sum_{\lambda\in\mathbb{Y}(N)}{c^{\lambda}_{\mu,\nu}(\theta)}=1,$
as deduced by setting $x_{1}=\cdots=x_{N}=1$ in (82).
Let $\mathbb{Y}(N)^{\leq L}$ denote the finite subset of partitions in $\mathbb{Y}(N)$ with longest part at most $L$. If $\mu,\nu\in\mathbb{Y}(N)^{\leq L}$, then $c^{\lambda}_{\mu,\nu}(\theta)\neq 0$ implies that $\lambda\in\mathbb{Y}(N)^{\leq 2L}$. This follows because the largest monomials (in the lexicographic order) in $P_{\mu}(x_{1},\dots,x_{N};\theta)$ and $P_{\nu}(x_{1},\dots,x_{N};\theta)$ are multiples of $x_{1}^{\mu_{1}}\cdots x_{N}^{\mu_{N}}$ and $x_{1}^{\nu_{1}}\cdots x_{N}^{\nu_{N}}$, respectively. Then the largest monomial in the LHS of (82) is a multiple of $x_{1}^{\mu_{1}+\nu_{1}}\cdots x_{N}^{\mu_{N}+\nu_{N}}$. Since $\mu_{1}+\nu_{1}\leq L+L=2L$, by assumption, it follows that any $\lambda\in\mathbb{Y}(N)$ with nonzero $c^{\lambda}_{\mu,\,\nu}(\theta)$ must have $\lambda_{1}\leq 2L$.
Definition 5.1.
For any $N\in\mathbb{Z}_{\geq 1}$ and $\mathbb{P}_{N}^{(1)},\mathbb{P}_{N}^{(2)}$ probability measures on $\mathbb{Y}(N)$, define the signed measure $\mathbb{P}_{N}^{(1)}\boxplus_{\theta}\mathbb{P}_{N}^{(2)}$ on $\mathbb{Y}(N)$ by
$\mathbb{P}_{N}^{(1)}\boxplus_{\theta}\mathbb{P}_{N}^{(2)}(\lambda)=\sum_{\mu,\nu\in\mathbb{Y}(N)}c^{\lambda}_{\mu,\,\nu}(\theta)\mathbb{P}_{N}^{(1)}(\mu)\mathbb{P}_{N}^{(2)}(\nu),\quad\lambda\in\mathbb{Y}(N).$
Due to (83), the signed measure $\mathbb{P}_{N}^{(1)}\boxplus_{\theta}\mathbb{P}_{N}^{(2)}$ has total mass equal to $1$. Moreover, due to the previous discussion, if $\mathbb{P}_{N}^{(1)}$, $\mathbb{P}_{N}^{(2)}$ are supported on $\mathbb{Y}(N)^{\leq L}$, then the support of $\mathbb{P}_{N}^{(1)}\boxplus_{\theta}\mathbb{P}_{N}^{(2)}$ is contained in $\mathbb{Y}(N)^{\leq 2L}$, so that $\mathbb{P}_{N}^{(1)}\boxplus_{\theta}\mathbb{P}_{N}^{(2)}$ is also finitely supported.
Remark 3.
[…]
Remark 4.
When $\theta=\frac{1}{2},1,2$, it is known that the conjecture of Stanley, mentioned in Remark 3, is true. This follows from a representation-theoretic meaning of the Jack Littlewood–Richardson coefficients: up to normalization, $c^{\lambda}_{\mu,\,\nu}(\theta)$ is equal to the multiplicity of the spherical function with highest weight $\lambda$ in the decomposition of the product of the spherical functions of highest weights $\mu,\nu$ for the Gelfand pairs $(\operatorname{GL}_{N}(\mathbb{R}),\operatorname{O}_{N}(\mathbb{R}))$, $(\operatorname{GL}_{N}(\mathbb{C}),\operatorname{U}_{N}(\mathbb{C}))$, and $(\operatorname{GL}_{N}(\mathbb{H}),\operatorname{U}_{N}(\mathbb{H}))$, respectively. Here, $\mathbb{H}$ denotes the skew-field of quaternions.
5.1.2. LLN for $\theta$-tensor products
Fix $L\in\mathbb{Z}_{\geq 1}$, let $\big{\{}\mathbb{P}_{N}^{(1)}\big{\}}_{N\geq 1}$ and $\big{\{}\mathbb{P}_{N}^{(2)}\big{\}}_{N\geq 1}$ be sequences of probability measures on $\mathbb{Y}(N)^{\leq L}$, and let $\mu^{(1)}_{N},\mu^{(2)}_{N}$ be the associated empirical measures. Recall the terminology used in Section 1.2. For a probability measure $\mu$ on $\mathbb{R}$ with finite moments $m_{1},m_{2},\dots$ of all orders (e.g. if $\mu$ is compactly supported), its quantized $\gamma$-cumulants $\kappa_{1},\kappa_{2},\dots$ are defined from the relation $(\kappa_{1},\kappa_{2},\dots)=\mathcal{J}_{\gamma}^{m\mapsto\kappa}\big{(}m_{1},m_{2},\dots\big{)}$.
Theorem 5.2.
Let $\{\theta_{N}>0\}_{N\geq 1}$ be such that $N\theta_{N}\to\gamma$, as $N\to\infty$. Assume that we are in the setting above and that we have the weak limits
(84) $\mu^{(1)}_{N}\to\mu^{(1)},\qquad\mu^{(2)}_{N}\to\mu^{(2)},\qquad\textrm{as $N\to\infty$},$
for certain probability measures $\mu^{(1)}$, $\mu^{(2)}$ on $\mathbb{R}$. Let $\kappa_{n}\big{[}\mu^{(1)}\big{]}$, $\kappa_{n}\big{[}\mu^{(2)}\big{]}$ be the corresponding quantized $\gamma$-cumulants. Then the sequence of signed measures $\big{\{}\mathbb{P}_{N}^{(1)}\boxplus_{\theta_{N}}\mathbb{P}_{N}^{(2)}\big{\}}_{N\geq 1}$ satisfies the LLN, in the sense of Definition 3.2, with corresponding quantized $\gamma$-cumulants, denoted by $\kappa_{n}\big{[}\mu^{(1)}\boxplus^{(\gamma)}\mu^{(2)}\big{]}$, being equal to
(85) $\kappa_{n}\big{[}\mu^{(1)}\boxplus^{(\gamma)}\mu^{(2)}\big{]}=\kappa_{n}\big{[}\mu^{(1)}\big{]}+\kappa_{n}\big{[}\mu^{(2)}\big{]},\quad\textrm{for all }n\in\mathbb{Z}_{\geq 1}.$
If the conjecture from Remark 3 were true, then $\mathbb{P}_{N}^{(1)}\boxplus_{\theta_{N}}\mathbb{P}_{N}^{(2)}$ would be probability measures and we would be able to define the random atomic measures
[…]
$\mu_{N}\to\mu^{(1)}\boxplus^{(\gamma)}\mu^{(2)},\qquad\textrm{as }N\to\infty,$
where $\mu^{(1)}\boxplus^{(\gamma)}\mu^{(2)}$ would be the unique probability measure with quantized $\gamma$-cumulants $\kappa_{n}\big{[}\mu^{(1)}\boxplus^{(\gamma)}\mu^{(2)}\big{]}$, and it would be supported on $[-\gamma,2L]$. In general, if given probability measures $\mu^{(1)}$, $\mu^{(2)}$ (compactly supported or not), there exists a probability measure with quantized $\gamma$-cumulants $\kappa_{n}\big{[}\mu^{(1)}\big{]}+\kappa_{n}\big{[}\mu^{(2)}\big{]}$ and which is uniquely determined by them, then it will be called the quantized $\gamma$-convolution of $\mu^{(1)}$ and $\mu^{(2)}$, and will be denoted by $\mu^{(1)}\boxplus^{(\gamma)}\mu^{(2)}$.
Proof of Theorem 5.2.
Since $N\theta_{N}\to\gamma$, as $N\to\infty$, there exists a constant $C>0$ such that $\sup_{N\geq 1}(\theta_{N}(N-1))\leq C$, and all measures $\mu^{(1)}_{N}$, $\mu^{(2)}_{N}$ are supported on the same compact interval $[-C,L]$. Thus, the weak limits (84) imply that the sequences $\{\mathbb{P}_{N}^{(1)}\}_{N\geq 1}$ and $\{\mathbb{P}_{N}^{(2)}\}_{N\geq 1}$ satisfy the LLN. Therefore Theorem 3.6 implies HT-appropriateness, meaning that the logarithms
$F_{\mathbb{P}_{N}^{(1)},\theta_{N}}:=\ln\left(G_{\mathbb{P}_{N}^{(1)},\theta_{N}}\right),\quad F_{\mathbb{P}_{N}^{(2)},\theta_{N}}:=\ln\left(G_{\mathbb{P}_{N}^{(2)},\theta_{N}}\right),$
have Taylor coefficients with certain limits (as in Definition 3.1), in particular,
(86) $\lim_{N\to\infty}\frac{\partial_{1}^{n}F_{\mathbb{P}_{N}^{(1)},\theta_{N}}}{(n-1)!}\Bigg{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\kappa_{n}\big{[}\mu^{(1)}\big{]},\qquad\lim_{N\to\infty}\frac{\partial_{1}^{n}F_{\mathbb{P}_{N}^{(2)},\theta_{N}}}{(n-1)!}\Bigg{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\kappa_{n}\big{[}\mu^{(2)}\big{]},$
for all $n\in\mathbb{Z}_{\geq 1}$.
Next, since the support of $\mathbb{P}_{N}^{(1)}\boxplus_{\theta_{N}}\mathbb{P}_{N}^{(2)}$ is the finite set $\mathbb{Y}(N)^{\leq 2L}$, this measure is finitely supported and, therefore, it has the small tails property. By Definition 5.1, the Jack generating function of $\mathbb{P}_{N}^{(1)}\boxplus_{\theta_{N}}\mathbb{P}_{N}^{(2)}$, to be denoted by $G_{N,\theta_{N}}(x_{1},\dots,x_{N})$, is equal to
$G_{N,\theta_{N}}(x_{1},\dots,x_{N})$ $=\sum_{\lambda\in\mathbb{Y}(N)}{\mathbb{P}_{N}^{(1)}\boxplus_{\theta_{N}}\mathbb{P}_{N}^{(2)}(\lambda)\,\frac{P_{\lambda}(x_{1},\dots,x_{N};\theta)}{P_{\lambda}(1^{N};\theta)}}$
$=\sum_{\mu,\nu\in\mathbb{Y}(N)}\mathbb{P}_{N}^{(1)}(\mu)\mathbb{P}_{N}^{(2)}(\nu)\sum_{\lambda\in\mathbb{Y}(N)}c^{\lambda}_{\mu,\,\nu}(\theta_{N})\,\frac{P_{\lambda}(x_{1},\dots,x_{N};\theta)}{P_{\lambda}(1^{N};\theta)}$
$=G_{\mathbb{P}_{N}^{(1)},\theta_{N}}(x_{1},\dots,x_{N})\cdot G_{\mathbb{P}_{N}^{(2)},\theta_{N}}(x_{1},\dots,x_{N}),$
and therefore $F_{N,\theta_{N}}(x_{1},\dots,x_{N}):=\ln\left(G_{N,\theta_{N}}(x_{1},\dots,x_{N})\right)$ equals
(87) $F_{N,\theta_{N}}(x_{1},\dots,x_{N})=F_{\mathbb{P}_{N}^{(1)},\theta_{N}}(x_{1},\dots,x_{N})+F_{\mathbb{P}_{N}^{(2)},\theta_{N}}(x_{1},\dots,x_{N}).$
Since $\big{\{}F_{\mathbb{P}_{N}^{(1)},\theta_{N}}\big{\}}_{N\geq 1}$ and $\big{\{}F_{\mathbb{P}_{N}^{(1)},\theta_{N}}\big{\}}_{N\geq 1}$ satisfy the conditions of HT-appropriateness from Definition 3.1, it readily follows that the same conditions are true for $\big{\{}F_{N,\theta_{N}}\big{\}}_{N\geq 1}$, in particular, thanks to (86) and (87), we have
$\lim_{N\to\infty}\frac{\partial_{1}^{n}F_{N,\theta_{N}}}{(n-1)!}\Bigg{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\kappa_{n}\big{[}\mu^{(1)}\big{]}+\kappa_{n}\big{[}\mu^{(2)}\big{]},\quad\textrm{for all $n\in\mathbb{Z}_{\geq 1}$},$
proving the desired (85). Finally, the fact that $\big{\{}\mathbb{P}_{N}^{(1)}\boxplus_{\theta_{N}}\mathbb{P}_{N}^{(2)}\big{\}}_{N\geq 1}$ satisfies the LLN follows from Theorem 3.6. ∎
5.2. Application 2: Jack measures and nonintersecting random walks
Let us discuss the general setting first and then specialize to three special (“pure”) families of examples.
Definition 5.3 ([BO05]).
For any Jack-positive specializations $\rho_{1},\rho_{2}\colon\operatorname{Sym}\to\mathbb{R}$ satisfying the finiteness assumption
(88) $\sum_{\lambda\in\mathbb{Y}}{Q_{\lambda}(\rho_{1};\theta)P_{\lambda}(\rho_{2};\theta)}<\infty,$
the Jack measure $\,\mathbb{P}^{(\theta)}_{\rho_{1};\rho_{2}}(\cdot)$ is the probability measure on the set of all partitions, defined by
(89) $\mathbb{P}^{(\theta)}_{\rho_{1};\rho_{2}}(\lambda):=\frac{1}{H_{\theta}(\rho_{1};\rho_{2})}\cdot Q_{\lambda}(\rho_{1};\theta)P_{\lambda}(\rho_{2};\theta),\quad\lambda\in\mathbb{Y},$
where $H_{\theta}(\rho_{1};\rho_{2})$ and $Q_{\lambda}(;\theta)$ are defined by (13) and (14), respectively.
We are interested in the Jack measures when $\rho_{2}$ is the pure alpha specialization with exactly $N$ parameters $\alpha_{i}$ being equal to $1$, and $\delta=N$, and the rest of Thoma parameters are equal to 00; this specialization was denoted $\operatorname{Alpha}(1^{N})$ in Example 2. In other words, $\rho_{2}$ is defined by
$p_{k}(\rho_{2})=N,\quad\textrm{for all }k\geq 1.$
Then the Jack measure $\mathbb{P}^{(\theta)}_{\rho_{1};\rho_{2}}$ depends only on $N\in\mathbb{Z}_{\geq 1}$ and one specialization $\rho_{1}$, to be denoted $\rho$; in this case, the Jack measure is supported on the subset $\mathbb{Y}(N)\subset\mathbb{Y}$, since $P_{\lambda}(1^{N};\theta)\neq 0$, only if $\lambda\in\mathbb{Y}(N)$. Moreover, observe that the finiteness condition (88) is satisfied if the series
(90) $\sum_{k\geq 1}\frac{|p_{k}(\rho)|}{k}z^{k}$
is absolutely convergent on some open annulus containing the unit circle.
Definition 5.4.
We call a Jack-positive specialization $\rho$ stable if the series (90) is absolutely convergent on $R^{-1}<z<R$, for some $R>1$. In this case, the Jack measure $\mathbb{P}^{(\theta)}_{\rho;\operatorname{Alpha}(1^{N})}$ (which is well-defined and supported on $\mathbb{Y}(N)$) will be denoted by $\mathbb{P}_{\rho;N}^{(\theta)}$ and its Jack generating function will be denoted by $G_{\rho;N,\theta}(x_{1},\dots,x_{N})$.
As an application of the Cauchy identity (13), we deduce
(91) $G_{\rho;N,\theta}(x_{1},\dots,x_{N})=\frac{H_{\theta}(\rho;\mathbf{x}_{N})}{H_{\theta}(\rho;1^{N})}=\prod_{i=1}^{N}\exp\Bigg{\{}\theta\sum_{k\geq 1}\frac{p_{k}(\rho)}{k}\,(x_{i}^{k}-1)\Bigg{\}},$
where $\mathbf{x}_{N}:=(x_{1},\dots,x_{N})$, so its natural logarithm will be
$F_{\rho;N,\theta}(x_{1},\dots,x_{N}):=\ln{G_{\rho;N,\theta}(x_{1},\dots,x_{N})}=\theta\sum_{i=1}^{N}{\sum_{k\geq 1}\frac{p_{k}(\rho)}{k}\,(x_{i}^{k}-1)}.$
As an immediate consequence, we have the following lemma.
Lemma 5.5.
Let $\rho$ be stable and let $F_{\rho;N,\theta}(x_{1},\dots,x_{N})$ be the logarithm of the Jack generating function of $\mathbb{P}_{\rho;N}^{(\theta)}$. Then $\mathbb{P}_{\rho;N}^{(\theta)}$ has the small tails property, and for all $r\geq 2$ and $i_{1},\dots,i_{r}\in\mathbb{Z}_{\geq 1}$ with at least two distinct indices among $i_{1},\dots,i_{r}$, we have $\partial_{i_{1}}\cdots\partial_{i_{r}}F_{\rho;N,\theta}=0$. Moreover, for any $n\in\mathbb{Z}_{\geq 1}$,
(92) $\frac{\partial_{1}^{n}F_{\rho;N,\theta}}{(n-1)!}\,\bigg{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\frac{\theta}{(n-1)!}\cdot\partial_{x}^{n}\sum_{k\geq 1}\frac{p_{k}(\rho)}{k}x^{k}\bigg{|}_{x=1}.$
This lemma shows that the second condition in the definition of HT-appropriateness (namely, Definition 3.1) always holds for $\big{\{}\mathbb{P}^{(\theta)}_{\rho;N}\big{\}}_{N\geq 1}$ with arbitrary stable $\rho$. Moreover, the first condition is also satisfied if (92) has a limit in the regime (17), for all $n\in\mathbb{Z}_{\geq 1}$.
We now switch our focus to Markov chains. Recall from Section 2.7 the definitions of union $(\rho_{1},\rho_{2})$ and $n$-fold union $n\rho$, for specializations $\rho,\rho_{1},\rho_{2}$. These operations preserve the property of Jack-positiveness.
Proposition 5.6 ([GS15]).
For any Jack-positive specializations $\rho_{1},\rho_{2}\colon\operatorname{Sym}\to\mathbb{R}$ satisfying the finiteness assumption (88), define
$p_{\lambda\to\mu}(\rho_{1},\rho_{2};\theta):=\frac{1}{H_{\theta}(\rho_{1};\rho_{2})}\cdot\frac{P_{\mu}(\rho_{2};\theta)}{P_{\lambda}(\rho_{2};\theta)}\cdot Q_{\mu/\lambda}(\rho_{1};\theta),\quad\lambda,\mu\in\mathbb{Y},$
where $Q_{\mu/\lambda}(;\theta)$ is defined by (15). Then $\left(p_{\lambda\to\mu}(\rho_{1},\rho_{2};\theta)\right)_{\lambda,\mu\in\mathbb{Y}}$ is a stochastic matrix, so that its entries define transition probabilities for a Markov chain on the set of partitions $\mathbb{Y}$. Moreover, if $\rho_{3}$ is another Jack-positive specialization such that (88) is satisfied for $\rho_{2},\rho_{3}$, then
$\sum_{\lambda\in\mathbb{Y}}\mathbb{P}^{(\theta)}_{\rho_{1};\rho_{2}}(\lambda)p_{\lambda\to\mu}(\rho_{3},\rho_{2};\theta)$ $=\mathbb{P}^{(\theta)}_{(\rho_{1},\rho_{2});\rho_{3}}(\mu),$
$\sum_{\mu\in\mathbb{Y}}p_{\lambda\to\mu}(\rho_{1},\rho_{2};\theta)p_{\mu\to\rho}(\rho_{3},\rho_{2};\theta)$ $=p_{\lambda\to\rho}((\rho_{1},\rho_{3}),\rho_{2};\theta).$
We are interested in the Markov chain defined in [Hua21] by means of the transition probabilities $p_{\lambda\to\mu}(\rho_{1},\rho_{2};\theta)$ when $\rho_{2}$ is the pure alpha specialization $p_{k}(\rho_{2})=N$, for all $k\geq 1$. Then $p_{\lambda\to\mu}(\rho_{1},\rho_{2};\theta)$ depend only on $N\in\mathbb{Z}_{\geq 1}$ and on one specialization $\rho_{1}$, to be denoted $\rho$; we will denote these quantities by $p_{\lambda\to\mu}(\rho;N;\theta)$.
Definition 5.7 ([Hua21]).
Let $\mathbb{P}_{N}$ be a fixed probability measure on $\mathbb{Y}(N)$ and $\rho$ a Jack-positive specialization. Define the associated Markov chain $\lambda^{(0)},\lambda^{(1)},\lambda^{(2)},\dots\in\mathbb{Y}(N)$ by
$\mathbb{P}\big{(}\lambda^{(0)}=\lambda\big{)}=\mathbb{P}_{N}(\lambda),\quad\mathbb{P}\big{(}\lambda^{(n+1)}=\mu\,\big{|}\,\lambda^{(n)}=\lambda\big{)}=p_{\lambda\to\mu}(\rho;N;\theta),\quad n\in\mathbb{Z}_{\geq 0}.$
Also, define the auxiliary Markov chain
$\mathscr{L}^{(0)},\mathscr{L}^{(1)},\mathscr{L}^{(2)},\dots,\quad\mathscr{L}^{(n)}=\Big{(}\mathscr{L}^{(n)}_{1},\dots,\mathscr{L}^{(n)}_{N}\Big{)},\quad n\in\mathbb{Z}_{\geq 0},$
by $\mathscr{L}^{(n)}_{i}:=\lambda^{(n)}_{i}-\theta(i-1)$, for all $i=1,\dots,N$ and $n\in\mathbb{Z}_{\geq 0}$.
The sequence $\mathscr{L}^{(0)},\mathscr{L}^{(1)},\mathscr{L}^{(2)},\dots$ is a discrete Markov chain of $N$ nonintersecting particles. The initial condition $\mathscr{L}^{(0)}$ is random and determined by $\mathbb{P}_{N}$. In three special cases, the nonintersecting random walks have interesting interpretations: [^8] Note that some interpretations are given in [Hua21] (see immediately before his Theorem 5.8), but they are slightly inaccurate, as the Markov processes are not simply independent walks conditioned on not intersecting.
  • When $\theta=1$ and $\rho$ is the pure alpha specialization (see Example 2) $p_{k}(\rho)=p^{k}$, for some $10<p<$ 1, we have that $\big{\{}\mathscr{L}^{(n)}\big{\}}_{n\geq 0}$ is a geometric-type random walk on $\mathbb{Z}$ in the sense that the $N$ particles can jump to the right by $k$ steps, where $k\in\mathbb{Z}_{\geq 0}$, but subject to certain conditions including the preservation of the nonintersecting property; moreover, when $N=1$, the process is just one particle doing independent jumps with geometric law of parameter $p$.
  • When $\theta=1$ and $\rho$ is the pure beta specialization (see Example 3) $p_{k}(\rho)=(-1)^{k-1}p^{k}$, for some $10<p<$ 1, we have that $\big{\{}\mathscr{L}^{(n)}\big{\}}_{n\geq 0}$ is a Bernoulli-type random walk on $\mathbb{Z}$ in the sense that the $N$ particles can jump to the right by $1$ or do not move, as long as the nonintersecting property is preserved; moreover, when $N=1$, the process is just one particle in $\mathbb{Z}$ moving at each time to the right by one with probability $p$.
  • When $\theta=1$ and $\rho$ is the Plancherel specialization (see Example 4) $p_{k}(\rho)=\chi\cdot\delta_{k,1}$, for some $\chi>0$, we have that $\big{\{}\mathscr{L}^{(n)}\big{\}}_{n\geq 0}$ is a nonintersecting Poisson-type random walk on $\mathbb{Z}$, in the sense that when $N=1$, the process is just one particle doing independent Poisson jumps of intensity $\chi$.
For a general $\theta>0$, the defined Markov chain is a natural one-parameter deformation of log-gas type of the above examples.
We are interested in the fixed-time distribution of the Markov chain in the high temperature regime. In other words, we will study the empirical measure
$\mu^{(t)}_{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{\mathscr{L}^{(t)}_{i}},$
where $t\in\mathbb{Z}_{\geq 0}$ possibly depends on $N$, in the limit as $N\to\infty$, $N\theta\to\gamma$. For this, let us compute the Jack generating function $G^{(t)}_{\rho;N,\theta}(x_{1},\dots,x_{N})$ of the law of $\lambda^{(t)}$. From Proposition 5.6 and Eq. 15, we have that
(93) $\mathbb{P}\left(\lambda^{(t)}=\lambda\right)=\frac{1}{H_{\theta}(t\rho;1^{N})}\cdot\sum_{\lambda^{(0)}\in\mathbb{Y}(N)}\mathbb{P}_{N}(\lambda^{(0)})\cdot\frac{P_{\lambda}(1^{N};\theta)}{P_{\lambda^{(0)}}(1^{N};\theta)}\cdot Q_{\lambda/\lambda^{(0)}}(t\rho;\theta).$
Note that when $\mathbb{P}_{N}(\lambda)=\delta_{\lambda,\emptyset}$, the RHS of (93) is the probability $\mathbb{P}^{(\theta)}_{t\rho;N}(\lambda)$, according to the Jack measure $\mathbb{P}^{(\theta)}_{t\rho;N}$ from Definition 5.4. Let us record this fact, as it will be useful later.
Lemma 5.8.
Let $\{\lambda^{(t)}\}_{t=0,1,2,\dots}$ be the Markov chain from Definition 5.7 with Jack-positive specialization $\rho$ and initial law of $\lambda^{(0)}$ being $\mathbb{P}_{N}(\lambda)=\delta_{\lambda,\emptyset}$. Then the law of $\lambda^{(t)}$ is the Jack measure $\mathbb{P}^{(\theta)}_{t\rho;N}$, where $t\rho=(\rho,\dots,\rho)$ ($t$-fold union).
Remark 5.
Gorin and Shkolnikov defined in [GS15] the continuous Markov process $\{X_{N}^{disc}(s)\colon s\in\mathbb{R}_{\geq 0}\}$ on the discrete space $\mathbb{Y}(N)$, with the law given by (93) with $\rho$ being the pure Plancherel specialization with parameter $\delta=1$, so that $t\rho$ is the pure Plancherel specialization with parameter $\delta=t$. In our case, $t\in\mathbb{Z}_{\geq 0}$, so the process $X_{N}^{disc}(t)$ can be regarded as a continuous version of ours in the special case when $\rho$ is pure Plancherel.
Let us return to (93) again with a general measure $\mathbb{P}_{N}$. The Jack generating function $G^{(t)}_{\rho;N,\theta}(x_{1},\dots,x_{N})$ of the law of $\lambda^{(t)}$ is then calculated as
$G^{(t)}_{\rho;N,\theta}(x_{1},\dots,x_{N})$ $=\frac{1}{H_{\theta}(t\rho;1^{N})}$
×$\quad\times\sum_{\lambda,\lambda^{(0)}\in\mathbb{Y}(N)}\mathbb{P}_{N}(\lambda^{(0)})\cdot\frac{P_{\lambda}(1^{N};\theta)}{P_{\lambda^{(0)}}(1^{N};\theta)}\cdot Q_{\lambda/\lambda^{(0)}}(t\rho;\theta)\cdot\frac{P_{\lambda}(x_{1},\dots,x_{N};\theta)}{P_{\lambda}(1^{N};\theta)}$
$=\frac{1}{H_{\theta}(t\rho;1^{N})}\cdot\sum_{\lambda,\lambda^{(0)}\in\mathbb{Y}(N)}\mathbb{P}_{N}(\lambda^{(0)})\,\frac{Q_{\lambda/\lambda^{(0)}}(t\rho;\theta)P_{\lambda}(x_{1},\dots,x_{N};\theta)}{P_{\lambda^{(0)}}(1^{N};\theta)}$
$=\frac{H_{\theta}(t\rho;\mathbf{x}_{N})}{H_{\theta}(t\rho;1^{N})}\cdot\sum_{\lambda^{(0)}\in\mathbb{Y}(N)}\mathbb{P}_{N}(\lambda^{(0)})\,\frac{P_{\lambda^{(0)}}(x_{1},\dots,x_{N};\theta)}{P_{\lambda^{(0)}}(1^{N};\theta)}$
(94) $=G_{t\rho;N,\theta}(x_{1},\dots,x_{N})\cdot G_{\mathbb{P}_{N},\theta}(x_{1},\dots,x_{N}),$
where $G_{t\rho;N,\theta}(x_{1},\dots,x_{N})$ is the Jack generating function of the Jack measure $\mathbb{P}_{t\rho;N}^{(\theta)}$. We remark that the third equality is a consequence of the identity (16), while the last one follows from (91) and the definition (10) of Jack generating functions.
Theorem 5.9.
(1) Let $\{\theta_{N}>0\}_{N\geq 1}$ be such that $N\theta_{N}\to\gamma$, as $N\to\infty$. Let $\rho=\{\rho_{N}\}_{N\geq 1}$ be a sequence of stable Jack-positive specializations and assume that $\{t_{N}\in\mathbb{Z}_{\geq 1}\}_{N\geq 1}$ are such that the limits
(95) $\kappa_{n}[\mu_{\rho}]:=\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\frac{t_{N}\theta_{N}}{(n-1)!}\cdot\partial_{x}^{n}\sum_{k\geq 1}\frac{p_{k}(\rho_{N})}{k}x^{k}\bigg{|}_{x=1}}$
exist, for all $n\in\mathbb{Z}_{\geq 1}$. Then the sequence of Jack measures $\mathbb{P}_{t_{N}\rho_{N};N}^{(\theta_{N})}$ satisfies the LLN; moreover, if we let $\mu_{\rho}$ be the limit of the empirical measures, then the quantized $\gamma$-cumulants of $\mu_{\rho}$ are exactly the quantities (95).
(2) Additionally, let $\,\mathbb{P}_{N}$ be probability measures on $\mathbb{Y}(N)$ with the small tails property, and let $\{\lambda^{(0)}(N)\in\mathbb{Y}(N)\}_{N\geq 1}$ be a sequence of $\mathbb{P}_{N}$-distributed partitions. Assume that their empirical measures are all supported on the same compact interval $[-L,L]$, and their weak limit, in probability,
(96) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\mu^{(0)}_{N}=\mu^{(0)}$
exists. Then the law of $\lambda^{(t_{N})}(N)$ from the Markov chain $\lambda^{(0)}(N),\lambda^{(1)}(N),\lambda^{(2)}(N),\cdots$, given by Definition 5.7, satisfies the LLN with corresponding quantized $\gamma$-cumulants, denoted $\kappa_{n}[\mu^{(0)}\boxplus^{(\gamma)}\mu_{\rho}]$, being equal to
(97) $\kappa_{n}\big{[}\mu^{(0)}\boxplus^{(\gamma)}\mu_{\rho}\big{]}=\kappa_{n}\big{[}\mu^{(0)}\big{]}+\kappa_{n}\big{[}\mu_{\rho}\big{]},\quad\textrm{for all }n\in\mathbb{Z}_{\geq 1}.$
Proof.
Part (1) follows directly from our main Theorem 3.6, Lemma 5.5, and our assumption (95).
For part (2), due to Theorem 3.6, it will be enough to show that the sequence of laws of $\lambda^{(t_{N})}(N)$ is HT-appropriate. Equations (94), (91), together with the assumptions that $\rho_{N}$ is stable and $\mathbb{P}_{N}$ has the small tails property imply the small tails property for the law of $\lambda^{(t_{N})}(N)$. Moreover, (94) implies that
(98) $\ln G^{(t_{N})}_{\rho_{N};\,N,\theta_{N}}(x_{1},\dots,x_{N})=\ln{G_{\mathbb{P}_{N},\theta_{N}}(x_{1},\dots,x_{N})}+F_{t_{N}\rho_{N};\,N,\theta_{N}}(x_{1},\dots,x_{N}).$
Equation (96) and the fact that all $\mu^{(0)}_{N},\mu^{(0)}$ are compactly supported on the same interval imply also the limit $\mu_{N}^{(0)}\to\mu^{(0)}$ in the sense of moments, in probability. Therefore, Theorem 3.6 shows that the sequence $\{\mathbb{P}_{N}\}_{N\geq 1}$ is HT-appropriate. Combined with Lemma 5.5 and the assumption (95), we deduce that (98) implies that the sequence of laws of $\lambda^{(t_{N})}(N)$ is HT-appropriate with corresponding quantized $\gamma$-cumulants (97), and the proof is finished. ∎
A few remarks are in order. First of all, this theorem shows that the empirical measures of the laws of $\lambda^{(t_{N})}(N)$ converge weakly, in probability, to a probability measure with quantized $\gamma$-cumulants being $\kappa_{n}[\mu^{(0)}]+\kappa_{n}[\mu_{\rho}]$. In the terminology of Section 5.1.1, this means that the quantized $\gamma$-convolution $\mu^{(0)}\boxplus^{(\gamma)}\mu_{\rho}$ exists. Below, we discuss examples when the limits (95) exist for $\rho_{N}$ belonging to the three families of pure specializations, and prove in those cases the existence of $\mu^{(0)}\boxplus^{(\gamma)}\mu_{\rho}$, for all $\mu^{(0)}$ arising as a limit of the form (96).
Secondly, note that the sequence $\{\mathbb{P}_{N}\}_{N\geq 1}$ of probability measures $\mathbb{P}_{N}(\lambda)=\delta_{\lambda,\emptyset}$ have Jack generating functions $G_{\mathbb{P}_{N},\theta}(x_{1},\dots,x_{N})=1$ and the limit of their empirical measures is
(99) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\frac{1}{N}\sum_{i=0}^{N-1}\delta_{-i\theta}=\gamma^{-1}\cdot\mathbf{1}_{[-\gamma,0]}(x)\text{d}x.$
By Theorem 3.6, the corresponding quantized $\gamma$-cumulants are $\kappa_{n}[\gamma^{-1}\cdot\mathbf{1}_{[-\gamma,0]}(x)\text{d}x]=0$, for all $n\in\mathbb{Z}_{\geq 1}$. This could have also been deduced by (97) and the fact that if the law of $\lambda^{(0)}$ is $\mathbb{P}_{N}(\lambda)=\delta_{\lambda,\emptyset}$, then the law of $\lambda^{(t_{N})}$ is the Jack measure $\mathbb{P}^{(\theta_{N})}_{t_{N}\rho_{N};\,N}$ (see Lemma 5.8).
As an interesting consequence, we have the quantized $\gamma$-convolution identity
(100) $\mu\boxplus^{(\gamma)}\big{(}\gamma^{-1}\cdot\mathbf{1}_{[-\gamma,0]}(x)\text{d}x\big{)}=\mu,$
for any probability measure $\mu$. So the measure $\gamma^{-1}\cdot\mathbf{1}_{[-\gamma,0]}(x)\text{d}x$ is a fixed point for the operation of quantized $\gamma$-convolution; in particular, it is also an $\boxplus^{(\gamma)}$-infinitely divisible distribution.
5.2.1. Pure-alpha Jack measure
Let $c\in(0,1)$, $\eta>0$ be arbitrary constants, and let $N\in\mathbb{Z}_{\geq 1}$. Let us consider the pure alpha specialization $\operatorname{Alpha}(c^{\lfloor N\eta\rfloor})$ (see Example 2) defined by $f\big{(}\operatorname{Alpha}(c^{\lfloor N\eta\rfloor})\big{)}:=f(c^{\lfloor N\eta\rfloor})$, for all $f\in\operatorname{Sym}$, or equivalently,
$p_{k}\big{(}\operatorname{Alpha}(c^{\lfloor N\eta\rfloor})\big{)}=\lfloor N\eta\rfloor\cdot c^{k},\ \text{ for all $k\geq 1$}.$
Then the normalization constant for the Jack measure $\mathbb{P}^{(\theta)}_{\operatorname{Alpha}(c^{\lfloor N\eta\rfloor});\,N}$, given by (89), is
$\exp\Bigg{\{}N\theta\sum_{k\geq 1}{\frac{p_{k}\big{(}\operatorname{Alpha}(c^{\lfloor N\eta\rfloor})\big{)}}{k}}\Bigg{\}}=\exp\big{(}-N\theta\lfloor N\eta\rfloor\cdot\ln(1-c)\big{)}=(1-c)^{-N\theta\lfloor N\eta\rfloor}.$
It follows that
$\mathbb{P}^{(\theta)}_{\operatorname{Alpha}(c^{\lfloor N\eta\rfloor});\,N}(\lambda)=(1-c)^{N\theta\lfloor N\eta\rfloor}\,c^{|\lambda|}\,Q_{\lambda}(1^{\lfloor N\eta\rfloor};\theta)P_{\lambda}(1^{N};\theta),$
and this is nonzero whenever $\ell(\lambda)\leq\min\{N,\lfloor N\eta\rfloor\}$. The previous equation can be made more explicit with the help of (14) and the following evaluation identity (see [Mac95, Ch. VI.10, Eqn. (10.20)] or [Sta89, Thm. 5.4]):
(101) $P_{\lambda}(1^{N};\theta)=\prod_{(i,j)\in\lambda}\frac{N\theta+(j-1)-\theta(i-1)}{(\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+\theta},$
valid for all $N\in\mathbb{Z}_{\geq 1}$ and $\lambda\in\mathbb{Y}$. We summarize this discussion in the following.
Definition 5.10.
Let $c\in(0,1)$, $\eta,\theta>0$, and $N\in\mathbb{Z}_{\geq 1}$. The pure-alpha Jack measure associated to these parameters is the probability measure on partitions $\lambda$ with $\ell(\lambda)\leq\min\{N,\lfloor N\eta\rfloor\}$, defined by
$\mathbb{P}^{(\theta)}_{\operatorname{Alpha}(c^{\lfloor N\eta\rfloor});\,N}(\lambda):=(1-c)^{N\theta\lfloor N\eta\rfloor}c^{|\lambda|}\\ \times\prod_{(i,j)\in\lambda}{\frac{(\lfloor N\eta\rfloor\theta+(j-1)-\theta(i-1))(N\theta+(j-1)-\theta(i-1))}{((\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+\theta)((\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+1)}},$
for all $\lambda\in\mathbb{Y}\big{(}\min\{N,\lfloor N\eta\rfloor\}\big{)}$.
We claim that $\left\{\mathbb{P}^{(\theta)}_{\operatorname{Alpha}(c^{\lfloor N\eta\rfloor});\,N}\right\}_{N\geq 1}$ satisfies the two conditions of HT-appropriateness from Definition 3.1, when $c\in(0,1)$, $\eta>0$ are fixed. By Lemma 5.5, the second condition is automatic, while the first follows from Equation (92) and the following calculation:
(102) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\frac{\theta}{(n-1)!}\partial_{x}^{n}\sum_{k\geq 1}\frac{p_{k}\big{(}\operatorname{Alpha}(c^{\lfloor N\eta\rfloor})\big{)}}{k}x^{k}\bigg{|}_{x=1}}=\frac{1}{(n-1)!}\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\theta\lfloor N\eta\rfloor\,\partial_{x}^{n}\sum_{k\geq 1}\frac{(cx)^{k}}{k}\bigg{|}_{x=1}}\\ =-\frac{\gamma\eta}{(n-1)!}{\partial_{x}^{n}\,\ln(1-cx)\Big{|}_{x=1}}=-\frac{\gamma\eta}{(n-1)!}{\frac{-(n-1)!\cdot c^{n}}{(1-cx)^{n}}\bigg{|}_{x=1}}=\frac{\gamma\eta c^{n}}{(1-c)^{n}}.$
Theorem 5.11.
Let $c\in(0,1)$ and $\gamma,\eta>0$ be fixed. Let $\{\theta_{N}>0\}_{N\geq 1}$ be such that $N\theta_{N}\to\gamma$, as $N\to\infty$. Let $\big{\{}\lambda^{(t)}(N)\big{\}}_{t=0,1,2,\dots}$ be the Markov chain from Theorem 5.9 with $\rho_{N}=\operatorname{Alpha}(c)$ (independent of $N$), and initial laws of $\lambda^{(0)}(N)$ given by $\mathbb{P}_{N}$.
(1) If $\mathbb{P}_{N}(\lambda)=\delta_{\lambda,\emptyset}$, then at time $t=\lfloor N\eta\rfloor$, the law of $\lambda^{(t)}(N)=\lambda^{(\lfloor N\eta\rfloor)}(N)$ is the pure-alpha Jack measure $\mathbb{P}^{(\theta_{N})}_{\operatorname{Alpha}(c^{\lfloor N\eta\rfloor});\,N}$. Moreover, this sequence of Jack measures satisfies the LLN and the corresponding empirical measures converge weakly, in probability, to the unique probability measure $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$ with moments
$\int_{\mathbb{R}}{x^{\ell}\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}(\operatorname{d}\!x)}=\sum_{\Gamma\in\mathbf{L}(\ell)}(a+ab)^{\text{\#\,up steps}}\,b^{\text{\#\,down steps}}\ \frac{\Delta_{\gamma}\left(x^{1+\text{\#\,horizontal steps at height $0$}}\right)(ab)}{1+\text{\#\,horizontal steps at height $0$}}\\ \cdot\prod_{j\geq 1}(j+ab)^{\text{\#\,horizontal steps at height $j$}}(j+\gamma)^{\text{\#\,down steps from height $j$}},\quad\text{for all $\ell\in\mathbb{Z}_{\geq 1}$},$
where we set $a:=\gamma\eta>0$ and $b:=c/(1-c)>0$, for convenience.
(2) In the general case, when $\mathbb{P}_{N}$ are such that the empirical measures $\mu_{N}^{(0)}$ of $\lambda^{(0)}(N)$ converge to $\mu^{(0)}$, as in (96), then the empirical measures of $\lambda^{(\lfloor N\eta\rfloor)}(N)$ converge weakly, in probability, to
$\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\mu_{N}^{(\lfloor N\eta\rfloor)}=\mu^{(0)}\boxplus^{(\gamma)}\mu_{\operatorname{Alpha}}^{\gamma;c;\eta},$
where the RHS is understood as the measure with quantized $\gamma$-cumulants $\kappa_{n}\big{[}\mu^{(0)}\big{]}+\kappa_{n}\big{[}\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}\big{]}$. The existence and uniqueness of this probability measure is part of the statement of the theorem.
Proof.
The fact that the law of $\lambda^{(\lfloor N\eta\rfloor)}(N)$ is $\mathbb{P}^{(\theta_{N})}_{\operatorname{Alpha}(c^{\lfloor N\eta\rfloor});\,N}$ if $\mathbb{P}_{N}(\lambda)=\delta_{\lambda,\emptyset}$ follows from Lemma 5.8. Next, thanks to Equations (92) and (102), as well as Lemma 5.5, it follows that the sequence $\Big{\{}\mathbb{P}^{(\theta_{N})}_{\operatorname{Alpha}(c^{\lfloor N\eta\rfloor});\,N}\Big{\}}_{N\geq 1}$ is HT-appropriate with $\kappa_{n}=ab^{n}$. Note that $|\kappa_{n}|=ab^{n}\leq(\max\{a,1\}\cdot b)^{n}$, for all $n\geq 1$, therefore we can apply Theorem 3.6, which completes the proof of part (1). Finally, part (2) follows from Theorem 5.9. ∎
Corollary 5.12.
Assume we are in the context of Theorem 5.11. If $\eta=(1-c)/c$, then the moments of $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}=\mu_{\operatorname{Alpha}}^{\gamma;c;\frac{1-c}{c}}$ can be expressed as
[…]
Proof.
By the proof of Theorem 5.11, $\kappa_{1}=\gamma$ iff $\eta=(1-c)/c$, therefore Lemma 3.5 gives the desired formula. ∎
5.2.2. Pure-beta Jack measure
Let $c>0$ and $N,M\in\mathbb{Z}_{\geq 1}$. Let us consider the pure beta specialization $\operatorname{Beta}((c/\theta)^{M})$ (see Example 3), defined by
$p_{k}\big{(}\operatorname{Beta}((c/\theta)^{M})\big{)}:=(-1)^{k-1}\theta^{-1}M\,c^{k},\ \text{ for all $k\geq 1$}.$
Then the normalization constant for the Jack measure $\mathbb{P}^{(\theta)}_{\operatorname{Beta}((c/\theta)^{M});\,N}$, defined in (93), is
(103) $\exp\Bigg{\{}N\theta\sum_{k\geq 1}{\frac{p_{k}\big{(}\operatorname{Beta}((c/\theta)^{M})\big{)}}{k}}\Bigg{\}}=\exp\Bigg{\{}NM\sum_{k\geq 1}{(-1)^{k-1}\,\frac{c^{k}}{k}}\Bigg{\}}=(1+c)^{NM}.$
Recall that the pure beta specialization $\operatorname{Beta}((c/\theta)^{M})$ is equal to the composition of the $\operatorname{Sym}$-automorphism $\omega_{\theta^{-1}}$ defined by $\omega_{\theta^{-1}}(p_{r})=(-1)^{r-1}\theta^{-1}p_{r}$, for all $r\geq 1$, and the pure alpha specialization $\operatorname{Alpha}(c^{M})$. In particular, since $\omega_{\theta^{-1}}(Q_{\lambda}(;\theta))=P_{\lambda^{\prime}}(;\theta^{-1})$ (see [Mac95, Ch. VI.10, Eqn. (10.17)]), then
(104) $Q_{\lambda}\big{(}\operatorname{Beta}\big{(}(c/\theta)^{M}\big{)};\theta\big{)}=c^{|\lambda|}\cdot P_{\lambda^{\prime}}(1^{M};\theta^{-1}).$
The identities (103)-(104) show that the Jack measure (89) admits the following formula
$\mathbb{P}^{(\theta)}_{\operatorname{Beta}((c/\theta)^{M});\,N}(\lambda)=(1+c)^{-NM}\,c^{|\lambda|}P_{\lambda}(1^{N};\theta)P_{\lambda^{\prime}}(1^{M};\theta^{-1}).$
This is nonzero whenever $\ell(\lambda)\leq N$ and $\ell(\lambda^{\prime})\leq M$, i.e. when the Young diagram of $\lambda$ fits into an $N\times M$ rectangle. By virtue of (101), we obtain the explicit formula in the following definition.
Definition 5.13.
Let $c,\theta>0$, and $N,M\in\mathbb{Z}_{\geq 1}$. Then the pure-beta Jack measure associated to these parameters is the probability measure on partitions $\lambda\subseteq(M^{N})=(\underbrace{M,M,\dots,M}_{N\text{ times}})$, defined by
$\mathbb{P}^{(\theta)}_{\operatorname{Beta}((c/\theta)^{M});\,N}(\lambda):=(1+c)^{-NM}\,c^{|\lambda|}\prod_{(i,j)\in\lambda}\frac{(N\theta+(j-1)-\theta(i-1))(M+\theta(i-1)-(j-1))}{((\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+\theta)((\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+1)},$
for all $\lambda\subseteq(M^{N})$.
We want to prove the HT-appropriateness. The second condition is always true, while the first one follows from Eq. 92 and the following calculation, valid for all $n\in\mathbb{Z}_{\geq 1}$:
(105) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\frac{\partial_{1}^{n}F_{\operatorname{Beta}((c/\theta)^{M});N,\theta}}{(n-1)!}\bigg{|}_{(x_{1},\dots,x_{N})=(1^{N})}}=\frac{M}{(n-1)!}\partial_{x}^{n}\sum_{k\geq 1}\frac{(-1)^{k-1}(cx)^{k}}{k}\bigg{|}_{x=1}=\frac{(-1)^{n-1}Mc^{n}}{(1+c)^{n}}.$
Theorem 5.14.
Let $\gamma,c>0$ and $M\in\mathbb{Z}_{\geq 1}$ be fixed. Let $\{\theta_{N}>0\}_{N\geq 1}$ be such that $N\theta_{N}\to\gamma$, as $N\to\infty$. Let $\big{\{}\lambda^{(t)}(N)\big{\}}_{t=0,1,2,\dots}$ be the Markov chain from Theorem 5.9 with $\rho_{N}=\operatorname{Beta}(c/\theta_{N})$ and initial laws of $\lambda^{(0)}(N)$ given by $\mathbb{P}_{N}$.
(1) If $\mathbb{P}_{N}(\lambda)=\delta_{\lambda,\emptyset}$, then the law of $\lambda^{(M)}(N)$ is the pure-beta Jack measure $\mathbb{P}^{(\theta_{N})}_{\operatorname{Beta}((c/\theta_{N})^{M});\,N}$. Moreover, these Jack measures satisfy the LLN and the corresponding empirical measures converge weakly, in probability, to the unique probability measure $\mu_{\operatorname{Beta}}^{\gamma;c;M}$ with moments
$\int_{\mathbb{R}}{x^{\ell}\mu_{\operatorname{Beta}}^{\gamma;c;M}(\operatorname{d}\!x)}=\sum_{\Gamma\in\mathbf{L}(\ell)}(Mb-M)^{\text{\#\,up steps}}\,(-b)^{\text{\#\,down steps}}\ \frac{\Delta_{\gamma}\left(x^{1+\text{\#\,horizontal steps at height $0$}}\right)(Mb)}{1+\text{\#\,horizontal steps at height $0$}}\\ \cdot\prod_{j\geq 1}(j+Mb)^{\text{\#\,horizontal steps at height $j$}}(j+\gamma)^{\text{\#\,down steps from height $j$}},\quad\text{ for all }\ell\in\mathbb{Z}_{\geq 1},$
where we set $b:=c/(1+c)\in(0,1)$, for convenience.
(2) In the general case, when $\mathbb{P}_{N}$ are such that the empirical measures $\mu_{N}^{(0)}$ of $\lambda^{(0)}(N)$ converge to $\mu^{(0)}$, as in (96), then the empirical measures of $\lambda^{(M)}(N)$ converge weakly, in probability, to
$\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\mu_{N}^{(M)}}=\mu^{(0)}\boxplus^{(\gamma)}\mu_{\operatorname{Beta}}^{\gamma;c;M}.$
Proof.
The law of $\lambda^{(M)}(N)$ is $\mathbb{P}^{(\theta_{N})}_{\operatorname{Beta}((c/\theta_{N})^{M});\,N}$ if $\mathbb{P}_{N}(\lambda)=\delta_{\lambda,\emptyset}$, because of Lemma 5.8. By (92), (105) and Lemma 5.5, this sequence of Jack measures is HT-appropriate with $\kappa_{n}=(-1)^{n-1}Mb^{n}$. Therefore the first statement follows directly from Theorem 5.9. Since $|\kappa_{n}|=Mb^{n}\leq(Mb)^{n}$, for all $n\geq 1$, Theorem 3.6 ends the proof of (1). Part (2) follows from Theorem 5.9. ∎
Corollary 5.15.
In the setting of Theorem 5.14, if $\gamma=Mc/(1+c)$, or equivalently, if $\gamma<M$ and $c=\gamma/(M-\gamma)$, then the moments of the measure $\mu_{\operatorname{Beta}}^{\gamma;c;M}=\mu_{\operatorname{Beta}}^{\gamma;\frac{\gamma}{M-\gamma};M}$ can be expressed as
[…]
Proof.
According to the proof of Theorem 5.14, we have $\kappa_{1}=\gamma$ iff $\gamma<M$ and $c=\gamma/(M-\gamma)$. As a result, Lemma 3.5 gives the desired formula. ∎
Example 6.
As $\mathbb{P}^{(\theta)}_{\operatorname{Beta}((c/\theta)^{M});\,N}$ is supported on partitions $\lambda\subseteq(M^{N})$, it can be regarded as a probability measure on $\{0,1,\dots,N\}^{M}$, by considering the lengths of the columns $(\lambda_{1}^{\prime},\dots,\lambda_{M}^{\prime})$. In particular, when $M=1$, then $\mathbb{P}^{(\theta)}_{\operatorname{Beta}((c/\theta)^{M=1});\,N}$ can be regarded as a probability measure on $\{0,1,\dots,N\}$. We will further consider parameters $\gamma\in(0,1)$ and $c=\gamma/(1-\gamma)$, so that Corollary 5.15 applies. From the evaluation formula (101), we have
$P_{(1^{k})}(1^{N};\theta)=\binom{N}{k},\qquad P_{(k)}(1;\theta^{-1})=1.$
It follows that $\mathbb{P}^{(\theta)}_{\operatorname{Beta}((c/\theta)^{M=1});\,N}$, to be denoted $\mathbb{P}^{(\theta)}_{\operatorname{Beta};\,\gamma,N}$ in this special setting, when regarded as a measure on $\{0,1,\dots,N\}$, turns out to be the binomial distribution $\textrm{Bin}(N,\gamma)$ with parameters $\gamma\in(0,1)$ and $N\in\mathbb{Z}_{\geq 1}$:
$\mathbb{P}^{(\theta)}_{\operatorname{Beta};\,\gamma,N}\{k\}={N\choose k}\gamma^{k}(1-\gamma)^{N-k},\qquad k\in\{0,1,\dots,N\}.$
According to Theorem 5.14 and Corollary 5.15, the random atomic measures
[…]
converge weakly, in probability, in the high temperature regime (17), to the unique probability measure $\mu_{\operatorname{Beta}}^{\gamma;\frac{\gamma}{1-\gamma};1}$ with moments given by (106) with $M=1$, for example:
$m_{1}\Big{[}\mu_{\operatorname{Beta}}^{\gamma;\frac{\gamma}{1-\gamma};1}\Big{]}=\frac{\gamma}{2},\quad m_{2}\Big{[}\mu_{\operatorname{Beta}}^{\gamma;\frac{\gamma}{1-\gamma};1}\Big{]}=\gamma+\frac{\gamma^{2}}{3}-\gamma^{3},\quad m_{3}\Big{[}\mu_{\operatorname{Beta}}^{\gamma;\frac{\gamma}{1-\gamma};1}\Big{]}=\gamma-\frac{7}{4}\gamma^{3}+\gamma^{5}.$
It turns out that
(107) $\mu_{\operatorname{Beta}}^{\gamma;\frac{\gamma}{1-\gamma};1}=\gamma^{-1}\cdot\big{(}\mathbf{1}_{[1-\gamma^{2},1]}(x)+\mathbf{1}_{[-\gamma,-\gamma^{2}]}(x)\big{)}dx,$
as can be proved by computing the (random) characteristic function $\phi_{N,\theta}(s)$ of $\mu_{N,\theta}$,
$\phi_{N,\theta}(s)=\frac{1}{N}\left(\sum_{j=1}^{k}{e^{i(1-\theta(j-1))s}}+\sum_{j=k+1}^{N}{e^{-is\theta(j-1)}}\right)=\frac{e^{is}-e^{-i\theta Ns}+(1-e^{is})e^{-i\theta ks}}{N(1-e^{-i\theta s})},$
and taking the limit of its expectation as $N\to\infty$ and $N\theta\to\gamma$:
(108) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\mathbb{E}_{k\sim\text{Bin}(n,p)}\left[\phi_{N,\theta}(s)\right]}=\gamma\cdot\frac{e^{is}-e^{i(1-\gamma^{2})s}}{i\gamma^{2}s}+(1-\gamma)\cdot\frac{e^{-i\gamma^{2}s}-e^{-i\gamma s}}{i(\gamma-\gamma^{2})s}.$
The RHS of (108) is the characteristic function of the mixture of two continuous uniform distributions that appears in the RHS of (107).
Remark 6.
Theorem 3.6 describes the limiting measure $\mu^{\gamma}$ in the LLN as the unique probability measure with a prescribed moment sequence. In theory, the density of $\mu^{\gamma}$ can then be recovered, e.g. by means of some transform inversion formula; this general procedure is, however, often hard to carry out in practice. In Example 6, we were able to find the density (107) in an alternative way, but the fact that this is the density of $\mu^{\gamma}$ could also have been proven from our general Theorem 3.6 by verifying the combinatorially interesting equalities between the formula (106) for the $\ell$-th moment of $\mu^{\gamma}$ (in terms of Łukasiewicz paths) and
[…]
Carrying out this strategy to calculate the densities of the limiting measures $\mu^{\gamma}$ is beyond the scope of this paper.
5.2.3. Pure-Plancherel Jack measure
Consider the Plancherel specialization $\operatorname{Planch}(t/\theta)$, given by $p_{k}\big{(}\operatorname{Planch}(t/\theta)\big{)}=\frac{t}{\theta}\cdot\delta_{k,1}$, for all $k\geq 1$. As the coefficient of $p_{1}^{|\lambda|}$ in $P_{\lambda}(;\theta)$ is equal to $\prod_{(i,j)\in\lambda}\frac{\theta}{(\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+\theta}$ (see [Mac95, Ch. VI, (10.27),(10.29)]), it follows that
(109) $P_{\lambda}\big{(}\operatorname{Planch}(t/\theta);\theta\big{)}=\prod_{(i,j)\in\lambda}\frac{t}{(\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+\theta},\quad\text{for all }\lambda\in\mathbb{Y}.$
The normalization constant corresponding to the Jack measure $\mathbb{P}^{(\theta)}_{\operatorname{Planch}(t/\theta);\,N}$ in (93) is
(110) $\exp\Bigg{\{}N\theta\sum_{k\geq 1}{\frac{p_{k}\big{(}\operatorname{Planch}(t/\theta)\big{)}}{k}}\Bigg{\}}=\exp(Nt).$
The identities (109)–(110) turn the Jack measure (93) into the following explicit formula.
Definition 5.16.
Let $t,\theta>0$ and $N\in\mathbb{Z}_{\geq 1}$. Then the pure-Plancherel Jack measure associated to these parameters is the probability measure on partitions of length $\leq N$, given by
(111) $\mathbb{P}^{(\theta)}_{\operatorname{Planch}(t/\theta);\,N}(\lambda)=e^{-Nt}t^{|\lambda|}\prod_{(i,j)\in\lambda}{\frac{N\theta+(j-1)-\theta(i-1)}{((\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+\theta)((\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+1)}},$
for all $\lambda\in\mathbb{Y}(N)$.
We show that $\big{\{}\mathbb{P}^{(\theta)}_{\operatorname{Planch}(t/\theta);\,N}\big{\}}_{N\geq 1}$ is HT-appropriate in the sense of Definition 3.1. The second condition is automatic; as for the first, we need (92) and the following calculation:
(112) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\frac{\partial_{1}^{n}\,F_{\operatorname{Planch}(t/\theta);\,N,\theta}}{(n-1)!}\bigg{|}_{(x_{1},\dots,x_{N})=(1^{N})}}=\frac{1}{(n-1)!}\cdot\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\theta\cdot\partial_{x}^{n}\left(\frac{t}{\theta}\,x\right)\Big{|}_{x=1}}=t\cdot\delta_{n,1}.$
Theorem 5.17.
Let $\gamma,t>0$ be fixed. Let $\{\theta_{N}>0\}_{N\geq 1}$ be such that $N\theta_{N}\to\gamma$, as $N\to\infty$. Let $\big{\{}\lambda^{(t)}(N)\big{\}}_{t=0,1,2,\dots}$ be the Markov chain from Theorem 5.9 with $\rho_{N}=\operatorname{Planch}(1/\theta_{N})$ and initial laws of $\lambda^{(0)}(N)$ given by $\mathbb{P}_{N}$.
(1) If $\mathbb{P}_{N}(\lambda)=\delta_{\lambda,\emptyset}$, then the law of $\lambda^{(t)}(N)$ is the pure-Plancherel Jack measure $\mathbb{P}^{(\theta_{N})}_{\operatorname{Planch}(t/\theta_{N});\,N}$. Moreover, these Jack measures satisfy the LLN and the corresponding empirical measures converge weakly, in probability, to the unique probability measure $\mu_{\operatorname{Planch}}^{\gamma;t}$ with moments
(113) $\int_{\mathbb{R}}{x^{\ell}\mu_{\operatorname{Planch}}^{\gamma;t}(\operatorname{d}\!x)}=\sum_{\Gamma\in\mathbf{M}(\ell)}t^{\text{\#\,up steps}}\ \frac{\Delta_{\gamma}\left(x^{1+\text{\#\,horizontal steps at height $0$}}\right)(t)}{1+\text{\#\,horizontal steps at height $0$}}\\ \cdot\prod_{j\geq 1}(j+t)^{\text{\#\,horizontal steps at height $j$}}\,(j+\gamma)^{\text{\#\,down steps from height $j$}},\quad\text{for all }\ell\in\mathbb{Z}_{\geq 1}.$
In this formula, $\mathbf{M}(\ell)$ is the set of all Motzkin paths of length $\ell$; by definition, they are the Łukasiewicz paths of length $\ell$ whose up steps are all $(1,1)$. [^9] In other words, Motzkin paths of length $\ell$, by definition, are the lattice paths with steps $(1,1),(1,0),(1,-1)$, that stay above the $x$-axis, begin at $(0,0)$ and end at $(\ell,0)$.
(2) In the general case, when $\mathbb{P}_{N}$ are such that the empirical measures $\mu_{N}^{(0)}$ of the laws of $\lambda^{(0)}(N)$ converge to $\mu^{(0)}$, as in (96), then the empirical measures of $\lambda^{(t)}(N)$ converge weakly, in probability, to
$\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\mu_{N}^{(t)}}=\mu^{(0)}\boxplus^{(\gamma)}\mu_{\operatorname{Planch}}^{\gamma;t}.$
Proof.
The law of $\lambda^{(t)}(N)$ is $\mathbb{P}^{(\theta_{N})}_{\operatorname{Planch}(t/\theta_{N});\,N}$ if $\mathbb{P}_{N}(\lambda)=\delta_{\lambda,\emptyset}$, because of Lemma 5.8. By (92), (112) and Lemma 5.5, the sequence $\Big{\{}\mathbb{P}^{(\theta_{N})}_{\operatorname{Planch}(t/\theta_{N});\,N}\Big{\}}_{N\geq 1}$ is HT-appropriate with $\kappa_{n}=t\cdot\delta_{n,1}$. Therefore the first statement follows from Theorem 5.9. Since $|\kappa_{n}|=t\cdot\delta_{n,1}\leq t^{n}$, for all $n\geq 1$, Theorem 3.6 proves part (1). Part (2) follows from Theorem 5.9. ∎
Remark 7.
Note that $t$ in $\mathbb{P}^{(\theta_{N})}_{\operatorname{Planch}(t/\theta_{N});\,N}$ does not necessarily have to be an integer, but it can also be a positive real number. In this case, we obtain the LLN in the high temperature regime for the fixed-time distribution of the continuous Markov chain of Gorin–Schkolnikov mentioned in Remark 5.
Corollary 5.18.
In the setting of Theorem 5.17, if $t=\gamma$, the moments of $\mu_{\operatorname{Planch}}^{\gamma;\gamma}$ are equal to
[…]
Proof.
By the proof of Theorem 5.17, $\kappa_{1}=\gamma$ iff $t=\gamma$; then apply Lemma 3.5. ∎
6. The gamma-deformed R-transform and moment generating function
As a consequence of our main Theorem 3.6, we can define, for any $\gamma>0$, the sequence of quantized $\gamma$-cumulants $\{\kappa_{n}^{\gamma}[\mu]\}_{n\geq 1}$ for probability measures $\mu$ that are weak limits, in probability, of atomic measures of the form (21). In fact, the quantities $\kappa_{n}^{\gamma}[\mu]$ are obtained by first computing the moments $m_{n}(\mu)$ and then applying the recursive combinatorial formula from Definition 3.4. There is certain parallelism to Speicher’s combinatorial approach to Free Probability, where free cumulants of probability measures are obtained from their moments by a recursive formula that involves noncrossing set partitions [Spe94].
Speicher’s approach to Free Probability, however, cannot be used to define free cumulants for probability measures that do not have finite moments of all orders. This is where the analytic approach to Free Probability shines, as instead of working with the sequence of cumulants, one works with the $R$-transform and proves its existence as an analytic function for a larger class of probability measures (not only those having finite moments of all orders) by means of a functional equation between the $R$-transform and Cauchy transform. Envisioning a Quantized $\gamma$-deformed Probability Theory, we also want to find functional equations between the series
$R^{\gamma}_{\mu}(z):=\sum_{n\geq 1}{\frac{\kappa_{n}^{\gamma}[\mu]z^{n}}{n}},$
which we call the $\gamma$-quantized $R$-transform of $\mu$, and the Cauchy transform (or moment generating function) of $\mu$. This motivated the following result.
Theorem 6.1.
Let $\gamma>0$ and assume that $\vec{\kappa}=(\kappa_{1},\kappa_{2},\dots)$ and $\vec{m}=(m_{1},m_{2},\dots)$ are related by the combinatorial identity from Definition 3.4, i.e. $\vec{\kappa}=\mathcal{J}_{\gamma}^{\kappa\mapsto m}(\vec{m})$. Then there exists a unique $\vec{c}=(c_{1},c_{2},\dots)$ such that the following relations between generating functions hold:
(114) $\exp\left(\sum_{n\geq 1}\frac{\kappa_{n}z^{n}}{n}\right)$ $=1+\sum_{n\geq 1}\frac{c_{n}}{\gamma^{\uparrow n}}z^{n},$
(115) $1+\sum_{n\geq 1}\frac{(-1)^{n}c_{n}}{z^{\uparrow n}}$ $=\exp\left(\gamma\cdot\widetilde{\mathcal{L}}\left\{\frac{t}{1-e^{-t}}\sum_{n\geq 1}\left(\frac{(-1)^{n}m_{n}}{n!}-\frac{\gamma^{n}}{(n+1)!}\right)t^{n-1}\right\}(z)\right).$
Recall that $z^{\uparrow n}=z(z+1)\cdots(z+n-1)$ is the raising factorial, while we define $\widetilde{\mathcal{L}}$ as the linear operator that transforms power series in $t$ into power series in $z^{-1}$, and is uniquely defined by
(116) ~$\widetilde{\mathcal{L}}\left\{\sum_{n\geq 0}{\frac{a_{n}}{n!}t^{n}}\right\}(z):=\sum_{n\geq 0}{a_{n}z^{-n-1}}.$
Note that (114) is an equality between power series in $z$, which furnishes formulas between $(\kappa_{1},\kappa_{2},\dots)$ and $(c_{1},c_{2},\dots)$. Likewise, (115) is an equality between power series in $z^{-1}$ that yields formulas between $(c_{1},c_{2},\dots)$ and $(m_{1},m_{2},\dots)$. Put together, these two equalities can be used to express each $m_{\ell}$ in terms of $\kappa_{1},\kappa_{2},\dots$ (and viceversa); Theorem 6.1 states that this gives an equivalent expression as the one in terms of Łukasiewicz paths from Definition 3.4.
Remark 8.
We call $\widetilde{\mathcal{L}}$ the formal Laplace transform because of its relation with the Laplace transform
[…]
Observe that if $|f(t)|<\exp(bt)$, for some $b>0$ and all large $t\gg 0$, then $\mathcal{L}\{f(t)\}(z)$ is analytic in the region $\Re z>b$. In particular, if $f(t)$ has subexponential growth, then $\mathcal{L}\{f(t)\}(z)$ is analytic in $\Re z>0$; this is the case, for instance, when $f(t)$ is a polynomial. Since $\mathcal{L}\{t^{n}\}(z)=n!\,z^{-n-1}$, for all $n\in\mathbb{Z}_{\geq 0}$ and $\Re z>0$, then $\widetilde{\mathcal{L}}$ coincides with $\mathcal{L}$ on polynomials in the region $\Re z>0$; this agreement can be extended to more general functions under suitable analytic conditions. In Theorem 6.1, we decided to use $\widetilde{\mathcal{L}}$ instead of $\mathcal{L}$ to avoid specifying any particular choice of analytic conditions and domains for $z$.
Remark 9.
Note that
(117) $\frac{t}{1-e^{-t}}=\sum_{m\geq 0}\frac{(-1)^{m}B_{m}}{m!}t^{m}$
is the exponential generating function of the Bernoulli numbers, the first few of them being: $B_{0}=1$, $B_{1}=-\frac{1}{2}$, $B_{2}=\frac{1}{6}$, $B_{3}=0$, $B_{4}=-\frac{1}{30}$, $B_{5}=0,\cdots$. This shows that the relation between $c_{1},c_{2},\dots$ and $m_{1},m_{2},\dots$ from (115) is expressed in terms of these quantities. The occurrence of the Bernoulli numbers in the theorem can be traced to an asymptotic expansion of the log gamma function in the proof below; see Lemma 6.4.
Remark 10.
Another reason for the interest in Theorem 6.1 is the (possible) connection with Finite Free Probability, see e.g. [MSS22]. In fact, similar identities were proved in the continuous settings of $\beta$-eigenvalues, see [BGCG22, Thm. 3.11] and $\beta$-singular values, see [Xu23, Thm. 5.8]. It is now known that the identities in those references are precisely the same as in the cumulant-based combinatorial approach to Finite Free Probability started in [AP18], but see also [AGVP23, Cue24], after the formal identification of the parameter $\gamma\mapsto\textrm{minus the degree of the polynomials}$. Whether our equations (114)–(115) have any relation to Finite Free Probability is an open problem.
6.1. Preliminary lemmas
We need a few lemmas for the proof of Theorem 6.1. For the first two of them, we denote by $Q^{*}_{\mu}(x_{1},\dots,x_{N};\theta)$, $\mu\in\mathbb{Y}(N)$, the $Q$-normalization of the shifted Jack polynomials, as defined in [OO97] or [OO98, Secs. 2.2–2.4]. In particular, for row partitions $\mu=(k)$, they admit the explicit expression
(118) $Q_{(k)}^{*}(x_{1},\dots,x_{N};\theta)=\sum_{1\leq i_{1}\leq\cdots\leq i_{k}\leq N}{\frac{\theta^{\uparrow m_{1}}\theta^{\uparrow m_{2}}\cdots}{m_{1}!m_{2}!\cdots}(x_{i_{1}}-k+1)\cdots(x_{i_{k-1}}-1)x_{i_{k}}},$
where $m_{l}:=\#\{r\mid i_{r}=l\}$ are the multiplicities of $l=1,2,\dots$ in the sequence $i_{1},\dots,i_{k}$.
Lemma 6.2.
If $|x_{i}|\leq A$, for all $i=1,\dots,N$ and some $A>0$, then
(119) $\big{|}Q_{(k)}^{*}(x_{1},\dots,x_{N};\theta)\big{|}\leq A^{\uparrow k}\cdot\sum_{1\leq i_{1}\leq\cdots\leq i_{k}\leq N}{\frac{\theta^{\uparrow m_{1}}\theta^{\uparrow m_{2}}\cdots}{m_{1}!m_{2}!\cdots}},$
for all $k\in\mathbb{Z}_{\geq 1}$, where $m_{l}:=\#\{r\mid i_{r}=l\}$ are the multiplicities of $l=1,2,\dots$ in the sequence $i_{1},\dots,i_{k}$.
Proof.
This follows immediately from (118). ∎
Lemma 6.3.
If $|x_{i}|\leq A$, for all $i=1,\dots,N$ and some $A>0$, then the series
(120) $\sum_{k\geq 0}{\frac{(-1)^{k}Q_{(k)}^{*}(x_{1},\dots,x_{N};\theta)}{z^{\uparrow k}}}$
converges absolutely in the region $\Re z>N\theta+A$, and in fact it is identically equal to
(121) $\prod_{i=1}^{N}{\frac{\Gamma(x_{i}+z-i\theta)}{\Gamma(x_{i}+z-(i-1)\theta)}\frac{\Gamma(z-(i-1)\theta)}{\Gamma(z-i\theta)}}.$
Proof.
This is proved in [OO98, Appendix], so we only comment on the bound $\Re z>N\theta+A$, which is not explicitly stated in that reference. Denote the RHS of (119) by $q_{A;\,\theta}^{(N)}(k)$, e.g. $q_{A;\,\theta}^{(1)}(k)=A^{\uparrow k}\theta^{\uparrow k}/k!$. To prove that (120) converges absolutely in the region $\Re z>N\theta+A$, it suffices to show by Lemma 6.2 that the series
$F^{(N)}_{A;\,\theta}(y):=\sum_{k\geq 0}{\frac{q_{A;\,\theta}^{(N)}(k)}{y^{\uparrow k}}}$
converges if $y>N\theta+A$; in fact, by induction on $N$, we show it converges to $\frac{\Gamma(y-A-N\theta)\Gamma(y)}{\Gamma(y-A)\Gamma(y-N\theta)}$.
For the base case $N=1$, note that $F^{(1)}_{A;\,\theta}(y)$ is Gauss’s hypergeometric series ${}_{2}F_{1}(\theta,A;y;1)$, for which it is well-known that it converges absolutely when $y>\theta+A$, and equals $\frac{\Gamma(y-A-\theta)\Gamma(y)}{\Gamma(y-A)\Gamma(y-\theta)}$, see e.g. [Bai35, Sec. 1.3]. For the inductive step, use the definitions and inductive hypothesis to deduce
(122) $F_{A;\,\theta}^{(N)}(y)=\sum_{k\geq 0}{F_{A+k;\,\theta}^{(N-1)}(y+k)\frac{q^{(1)}_{A;\,\theta}(k)}{y^{\uparrow k}}}=F_{A;\,\theta}^{(N-1)}(y)\sum_{k\geq 0}{\frac{F_{A+k;\,\theta}^{(N-1)}(y+k)}{F_{A;\,\theta}^{(N-1)}(y)}\frac{q^{(1)}_{A;\,\theta}(k)}{y^{\uparrow k}}}\\ =F_{A;\,\theta}^{(N-1)}(y)\sum_{k\geq 0}{\frac{\Gamma(y+k)\Gamma\big{(}y-(N-1)\theta\big{)}}{\Gamma\big{(}y+k-(N-1)\theta\big{)}\Gamma(y)}\frac{q^{(1)}_{A;\,\theta}(k)}{y^{\uparrow k}}}=F_{A;\,\theta}^{(N-1)}(y)\sum_{k\geq 0}{\frac{q^{(1)}_{A;\,\theta}(k)}{\big{(}y-(N-1)\theta\big{)}^{\uparrow k}}}.$
The last summation above is the hypergeometric series ${}_{2}F_{1}(\theta,A;y-(N-1)\theta;1)$, which converges if $y-(N-1)\theta>\theta+A$, and equals $\frac{\Gamma(y-A-N\theta)\Gamma(y-(N-1)\theta)}{\Gamma(y-(N-1)\theta-A)\Gamma(y-N\theta)}$ in that case. Plugging back into (122) and using the inductive hypothesis again concludes the induction.
A small refinement of this argument proves (121); see [OO98, Appendix] for details. ∎
Recall that for a function $F(z)$ defined on the half-plane $|\arg z|<\pi/2$, the asymptotic expansion
$F(z)\sim\sum_{m=0}^{\infty}{b_{m}z^{-m}},$
as $z\to\infty$, means that
(123) $\lim_{\begin{subarray}{c}|\arg(z)|<\pi/2\\ z\to\infty\end{subarray}}{z^{N}\left(F(z)-\sum_{m=0}^{N}{b_{m}z^{-m}}\right)}=0,\quad\textrm{for all }N\geq 0.$
Lemma 6.4.
The following is the asymptotic expansion of $\ln\Gamma(z)$, in the sense above,
(124) $\ln\Gamma(z)\sim\left(z-\frac{1}{2}\right)\ln z-z+\frac{\ln(2\pi)}{2}+\widetilde{\mathcal{L}}\left\{-\frac{1}{t^{2}}+\frac{1}{2t}+\frac{1}{t(e^{t}-1)}\right\}(z).$
Proof.
By Eq. 117 and the fact that $B_{2m+1}=0$, for all $m\geq 1$, we have
$-\frac{1}{t^{2}}+\frac{1}{2t}+\frac{1}{t(e^{t}-1)}=\sum_{m=1}^{\infty}{\frac{B_{2m}}{2m(2m-1)}t^{2m-2}},$
in a neighborhood of zero, therefore
~$\widetilde{\mathcal{L}}\left\{-\frac{1}{t^{2}}+\frac{1}{2t}+\frac{1}{t(e^{t}-1)}\right\}(z)=\sum_{m=1}^{\infty}{\frac{B_{2m}}{2m(2m-1)}z^{1-2m}}.$
In terms of these Bernoulli numbers, the asymptotic expansion (124) is stated in [Cop04, top of page 53]. [^10] There is a typo in the referenced equation from [Cop04]: $\frac{1}{p^{2m}}$ should be $\frac{1}{p^{2m-1}}$.
Lemma 6.5.
For any formal power series $f(t)$ and $x\in\mathbb{R}$, we have
(125) ~$\widetilde{\mathcal{L}}\{f(t)\}(z+x)=\widetilde{\mathcal{L}}\{e^{-tx}f(t)\}(z).$
Proof.
As mentioned in Remark 8, the formal Laplace transform and the usual Laplace transform agree on polynomials of arbitrary degree. They also agree on functions of $t$ of the form $e^{-xt}\sum_{i=0}^{N}{a_{i}t^{i}}$, for any $x>0$. So for any polynomial $f(t)$ and $x>0$, we deduce
[…]
i.e. (125) holds if $f(t)$ is a polynomial and $x>0$. The desired Eq. 125 is an identity of formal power series in $z^{-1}$, meaning that we want to prove that the coefficients on both sides agree. But such coefficients are just polynomials in $x$ and the coefficients of $f(t)$; since they agree for any $x>0$ and any polynomial $f(t)$, they must also agree for any $x\in\mathbb{R}$ and any formal power series $f(t)$, as needed. ∎
Lemma 6.6.
Let $\big{\{}L^{(N)}=\big{(}L^{(N)}_{1},\dots,L^{(N)}_{N}\big{)}\big{\}}_{N\geq 1}$ be a sequence of tuples of real numbers such that
(126) $\lim_{N\to\infty}{\frac{1}{N}\sum_{i=1}^{N}{\Big{(}L^{(N)}_{i}\Big{)}^{\ell}}}=M_{\ell},$
for some real numbers $M_{1},M_{2},\dots$ and all $\ell\in\mathbb{Z}_{\geq 1}$. Then
$\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\sum_{i=1}^{N}\left(\Big{(}L_{i}^{(N)}-\theta\Big{)}^{\ell}-\Big{(}L_{i}^{(N)}\Big{)}^{\ell}+\big{(}(1-i)\theta\big{)}^{\ell}-(-i\theta)^{\ell}\right)=(-1)^{\ell-1}\gamma^{\ell}-\ell\gamma M_{\ell-1},$
for all $\ell\in\mathbb{Z}_{\geq 1}$, where we set $M_{0}:=1$.
Proof.
By a simple calculation, the prelimit expression can be written as
$\sum_{i=1}^{N}\left(\Big{(}L_{i}^{(N)}-\theta\Big{)}^{\ell}-\Big{(}L_{i}^{(N)}\Big{)}^{\ell}+(-\theta)^{\ell}\Big{(}(i-1)^{\ell}-i^{\ell}\Big{)}\right)\\ =-\ell(\theta N)\cdot\frac{1}{N}\sum_{i=1}^{N}\left(L_{i}^{(N)}\right)^{\ell-1}+\sum_{k=2}^{\ell}\sum_{i=1}^{N}\binom{\ell}{k}\left(L_{i}^{(N)}\right)^{\ell-k}(-\theta)^{k}-(-\theta)^{\ell}N^{\ell}.$
The first term above converges to $-\ell\gamma M_{\ell-1}$, in the regime where $N\to\infty$, $N\theta\to\gamma$, due to the assumption (126). The same relation shows that the second term converges to zero. The last term evidently converges to $(-1)^{\ell-1}\gamma^{\ell}$, finishing the proof. ∎
We are finally ready to prove Theorem 6.1. It should be noted that the main tool will be Theorem 3.6, which in turn was proved by employing Jack generating functions and Cherednik operators. It would be desirable to have an elementary proof along the same lines as [Cue24, Main Result I]. We leave finding a combinatorial proof as an open problem.
6.2. Proof of Theorem 6.1
Let $\mu$ be any compactly supported probability measure on $\mathbb{R}$ for which there exists a sequence of partitions $\{\lambda^{(N)}\in\mathbb{Y}(N)\}_{N\geq 1}$ such that $C\geq\lambda^{(N)}_{1}\geq\cdots\geq\lambda^{(N)}_{N}\geq 0$, for some constant $C>0$, and such that we have the following weak limit
$\frac{1}{N}\sum_{i=1}^{N}{\delta_{\mathscr{L}_{i}^{(N)}}}\to\mu,\quad\text{as }N\to\infty,\ N\theta\to\gamma,$
where $\mathscr{L}_{i}^{(N)}=\lambda_{i}^{(N)}-\theta(i-1)$, for all $1\leq i\leq N$. As $\mu$ is compactly supported, it is implied that the sequence of probability measures
(127) $\mathbb{P}_{N}:=\frac{1}{N}\sum_{i=1}^{N}{\delta_{\mathscr{L}_{i}^{(N)}}}$
satisfies the LLN, in the sense of Definition 3.2. In particular, if we let $m_{1},m_{2},\cdots$ be the moments of $\mu$, then
(128) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\frac{1}{N}\sum_{i=1}^{N}{\Big{(}\mathscr{L}_{i}^{(N)}\Big{)}^{k}}}=m_{k},\quad k\in\mathbb{Z}_{\geq 1}.$
The Jack generating function of $\mathbb{P}_{N}$ is equal to
$G_{N,\theta}(x_{1}+1,\dots,x_{n}+1)$ $=\frac{P_{\lambda^{(N)}}(x_{1}+1,\dots,x_{N}+1;\theta)}{P_{\lambda^{(N)}}(1^{N};\theta)}$
(129) $=\sum_{\mu\in\mathbb{Y}(N)}\frac{Q_{\mu}^{*}\Big{(}\lambda_{1}^{(N)},\dots,\lambda_{N}^{(N)};\theta\Big{)}P_{\mu}\big{(}x_{1},\dots,x_{N};\theta\big{)}}{(N\theta)_{\mu}},$
where the second equality above is proved in [OO97, Thm. (3.2)]. In (129), the symbol in the denominator is $(g)_{\mu}:=\prod_{(i,j)\in\mu}\big{(}g+(j-1)-\theta(i-1)\big{)}$.
We want to specialize the variables $x_{1}=z$ and $x_{2}=\cdots=x_{N}=0$ in (129). Using the facts (i) $P_{\mu}(z,0^{N-1};\theta)=0$, if $\ell(\mu)>1$, (ii) $P_{(k)}(z,0^{N-1};\theta)=z^{k}$, for all $k\in\mathbb{Z}_{\geq 0}$, (iii) $Q_{(0)}^{*}(x_{1},\dots,x_{N};\theta)=1$, and (iv) $(N\theta)_{(k)}=(N\theta)^{\uparrow k}$, for all $k\in\mathbb{Z}_{\geq 0}$, we deduce:
[…]
Since $\{\mathbb{P}_{N}\}_{N\geq 1}$ satisfies the LLN, Theorem 3.6 implies HT-appropriateness. In particular, we have the following convergence of Taylor coefficients of the logarithms:
[…]
This has two implications. First, the functions $G_{N,\theta}(z+1,1^{N-1})$ themselves have converging Taylor coefficients, so due to the expansion (130), we deduce that the following limits exist:
(132) $c_{k}=\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}Q_{(k)}^{*}\Big{(}\lambda_{1}^{(N)},\dots,\lambda_{N}^{(N)};\theta\Big{)},\quad\text{for all }k\in\mathbb{Z}_{\geq 1}.$
Second, by (130) and (131), we actually deduce the following equality
$\exp\left(\sum_{n\geq 1}\frac{\kappa_{n}z^{n}}{n}\right)=1+\sum_{k\geq 1}\frac{c_{k}}{\gamma^{\uparrow k}}z^{k},$
that is to be interpreted coefficient-wise, i.e. as an equality between formal power series. This is precisely the first equation (114) in the statement of the theorem. We have thus proved that the quantities $c_{1},c_{2},\dots$, uniquely determined from $\kappa_{1},\kappa_{2},\dots$ by this equation, are exactly the limits (132) of certain evaluations of shifted Jack polynomials.
On the other hand, according to Lemma 6.6, we have the following generating function in $z^{-1}$ for shifted Jack polynomials:
(133) $1+\sum_{k\geq 1}{\frac{(-1)^{k}Q_{(k)}^{*}(x_{1},\dots,x_{N};\theta)}{z^{\uparrow k}}}=\prod_{i=1}^{N}{\frac{\Gamma(x_{i}+z-i\theta)}{\Gamma(x_{i}+z-(i-1)\theta)}\frac{\Gamma(z-(i-1)\theta)}{\Gamma(z-i\theta)}}.$
If we set $x_{i}=\lambda_{i}^{(N)}$, for $i=1,\dots,N$, then (133) is, in fact, an equality between analytic functions in the region $\Re z>N\theta+C$, as shown in Lemma 6.3. We will soon be concerned with the high temperature regime where $N\to\infty$, $N\theta\to\gamma$; in this regime, the identity (133) holds whenever $\Re z\gg 0$, i.e., in a region $\{z\in\mathbb{C}\colon\Re z>C^{\prime}\}$, for some large enough $C^{\prime}>0$.
By taking the logarithm of (133), we have
(134) $\ln\left(1+\sum_{k\geq 1}\frac{(-1)^{k}Q_{(k)}^{*}\Big{(}\lambda_{1}^{(N)},\dots,\lambda_{N}^{(N)};\theta\Big{)}}{z^{\uparrow k}}\right)=\sum_{i=1}^{N}\sum_{j=1}^{4}{(-1)^{j-1}\ln\Gamma\left(z+\nabla_{i;j}^{(N)}\right)},$
where
$\nabla_{i;1}^{(N)}:=\mathscr{L}_{i}^{(N)}-\theta,\quad\nabla_{i;2}^{(N)}:=\mathscr{L}_{i}^{(N)},\quad\nabla_{i;3}^{(N)}:=(1-i)\theta,\quad\nabla_{i;4}^{(N)}:=-i\theta,$
and (134) is valid whenever $\Re z\gg 0$. Next, we want to consider the asymptotic expansion of both sides in the region $|\arg z|<\pi/2$, as $z\to\infty$, in the sense of Eq. 123. The LHS is the same because it is an analytic function, while the asymptotic expansion of the RHS can be deduced from Lemmas 6.4 and 6.5. We end up with the following equality, which should be interpreted as an equality between formal power series in $z^{-1}$:
$\ln\left(1+\sum_{k\geq 1}\frac{(-1)^{k}Q_{(k)}^{*}\Big{(}\lambda_{1}^{(N)},\dots,\lambda_{N}^{(N)};\theta\Big{)}}{z^{\uparrow k}}\right)=\sum_{i=1}^{N}\sum_{j=1}^{4}(-1)^{j-1}\left(z+\nabla_{i;j}^{(N)}-\frac{1}{2}\right)\ln\left(1+\frac{\nabla_{i;j}^{(N)}}{z}\right)$
(135) +~$+\widetilde{\mathcal{L}}\left\{\sum_{i=1}^{N}\sum_{j=1}^{4}(-1)^{j-1}e^{-t\nabla_{i;j}^{(N)}}\left(-\frac{1}{t^{2}}+\frac{1}{2t}+\frac{1}{t(e^{t}-1)}\right)\right\}(z).$
We can further simplify this expression by denoting
~$\widetilde{\nabla}_{\ell}^{(N)}:=\sum_{i=1}^{N}\sum_{j=1}^{4}{(-1)^{j-1}\Big{(}\nabla_{i;j}^{(N)}\Big{)}^{\ell}},\quad\ell\in\mathbb{Z}_{\geq 1}.$
For example, note that ~$\widetilde{\nabla}_{1}^{(N)}=0$. Then the RHS of (135) can be rewritten as
(136) $\sum_{\ell\geq 1}(-1)^{\ell}\left(\frac{\widetilde{\nabla}^{(N)}_{\ell+1}}{\ell+1}-\frac{\widetilde{\nabla}^{(N)}_{\ell+1}}{\ell}+\frac{\widetilde{\nabla}^{(N)}_{\ell}}{2\ell}\right)z^{-\ell}+\widetilde{\mathcal{L}}\left\{\sum_{\ell\geq 1}(-t)^{\ell}\frac{\widetilde{\nabla}^{(N)}_{\ell}}{\ell!}\left(-\frac{1}{t^{2}}+\frac{1}{2t}+\frac{1}{t(e^{t}-1)}\right)\right\}(z).$
Next, we want to take the coefficient-wise limit of the formal power series (135)–(136) in the regime $N\to\infty$, $N\theta\to\gamma$. We will use (128) and Lemma 6.6, which imply
(137) $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}{\widetilde{\nabla}_{\ell}^{(N)}}=(-1)^{\ell-1}\gamma^{\ell}-\ell\gamma m_{\ell-1},$
for all $\ell\in\mathbb{Z}_{\geq 1}$, if we set $m_{0}:=1$. As a result, the limit of the first term in (136) equals
(138) $\gamma\sum_{\ell\geq 1}\left(\frac{(-1)^{\ell}m_{\ell}}{\ell}+\frac{(-1)^{\ell-1}m_{\ell-1}}{2}+\frac{\gamma^{\ell}}{\ell+1}-\frac{\gamma^{\ell}}{\ell}-\frac{\gamma^{\ell-1}}{2\ell}\right)z^{-\ell}\\ =\widetilde{\mathcal{L}}\left\{\left(\gamma\cdot M(-t)-\frac{e^{t\gamma}-1}{t}\right)\left(\frac{1}{t}+\frac{1}{2}\right)\right\}(z),$
where $M(s):=1+\sum_{n\geq 1}{\frac{m_{n}}{n!}s^{n}}$ is the moment generating function of $\mu$. Similarly, due to (137), the limit of the second term of (136) is equal to
(139) ~$\widetilde{\mathcal{L}}\left\{\left(\gamma\cdot M(-t)-\frac{e^{t\gamma}-1}{t}\right)\left(-\frac{1}{t}+\frac{1}{2}+\frac{1}{e^{t}-1}\right)\right\}(z).$
We conclude that the sum of (138) and (139) is the limit of (136) (the RHS of (135)). On the other hand, the limit of the LHS of (135) follows from (132). By putting both of these results together, we obtain the identity
$\ln\left(1+\sum_{n\geq 1}\frac{(-1)^{n}c_{n}}{z^{\uparrow n}}\right)=\widetilde{\mathcal{L}}\left\{\left(\gamma\cdot M(-t)-\frac{e^{t\gamma}-1}{t}\right)\frac{1}{1-e^{-t}}\right\}(z),$
as the limit of (135) in the regime $N\to\infty$, $N\theta\to\gamma$; this is precisely the desired (115).
In order to finish the proof, we need to conclude that the relation between generating functions holds true for arbitrary sequences $m_{k}$, and not only for those of the form (128). Indeed, note that the two equations (114)–(115) lead to an expression for $\kappa_{n}$ as a polynomial in $m_{1},\dots,m_{n}$ over $\mathbb{C}(\gamma)$. On the other hand, (26) also leads to an expression of $\kappa_{n}$ as a polynomial in $m_{1},\dots,m_{n}$ over $\mathbb{C}(\gamma)$. Fix $n\in\mathbb{Z}_{\geq 1}$ and denote these polynomials by $\tilde{P}$ and $\tilde{P}^{\prime}$. There exists a polynomial $f\in\mathbb{C}[\gamma]$ such that both $P:=f\cdot\tilde{P}$ and $P^{\prime}:=f\cdot\tilde{P}^{\prime}$ are polynomials in $\gamma,m_{1},\dots,m_{n}$ over $\mathbb{C}$. The proof so far has shown that $P$ and $P^{\prime}$ agree when $\gamma\in(0,\infty)$ and the $m_{k}$’s are of the form (128). We claim that $P$ and $P^{\prime}$ are in fact identical, whence $\tilde{P}=\tilde{P}^{\prime}$, and the proof would be finished.
Fix an arbitrary partition $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{Y}(n)$ and define $\lambda_{n}^{(N)}\in\mathbb{Y}(N)$ as
$\lambda_{n}^{(N)}:=(\underbrace{\lambda_{1},\dots,\lambda_{1}}_{\frac{N}{n}+O(1)\text{ terms}},\dots,\underbrace{\lambda_{n},\dots,\lambda_{n}}_{\frac{N}{n}+O(1)\text{ terms}}).$
It is easy to show that the empirical measures (127) corresponding to the sequence $\big{\{}\lambda_{n}^{(N)}\big{\}}_{N\geq 1}$ satisfy the LLN in the sense of Definition 3.2, with corresponding $m_{1},m_{2},\cdots\in\mathbb{R}$ given by
[…]
As a result, $P(\gamma,m_{1},\dots,m_{n})=P^{\prime}(\gamma,m_{1},\dots,m_{n})$ on the following subset of $\mathbb{C}^{n+1}$:
$X:=\{(x_{0},x_{1},\dots,x_{n})\colon x_{0}\in(0,\infty),\\ x_{k}=\frac{1}{n(k+1)}\sum_{j=1}^{n}\sum_{s=0}^{k}\binom{k+1}{s}\left(\lambda_{j}-\frac{j\gamma}{n}\right)^{s}\left(\frac{\gamma}{n}\right)^{k-s},\text{ for}1\leq k\leq n,\ (\lambda_{1},\dots,\lambda_{n})\in\mathbb{Y}(n)\bigg{\}}.$
Notice now that $X$ is the image of $(0,\infty)\times\mathbb{Y}(n)\subset\mathbb{C}^{n+1}$ through the polynomial map $F=(f_{0},f_{1},\dots,f_{n})\in\mathbb{C}[x_{0},x_{1},\dots,x_{n}]^{n+1}$ defined by
$f_{0}:=x_{0},\qquad f_{k}:=\frac{1}{n(k+1)}\sum_{j=1}^{n}\sum_{s=0}^{k}\binom{k+1}{s}\left(x_{j}-\frac{jx_{0}}{n}\right)^{s}\left(\frac{x_{0}}{n}\right)^{k-s},\text{ for }1\leq k\leq n.$
Observe that $f_{0},f_{1},\dots,f_{n}\in\mathbb{C}[x_{0},x_{1},\dots,x_{n}]$ are algebraically independent, therefore the map $F$ is dominant (its image is dense in the Zariski topology), so it maps Zariski dense subsets of $\mathbb{C}^{n+1}$ into Zariski dense subsets. Thus, $X$ is Zariski dense, since it is the image of the Zariski dense subset $(0,\infty)\times\mathbb{Y}(n)\subset\mathbb{C}^{n+1}$. Consequently, $P$ and $P^{\prime}$ are identical, for being polynomial functions that coincide on the Zariski dense subset $X$, and the proof is concluded.
7. Comparison with the previous related results
Here, we discuss the relations between our results and related ones from the literature.
Even though the main topic of this paper is the asymptotic behavior of measures in the high temperature regime $N\to\infty,N\theta\to\gamma$, using the verbatim arguments, one can prove the analogous results in the fixed temperature regime, i.e. when $N\to\infty$ and $\theta$ is fixed. As the proofs are the same, we will only sketch the arguments and leave the details to the interested reader. In the case of a fixed $\theta>0$, the correct scaling of the empirical measure $\mu_{N}$ of $\mathbb{P}_{N}$ is as follows:
(140) $\tilde{\mu}_{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{\frac{\mathscr{L}_{i}}{N\theta}},$
where $\mathscr{L}_{i}:=\lambda_{i}-(i-1)\theta$, for $i=1,2,\dots,N$.
Definition 7.1.
Given a sequence $\vec{\kappa}=(\kappa_{n})_{n\geq 1}$, define $\vec{m}=(m_{n})_{n\geq 1}$ by
(141) $m_{\ell}:=\sum_{\Gamma\in\mathbf{L}(\ell)}\frac{\kappa_{1}^{1+\text{\#\,horizontal steps at height $0$ in $\Gamma$}}-(\kappa_{1}-1)^{1+\text{\#\,horizontal steps at height $0$ in $\Gamma$}}}{1+\text{\#\,horizontal steps at height $0$ in $\Gamma$}}\\ \cdot\kappa_{1}^{\text{\#\,horizontal steps at height larger than $0$ in $\Gamma$}}\cdot\prod_{j\geq 1}(\kappa_{j}+\kappa_{j+1})^{\text{\#\,steps $(1,j)$ in $\Gamma$}},\quad\text{for all $\ell\in\mathbb{Z}_{\geq 1}$}.$
We denote the map $\vec{\kappa}\mapsto\vec{m}$ by $\mathcal{J}_{\infty}^{\kappa\mapsto m}\colon\mathbb{R}^{\infty}\to\mathbb{R}^{\infty}$ and the compositional inverse $\vec{m}\mapsto\vec{\kappa}$ by $\mathcal{J}_{\infty}^{m\mapsto\kappa}:=(\mathcal{J}_{\infty}^{\kappa\mapsto m})^{(-1)}$.
Then we have the following result.
Theorem 7.2.
Let $\theta>0$ be fixed. Let $\{\mathbb{P}_{N}\}_{N\geq 1}$ be a sequence of probability measures with total mass $1$ on $N$-signatures and with small tails, and let $\{\tilde{\mu}_{N}\}_{N\geq 1}$ be their empirical measures, given by (140). Also, recall that $G_{\mathbb{P}_{N},\theta}(x_{1},\dots,x_{N})$ denotes the Jack generating function of $\mathbb{P}_{N}$. Suppose that the following two conditions are satisfied, for some constants $\kappa_{1}^{\infty},\kappa_{2}^{\infty},\cdots$:
  1. (1)
    $\lim_{N\to\infty}\frac{1}{N\theta}\frac{\partial_{1}^{n}}{(n-1)!}\ln\big{(}G_{\mathbb{P}_{N},\theta}\big{)}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=\kappa_{n}^{\infty}$ exists and is finite, for all $n\in\mathbb{Z}_{\geq 1}$,
  2. (2)
    $\lim_{N\to\infty}\frac{1}{N\theta}\partial_{i_{1}}\cdots\partial_{i_{r}}\ln\big{(}G_{\mathbb{P}_{N},\theta}\big{)}\big{|}_{(x_{1},\dots,x_{N})=(1^{N})}=0$, for all $r\geq 2$ and $i_{1},\dots,i_{r}\in\mathbb{Z}_{\geq 1}$ with at least two distinct indices among $i_{1},\dots,i_{r}$.
Then $\tilde{\mu}_{N}$ converges, as $N\to\infty$, in the sense of moments, in probability, to a probability measure $\mu_{\infty}$ with finite moments, denoted by $m^{\infty}_{n}$. Moreover, the sequences $\vec{\kappa}=(\kappa^{\infty}_{n})_{n\geq 1}$ and $\vec{m}=(m^{\infty}_{n})_{n\geq 1}$ are related to each other by $\vec{m}=\mathcal{J}_{\infty}^{\kappa\mapsto m}(\vec{\kappa})$. Finally, if $\sup_{n\geq 1}\big{|}\kappa_{n}^{\infty}\big{|}^{1/n}<\infty$, then $\mu_{\infty}$ is uniquely determined by its moments.
This theorem can also be generalized to the case where $\mathbb{P}_{N}$ are finite signed measures, as in Theorem 3.6.
Example 7.
The first few moments $m_{n}^{\infty}$ in terms of $\kappa_{n}^{\infty}$ are
$m_{1}^{\infty}$ $=\kappa_{1}^{\infty}-\frac{1}{2},$
$m_{2}^{\infty}$ $=\kappa_{2}^{\infty}+\big{(}\kappa_{1}^{\infty}\big{)}^{2}+\frac{1}{3},$
$m_{3}^{\infty}$ $=\kappa_{3}^{\infty}+3\kappa_{2}^{\infty}\kappa_{1}^{\infty}+\big{(}\kappa_{1}^{\infty}\big{)}^{3}+\frac{3}{2}\big{(}\kappa_{1}^{\infty}\big{)}^{2}-\frac{1}{4}.$
The fixed temperature regime was studied previously by Bufetov–Gorin [BG15] for $\theta=1$, and by Huang [Hua21] for arbitrary $\theta>0$. Bufetov–Gorin employed certain differential operators on the space of symmetric polynomials that were designed from the explicit determinantal formula for Schur polynomials. As a result, there is no obvious extension of their ideas to our setting, since Jack polynomials do not have determinantal formulas; moreover, the Cherednik operators that we use are differential-difference operators. On the other hand, the Jack generating function in the work of Huang is different from ours, namely, his version is a formal powers series of symmetric functions, not symmetric polynomials. Moreover, the hypothesis on the asymptotic behavior of the Jack generating function studied by Huang is much more complicated and technical than ours and it seems more difficult to verify in practice; this is a consequence of the fact that the techniques in that paper are different and rely on the analysis of the Nazarov–Sklyanin operators [NS13] that do not directly reflect the behavior of the model.111111 Huang’s technical assumptions involve infinite sums of differential operators with respect to the power-sum symmetric functions. They were derived from the explicit form of Nazarov–Sklyanin operators, which are differential operators in these power-sum variables. These operators are very useful in studying other models of discrete random $\beta$-ensembles, as shown in [Mol23, CDM23], but the model studied here is intrinsically related to the monomial expansions of Jack generating functions, and therefore Cherednik operators are much better suited to its analysis. Nevertheless, it is proved in [Hua21, Appendix B] that his assumptions are actually equivalent to our conditions in Theorem 7.2; in particular, this last theorem recovers the LLN proved in [Hua21].
As a sanity check, let us next deduce the formula obtained in [BG18, Theorem 2.4], [BG19, Eqn. (2.7) in Theorem 2.6] and [Hua21, Theorem 2.4] (expressed by extracting coefficients of certain generating functions) from our combinatorial formula (141).
Proposition 7.3.
Let $\vec{\kappa}=(\kappa_{n})_{n\geq 1}$ be arbitrary, let $R(z):=\sum_{n\geq 1}{\kappa_{n}z^{n-1}}$ and $(m_{n})_{n\geq 1}:=\mathcal{J}_{\infty}^{\kappa\mapsto m}(\vec{\kappa})$. Then, for any $n\in\mathbb{Z}_{\geq 1}$, we have
(142) $m_{n}=\frac{1}{n+1}[z^{-1}]\frac{1}{1+z}\left(z^{-1}+(1+z)R(z)\right)^{n+1}.$
Proof.
Let $F(z)=1+\sum_{i\geq 1}f_{i}z^{i}$ be a formal power series. It is a classical result in enumerative combinatorics (see e.g. [Sta99, Sections 5.3–5.4]) that the generating function for Łukasiewicz paths
$L(t):=1+\sum_{n\geq 1}\sum_{\Gamma\in\mathbf{L}(n)}t^{n}\prod_{j\geq 0}(f_{j+1})^{\text{\#\,steps $(1,j)$ in $\Gamma$}}$
satisfies the following functional equation:
$t\cdot L(t)=t\cdot F(t\cdot L(t)).$
In particular, it follows from the Lagrange inversion formula (see [Sta99, Theorem 5.4.2]) that
(143) $\frac{1}{n+1}[z^{-1}]\frac{1}{1+z}\left(z^{-1}\cdot F(z)\right)^{n+1}=\sum_{i\geq 0}^{n}\frac{(-1)^{i}}{i+1}\cdot\frac{i+1}{n+1}[z^{n+1-(i+1)}]\left(F(z)\right)^{n+1}\\ =[t^{n+1}]\sum_{i\geq 0}^{n}\frac{(-1)^{i}}{i+1}\left(t\cdot L(t)\right)^{i+1}.$
Note that $\left(t\cdot L(t)\right)^{i+1}$ can be interpreted as the generating series for the sequences of $i+1$ukasiewicz paths, where each of them starts with a horizontal step at height 00. By concatenating these paths and then removing the first horizontal step at height 00, we obtain a Łukasiewicz path of length $n$ (if looking at the coefficient of $t^{n+1}$) with $i$ marked horizontal steps at height 00. Therefore, we can rewrite the RHS of (143) as a sum over all Łukasiewicz paths of length $n$ with some number of marked horizontal steps at height 00 (we can mark $i$ steps in $\binom{h_{0}(\Gamma)}{i}$ ways, where $h_{0}(\Gamma)$ is the number of horizontal steps at height 00 in $\Gamma$), and then assigning the corresponding weight:
$\sum_{\Gamma\in\mathbf{L}(n)}\sum_{i=0}^{h_{0}(\Gamma)}\frac{(-1)^{i}}{i+1}\binom{h_{0}(\Gamma)}{i}f_{1}^{(\text{\#\,steps $(1,0)$ in $\Gamma$})\,-\,i}\prod_{j\geq 1}(f_{j+1})^{\text{\#\,steps $(1,j)$ in $\Gamma$}}.$
By choosing $F(z)=1+z(1+z)R(z)=1+\kappa_{1}z+\sum_{j\geq 1}{(\kappa_{j}+\kappa_{j+1})z^{j+1}}$, and applying the identity $\frac{1}{i+1}\binom{h_{0}(\Gamma)}{i}=\frac{1}{1+h_{0}(\Gamma)}\binom{1+h_{0}(\Gamma)}{i+1}$ in the above expression, we can express the RHS of (142) as
$\sum_{\Gamma\in\mathbf{L}(n)}\frac{1}{1+h_{0}(\Gamma)}\sum_{i=0}^{h_{0}(\Gamma)}(-1)^{i}\binom{1+h_{0}(\Gamma)}{i+1}\kappa_{1}^{\text{(\#\,steps $(1,0)$ in $\Gamma$})\,-\,i}\prod_{j\geq 1}(\kappa_{j}+\kappa_{j+1})^{\text{\#\,steps $(1,j)$ in $\Gamma$}}.$
This latter formula is clearly equal to the expression from Definition 7.1, due to the binomial formula, thus the proof is completed. ∎
Remark 11.
The formula in [BG18, Theorem 2.4] and [BG19, Eqn. (2.7)] is actually:
(144) $m_{n}=\frac{1}{n+1}[z^{-1}]\frac{1}{1+z}\left(z^{-1}+1+(1+z)R(z)\right)^{n+1}.$
Note that Eq. 144 differs slightly from Eq. 142, but this difference is a result of the simple fact that (when $\theta=1$) our empirical measures are $\frac{1}{N}\sum_{i=1}^{N}\delta\big{(}\frac{\lambda_{i}+1-i}{N}\big{)}$, while the empirical measures used in the previous references are $\frac{1}{N}\sum_{i=1}^{N}\delta\big{(}\frac{\lambda_{i}+N-i}{N}\big{)}$.
Remark 12.
As discussed in Section 5.1, the quantized $\gamma$-cumulants linearize the quantized $\gamma$-convolution. The limits of the quantized $\gamma$-cumulants when $\gamma\to\infty$, denoted $\kappa_{n}^{\infty}$, should therefore linearize the quantized convolution from [BG15]. In that paper, a quantized $R$-transform was also introduced and shown to linearize the quantized convolution; see [BG15, Theorem 2.9]. It turns out that the coefficients of that transform are not exactly $\kappa_{n}^{\infty}$, but linear combinations of them. For example, let $\mathfrak{m}$ be a probability measure with finite moments $m_{n}^{\infty}$ and let $(\kappa_{n}^{\infty})_{n\geq 1}:=\mathcal{J}_{\infty}^{m\mapsto\kappa}\big{(}(m_{n}^{\infty})_{n\geq 1}\big{)}$. Further, let $R_{\mathfrak{m}}^{quant}(z)$ be the quantized $R$-transform; then $R_{\mathfrak{m}}^{quant}(z)=R_{\mathfrak{m}}(z)+z^{-1}-(1-e^{-z})^{-1}$, where $R_{\mathfrak{m}}(z)$ is the $R$-transform of $\mathfrak{m}$, by virtue of [BG15, (2.4)-(2.5)]. From this formula and Example 7, we deduce
$R_{\mathfrak{m}}^{quant}(z)=\kappa_{1}^{\infty}+(\kappa_{2}^{\infty}+\kappa_{1}^{\infty})z+\left(\kappa_{3}^{\infty}+\frac{3}{2}\kappa_{2}^{\infty}+\frac{1}{2}\kappa_{1}^{\infty}\right)z^{2}+O(z^{3}),$
from which it is clear that even the first few coefficients of $R_{\mathfrak{m}}^{quant}(z)$ are already nontrivial linear combinations of $\kappa_{n}^{\infty}$.
We claim that the map $\mathcal{J}_{\infty}^{\kappa\mapsto m}$ from Definition 7.1 can be regarded as the limit of $\mathcal{J}_{\gamma}^{\kappa\mapsto m}$, as $\gamma\to\infty$. Indeed, in the high temperature regime $N\to\infty$, $N\theta\to\gamma$, the moments $m^{N}_{\ell}$ of the empirical measure $\mu_{N}$ have limits $m_{\ell}$ if and only if the moments $\tilde{m}^{N}_{\ell}$ of the empirical measure $\tilde{\mu}_{N}$ have limits $\tilde{m}_{\ell}$. In that case, these limits are related to each other by $m_{\ell}=\gamma^{\ell}\cdot\tilde{m}_{\ell}$. Similarly, the relation between the $\kappa_{\ell}$’s that appear in Definition 3.1 and the $\tilde{\kappa}_{\ell}$’s from Theorem 7.2 is given by $\kappa_{\ell}=\gamma\cdot\tilde{\kappa}_{\ell}$. Therefore, the formula (22), rewritten in terms of $\tilde{\kappa}_{\ell}$ and $\tilde{m}_{\ell}$, reads
(145) $\tilde{m}_{\ell}=\sum_{\Gamma\in\mathbf{L}(\ell)}\frac{\tilde{\kappa}_{1}^{1+h_{0}(\Gamma)}-(\tilde{\kappa}_{1}-1)^{1+h_{0}(\Gamma)}}{1+h_{0}(\Gamma)}\cdot\prod_{i\geq 1}(\tilde{\kappa}_{1}+i/\gamma)^{\text{\#\,horizontal steps at height $i$ in $\Gamma$}}\\ \cdot\prod_{j\geq 1}(\tilde{\kappa}_{j}+\tilde{\kappa}_{j+1})^{\text{\#\,steps$(1,j)$ in $\Gamma$}}(j/\gamma+1)^{\text{\#\,down steps from height $j$ in $\Gamma$}},\quad\text{for all $\ell\in\mathbb{Z}_{\geq 1}$},$
where $h_{0}(\Gamma):=\text{\#\,horizontal steps at height $0$ in $\Gamma$}$. Evidently, the formula (145) in the limit $\gamma\to\infty$ is exactly the formula (141) for the moments $m^{\infty}_{\ell}$, as claimed. This limit can also be understood by introducing a certain gradation on the polynomial algebra $\mathbb{C}[\gamma,\kappa_{1},\kappa_{2},\dots]$. If we define
$\deg_{1}(\gamma)=\deg_{1}(\kappa_{1})=\deg_{1}(\kappa_{2})=\dotsm=1,$
then the combinatorial formula (22) for the moments $m_{\ell}$ implies that $m_{\ell}\in\mathbb{C}[\gamma,\kappa_{1},\kappa_{2},\dots]$ has $\deg_{1}(m_{\ell})=\ell$, and $m_{\ell}^{\infty}$ is equal to the top degree term of $m_{\ell}$ followed by the substitutions $\gamma\mapsto 1$ and $\kappa_{n}\mapsto\kappa^{\infty}_{n}$.
It turns out that there is another gradation $\deg_{2}$ on the polynomial algebra $\mathbb{C}[\gamma,\kappa_{1},\kappa_{2},\dots]$ that allows to deduce a recent result of [BGCG22] from our formula (22), namely:
$\deg_{2}(\gamma)=0,\quad\deg_{2}(\kappa_{n})=n,\quad\forall\,n\geq 1.$
In order to motivate this gradation, recall that in the proof of Theorem 3.6, the moments $m_{\ell}$ were calculated as the limits
$\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\sum_{i=1}^{N}\big{(}\xi_{i}+\tilde{\xi}_{i}\big{)}^{\ell}\exp\left(\hat{F}_{N}\right)\bigg{|}_{(x_{1},\dots,x_{N})=(0^{N})},$
where $\xi_{i}$ are the Cherednik operators, $\tilde{\xi}_{i}$ are the Dunkl operators, and $\hat{F}_{N}=\hat{F}_{N}(x_{1},\dots,x_{N})$ is a formal power series such that
  1. (1)
    $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\frac{1}{N\theta}\frac{\partial_{1}^{n}}{(n-1)!}\hat{F}_{N}\big{|}_{(x_{1},\dots,x_{N})=(0^{N})}=\kappa_{n}$ exists and is finite, for all $n\in\mathbb{Z}_{\geq 1}$,
  2. (2)
    $\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\frac{1}{N\theta}\,\partial_{i_{1}}\cdots\partial_{i_{r}}\hat{F}_{N}\big{|}_{(x_{1},\dots,x_{N})=(0^{N})}=0$, for all $r\geq 2$ and $i_{1},\dots,i_{r}\in\mathbb{Z}_{\geq 1}$ with at least two distinct indices among $i_{1},\dots,i_{r}$;
see Eq. 51. The Cherednik operators preserve the degree of power series, while the Dunkl operators decrease the degree by $1$, therefore by introducing a formal variable $t$, the limit
$\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\frac{1}{N}\sum_{i=1}^{N}\big{(}\xi_{i}+t\cdot\tilde{\xi}_{i}\big{)}^{\ell}\exp\left(\hat{F}_{N}\right)\bigg{|}_{(x_{1},\dots,x_{N})=(0^{N})},$
is equal to the following $t$-deformation of the formula for $m_{\ell}$:
$\sum_{\Gamma\in\mathbf{L}(\ell)}\frac{\Delta_{\gamma}\left(x^{1+\text{\#\,horizontal steps at height $0$ in $\Gamma$}}\right)(t\cdot\kappa_{1})}{1+\text{\#\,horizontal steps at height $0$ in $\Gamma$}}\cdot\prod_{i\geq 1}(t\cdot\kappa_{1}+i)^{\text{\#\,horizontal steps at height $i$ in $\Gamma$}}\\ \cdot\prod_{j\geq 1}(t^{j}\kappa_{j}+t^{j+1}\kappa_{j+1})^{\text{\#\,steps $(1,j)$ in $\Gamma$}}(j+\gamma)^{\text{\#\,down steps from height $j$ in $\Gamma$}}.$
In this formula, the coefficient of $t^{d}$ has degree $\deg_{2}$ equal to the exponent $d$. Define $m_{\ell}^{\operatorname{top}}$ as the top homogenous degree part of the RHS of Eq. 22 with respect to the gradation given by $\deg_{2}$. From the formula above, it follows that
(146) $m^{\operatorname{top}}_{\ell}=\sum_{\Gamma\in\mathbf{L}(\ell)}\prod_{j\geq 0}\kappa_{j+1}^{\text{\#\,steps $(1,j)$ in $\Gamma$}}(j+\gamma)^{\text{\#\,down steps from height $j$ in $\Gamma$}}.$
This formula corresponds exactly to the limit
$\lim_{\begin{subarray}{c}N\to\infty\\ N\theta\to\gamma\end{subarray}}\frac{1}{N}\sum_{i=1}^{N}\big{(}\tilde{\xi}_{i}\big{)}^{\ell}\exp\left(\hat{F}_{N}\right)\bigg{|}_{(x_{1},\dots,x_{N})=(0^{N})},$
with the above assumptions on $\hat{F}_{N}$. The calculation of such a limit was one of the main calculations in [BGCG22], where the authors showed that it is the $\ell$-th moment of the limiting measure of a sequence of empirical measures of certain continuous $\beta$-ensembles defined by the Bessel generating functions. Their combinatorial formula was formulated in terms of set partitions, but is, in fact, equivalent to Eq. 146; indeed, [Xu23, Proposition 5.11] formulates it in terms of noncrossing set partitions and it can be translated to Łukasiewicz paths by using the well-known bijection in [NS06, Proposition 9.8]. Our discrete model is a quantization of the continuous model from [BGCG22] (see the introduction of [BG15]), yet it is still quite interesting that we can recover their exact results as corollaries of our proofs.